A Bibliography of Papers in Lecture Notes in Computer Science (2012): Volumes 7350–7399

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

(0, δ) [487]. (F, G) [158]. (k, r) [114]. 0 [527]. 1 [261, 527, 487]. 2 [229, 775, 878, 658, 180]. 2.5 [229]. 3 [570, 609, 229, 571, 190, 577, 582, 584, 468, 567, 180]. ^{+} [210]. α [131]. ∩ [160]. ΔΣ [474]. EL [284]. ELH_{R^{+}} [283]. ϵ [690]. F [812]. F_{2^{n}} [392]. GF(2^{m}) [386]. k [829, 832, 97, 68, 688, 466, 438]. K_{4} [141]. L [23, 389]. L_{p} [527]. λ [789]. N [532, 111]. O(log f) [112]. O(n^{3} log log n/ log^{2} n) [142]. P [520]. Φ [440]. q [92]. R^{m} [787]. SHACUN [560]. U_{2} [61]. x^{2^{n}−1k+1} + ax [392]. x^{n} − a [391].

-Anonymity-Based [438]. -Center [97]. -Clustering [114]. -Completeness [23]. -Cover [789]. -D [229, 261]. -Differential [690]. -Dimensional [658]. -Fragment [158]. -Functions [389]. -Gram [92]. -Hiding [440]. -Indistinguishable [688]. -level [829]. -Median [832]. -Minor [141]. -norm [527]. -Pad [812]. -party [111]. -Round [487]. -Secure [487]. -Visibility [131]. -XORSAT [68].

11th [1028, 997, 998, 996, 1034]. 12th [1022, 1019, 1012]. 13th [1029, 992, 1017, 1018]. 14th [1021]. 15th [1030]. 16th [1032, 1008, 1030]. 17th [1016, 1007]. 19th [1020, 990, 997].

2.0 [963]. 2.4GHz [277]. 20th [1014]. 22nd [893]. 23rd [989]. 24th [993]. 256 [447]. 26th [1006].

3 [499, 501]. 39th [1026, 1027]. 3rd [35].

4 [497]. 4th [1004].

5th [994, 997, 1009, 991, 1001].

6/TC [1029]. 60th [995, 1005]. 65th [1000]. 6LoWPAN [273]. 6LoWPAN/CoAP [273]. 6th [999].

7th [986, 987, 988, 1013].

8th [1024, 1025, 1011].

9th [1002, 1003].

= [298].

Abduction [210]. Abductive [196, 197]. Abelian [820]. Ability [358]. Abnormal [552]. above [410]. Abramowitz [247]. Abstract [113, 19, 254, 908, 28, 168, 859]. Abstraction [167, 714]. Abstractions [288]. Abuse [963]. Academic [263, 810]. Accelerating [444]. Accelerometer [474]. Access [753, 421, 752, 420, 422, 962, 950, 949, 649]. Accessibility [916]. Accessible [659, 37, 663, 650, 660]. Accumulator [949]. Accuracy [595, 190]. ACGs [21]. Achieving [690]. ACISP [1007]. Acquisition [479, 812]. across [967]. Action [995, 870, 517, 519]. Actions [869, 870, 244]. Activation [320, 329]. Active [560, 533, 519, 963, 157]. Activity [328, 386, 576]. ACTL [160]. Acyclic [640]. Ad [593, 103, 259, 260, 277, 105, 998]. Ad-Hoc [593, 259, 105, 998]. Adaptation [510, 1014]. Adaption [363]. Adaptive [512, 692, 7, 783, 777]. Adding [624]. Adequate [910]. ADHOC [998]. ADHOC-NOW [998]. Adjunction [231]. Ads [684]. Advanced [384, 618]. Advances [1012, 1002, 1003, 1001]. Adventure [402]. Adversary [487]. Aesthetic [44]. Affective [314]. Affinity [792]. Afraid [209]. Africa [1009]. AFRICACRYPT [1009]. After [399, 677]. Against [356, 498, 449, 487, 436, 485]. Agency [869, 865]. Agent [712, 771]. Agents [202, 858, 204, 866]. Aggregation [552, 269, 948, 770, 268]. Agile [759]. Aging [884]. Agnostic [27]. Agreement [102]. Agreements [860]. Aho [83]. Aid [336, 359]. Aided [993, 882]. Air [275]. Air-Cooled [275]. Aircraft [776]. AISC [997]. Alcohol [332]. Algebra [52, 625]. Algebraic [843, 382, 449, 301, 255]. Algorithm [83, 143, 829, 463, 562, 633, 334, 790, 142, 391, 251, 141, 101, 809, 355, 372, 786, 443, 806, 409, 357, 802, 783, 803, 346, 799, 150, 992]. Algorithmic [1005, 163]. Algorithms [791, 80, 828, 186, 146, 473, 140, 100, 339, 533, 545, 841, 390, 748, 54]. Aligning [5]. Alignment [793, 353]. Alike [711]. Alliance [240]. Allies [240]. Allocation [925, 519]. Almost [818]. Alpha [977]. Alpha-Corecursion [977]. Alphabets [741]. Alternating [72]. Alternative [648]. Amalgamability [290]. Ambiguities [915]. AMDO [1013]. among [332]. Amorphous [217]. amortizing [826]. Analog [461, 466, 802]. Analyses [167]. Analysing [264]. Analysis [450, 499, 421, 150, 938, 47, 566, 183, 334, 689, 124, 687, 615, 445, 564, 893, 163, 757, 906, 531, 937, 548, 466, 726, 120, 494, 247, 877, 166, 880, 715]. Analyzer [666, 895]. Analyzing [127, 611, 118]. Anaphoric [916]. Anatomy [233]. Anatomy-Based [233]. Ancestral [67]. Andersonian [868]. Android [960]. Angiogenesis [216]. Animal [267]. Animals [804]. Animated [584]. Animations [583]. Anisotropic [373]. Annealing [800]. Annotating [147]. Annual [1006, 989]. Anomaly [959, 706]. Anonymity [438]. Anonymization [424, 699]. Anonymizing [697]. Anonymous [113]. Antennae [104, 262]. Antennas [148]. Anthropometric [577]. Anthropomorphic [234]. Anti [961, 877]. Anti-counterfeiting [877]. Anti-malware [961]. Any [818]. Anycast [791]. APIs [750]. App [678]. Append [945]. Append-Only [945]. Applicability [280]. Application [496, 964, 1000, 869, 547, 549, 1016, 446, 545, 804, 778]. Applications [135, 753, 237, 752, 201, 147, 670, 763, 1006, 759, 745, 988, 1025, 960, 614, 305, 688, 7, 1012, 275, 781, 985]. Applied [554]. Applying [651, 348]. Approach [780, 230, 319, 832, 542, 759, 745, 663, 650, 532, 761, 531, 563, 686, 919, 885, 888, 859, 480, 524, 708, 922, 851]. Approaches [212, 15, 648]. Approximate [86, 362, 429]. Approximating [846, 60]. Approximation [80, 828, 829, 775, 133, 259]. April [1034]. Arbitrage [341]. Arbitrary [59, 268]. Architectural [7]. Architecture [751, 503, 571, 309, 8, 437, 338]. Architectures [306, 384, 303, 11]. Area [228, 959, 809]. Argument [205, 225]. Argumentation [205, 206, 200, 207, 28]. Arguments [483]. Arithmetic [286, 473, 298, 1004]. ARM11 [501]. Arousal [329]. Arrows [45]. Art [618, 119]. Arteries [186]. Articulated [1013]. Artificial [509, 511, 876, 786, 803]. Artimetrics [612]. ASDs [678]. Aspects [986, 412, 263, 876, 1012]. Assertion [762]. Assertion-Based [762]. Assessing [588]. Assessment [675, 964, 652, 760]. Assisted [774]. Assistive [679, 664]. Association [562, 531]. Associative [504, 556]. Assumption [440]. Assurance [5]. Asynchronous [708]. Asynchrony [111]. AT-Free [144]. Atomicity [715]. Attaching [613]. Attack [447, 449, 495, 437, 490, 485]. Attackers [271]. Attacking [943]. Attacks [498, 435]. Attention [517, 603, 519, 329]. ATtiny [492]. Attractor [334].

Attribute [422]. Attribute-Based [422]. Attributes [421]. Auctions [837, 842]. Audit [945]. Auditory [658, 325]. Audits [944]. Augmentation [764]. Augmenting [125]. August [1023, 1030]. Australasian [1007]. Australia [1007]. Austria [1017, 1018, 991]. Authentication [879, 935, 878, 934, 889, 426, 884, 427]. Authoring [660]. Authorization [212, 896]. Autism [678]. Auto [157]. Auto-active [157]. Automata [83, 728, 702, 729, 628, 57, 851, 853, 629, 975, 631, 632, 633, 624, 769, 635, 639, 72, 19, 158, 741, 1016, 625, 742, 727, 1026, 1027]. Automated [421, 164, 926, 191, 280, 718, 292, 999]. Automatic [289, 175, 636, 580, 578, 182, 704]. Automatically [39, 184, 297]. Automaton [637]. Autonomic [266]. Autonomous [322, 321]. Avatar [612]. Averaging [235]. Avoidance [270]. Aware [925, 372, 8, 942]. Awareness [745]. Axiomatic [256]. aying [634].

Baby [518]. Back [12, 31, 129, 214, 311, 393, 417, 619, 700, 722, 772, 872, 917, 932, 968, 983, 157]. Backdoors [411]. Background [368, 346]. Badness [941]. Balanced [528]. BAMs [540]. Band [475, 277]. Bane [966]. Banking [936]. Bar [39, 111]. Barcelona [1010]. Barn [325]. Barrier [146, 606]. Barriers [654]. Based [560, 753, 552, 421, 306, 752, 167, 500, 546, 423, 80, 286, 588, 505, 851, 230, 679, 811, 186, 379, 590, 366, 420, 935, 670, 775, 566, 511, 592, 360, 296, 633, 316, 696, 712, 571, 92, 348, 362, 221, 784, 374, 344, 422, 474, 336, 650, 599, 554, 792, 795, 577, 809, 779, 355, 889, 7, 117, 786, 343, 222, 762, 261, 466, 119, 254, 327, 456, 892, 521, 539, 233, 428, 182, 310, 866, 11, 877, 949, 10, 438, 189, 806, 180, 939, 610, 541, 480, 748, 125, 485, 665, 358, 798, 326, 561, 427, 318, 608]. Based [793, 617, 746, 777, 781, 526, 373, 810, 803, 353, 346, 361, 337, 778, 922, 896]. Bases [59, 255]. Basic [405]. Basis [16, 147, 336, 359]. Batch [482, 555]. Batch-End [555]. Bayes [221]. Bayesian [538, 316, 178, 524]. Be [107, 518, 209]. Bed [553]. Bee [786]. Bees [786]. Before [399]. Behavior [594]. Behaviour [733]. Behaviours [469]. Belgrade [998]. below [410]. Benchmarking [923]. Bergen [1028]. Berkeley [993]. Berlekamp [391]. Berlin [1022, 1012, 1011]. Bernd [1000]. Betrayed [684]. Better [818]. between [676, 717, 568, 457, 219, 132, 880]. Beyond [702, 1005, 155, 68]. Bi [531]. Bi-clustering [531]. Bias [378]. Biased [478]. Biautomata [738]. Biblus [647]. Biclique [143]. BICS [1001]. Bidimensional [108]. Bilinear [387]. Bimonoids [632]. Binaries [388]. Binary [826, 388, 221]. Binding [792]. BINER [530]. Binomials [391, 392]. Biohybrid [1010]. Bioinspired [509]. Biology [153]. Biomechanical [222]. Biomedical [781, 994]. Biometric [884]. Biometrics [876]. Biomimetic [503, 514, 1010]. Bipartite [64]. Bipedal [505]. Birth [397]. Birthday [936, 995, 1005, 1000]. Bisection [180]. Bits [891]. Black [823]. Black-Box [823]. BLAKE [499]. Blanket [362]. Blind [676, 656, 670, 661, 654, 674, 664, 673, 665]. Block [489, 478, 492, 70, 445]. Blueprints [307]. Blum [733]. BLUP [804].

Bochum [1004]. Body [506, 573, 327]. Boltzmann [350]. Bonaire [1033, 1032]. bond [468]. Book [644]. Boolean [849, 383, 161, 709, 334, 638, 72, 381]. Boon [966]. Borders [967]. Bound [61]. Bounded [85, 831, 63, 296, 856, 635]. Bounding [103]. Bounds [816, 846, 96, 410, 259, 487, 440]. Box [823]. Braga [1021]. Brain [328, 319, 313, 324, 189, 1001]. Brain-Inspired [313]. Brandenburger [970]. Brandenburger-Keisler [970]. Breast [229, 235, 220, 232, 216, 217, 227, 224, 226, 233, 234, 219, 223, 231, 225, 996]. Breeding [804]. Bremen [997]. Bridging [717]. Broadcast [449, 70, 105]. Browser [696]. Browser-Based [696]. Bruijn [838]. Bruteforce [895]. Budgets [837]. Building [931, 912]. Bunching [551]. Burrows [94]. Bus [551]. Business [2, 4, 5, 1, 614, 8, 985]. Business-Aware [8].

C [157, 719]. C5.0 [534]. CA [993]. Cache [471]. Calculation [614]. Calculemus [997]. Calculi [285]. Calculus [898, 201, 900, 295, 200, 977, 23, 24, 29]. Call [552]. Camellia [447, 446]. Camellia-256 [447]. Camera [319, 670, 665]. Camera-Based [670, 665]. Campus [673]. Can [107, 518, 521, 54]. Canada [1014]. Cancer [220, 547]. Canterbury [987, 1029]. Capacitated [97]. Capacity [315]. Capsules [724]. Captioning [677]. Capturing [5]. Carbon [930, 788]. Card [426]. cardiology [374]. Cardiovascular [316]. Cards [427]. Caregiving [674]. Carotid [186]. Cascade [230]. Cascade-Based [230]. Case [964, 546, 66, 941, 554, 199, 543, 434]. Case-Based [546, 554]. Catalytic [326]. Categorial [19, 908]. Categorical [46]. Categories [980]. Category [22, 40]. Causal [845]. Causality [858]. CAV [993]. CCA2 [452]. CDCL [254]. CDCL-Based [254]. Cell [153, 476]. Cells [623]. Cellular [629, 614]. Censored [559]. Censorship [696]. Center [97, 929, 928, 921]. Centered [522]. Centers [1031, 920, 919, 930, 275]. Centrality [757]. Centralized [742]. Centres [922]. Cerebellar [504]. Cerebral [320]. Certain [71]. Certificate [238]. Certification [6]. Challenges [239, 1, 281, 152]. Challenging [648]. Change [871, 532]. Changes [578, 645]. Changing [99, 871, 648]. Channels [886]. Characteristic [472, 746]. Characteristic-2 [472]. Characterization [848, 220, 178]. Characterizing [71]. Characters [572]. Charge [646, 479]. Chargeback [929]. Charts [39]. Chatting [570, 571]. Chebyshev [444]. Checker [469, 713]. Checking [714, 59, 153, 1020, 296, 203, 169, 716, 719, 710, 715]. Chest [579]. Children [651, 329]. China [1024, 1025, 1002, 1003, 1001]. Chips [467]. Chisholm [864]. Choice [603, 604]. Chop [736]. Choquet [526]. Chosen [936]. Christofide [150]. Chromatography [793]. Church [855]. CIAA [1016]. CICM [997]. Cipher [488]. Ciphers [450, 489, 492, 445, 446]. Circuit [478, 61, 462, 952, 802]. Circuits [156, 480]. Circulating [553]. City [136]. Class [348, 356]. Classes [732, 145]. Classical [818, 766]. Classification [525, 360, 124, 568, 221, 374, 369, 533, 227, 355, 224, 545, 806, 534, 358, 338, 526, 805, 356]. Classifier [775, 221, 359, 556, 806, 557]. Classifiers [550, 162]. Client [685]. Client-Side [685]. Clients [694]. Clinical [218, 216]. Clock [114]. Closed [833, 558]. Clothes [304]. Cloud [679, 929, 926, 674, 749, 616, 307, 694]. Cloud-Based [679]. Clouds [927]. Cluster [895]. Clustering [560, 821, 137, 542, 348, 784, 336, 359, 114, 599, 749, 543, 561, 531]. Clustering-Based [561]. Clusters [230]. Clutter [566]. CMCS [1034]. CMOS [479, 226]. CMPs [471]. CMS [1029]. Co [525, 746]. Co-occurrence [746]. Co-training [525]. Coal [778]. Coalgebra [978]. Coalgebraic [982, 1034]. Coalgebraically [971]. Coalgebras [975]. Cochlear [372, 182]. Code [515, 433, 719]. Codes [379, 878, 658, 378, 377]. Coding [887, 376, 786]. Cofinite [29]. Cognitive [591, 41, 314, 522, 611, 1001]. COGPARSE [313]. Coherence [471]. Coherent [254]. Coinductive [981]. coli [814]. Collaboration [242, 920, 655]. Collaborative [750, 4, 959, 745, 654, 810]. Collapsible [853, 854]. Collecting [675]. Collection [269]. Collections [91]. Colliculus [325]. Collision [498, 497]. Colloquium [1026, 1027, 990]. Colluding [271]. Collusion [111]. Colocated [1034]. Colony [786]. Colorful [79]. Colouring [100]. Combination [186]. Combinational [480]. Combinations [849, 638]. Combinatorial [73, 989]. Combinatorics [636]. Combinators [250]. Combinatory [18, 26]. Combined [334, 518]. Combining [707, 576]. Commercial [807]. Committee [806]. Committees [966]. Common [85, 25, 314]. Communication [817, 676, 572, 947, 990, 301, 698, 571, 514, 757]. Communications [107, 277, 266, 1029]. Community [650]. Community-Based [650]. Commutative [885]. Compact [492]. Comparative [594, 784, 575]. Comparing [534]. Comparison [499, 228, 223, 804, 541, 231]. Complementary [494]. Complete [900]. Completeness [23, 29]. Complex [622, 126]. Complexities [732]. Complexity [817, 465, 824, 766, 730, 57, 733, 447, 407, 258, 87, 768, 734, 61, 397, 405, 990, 67, 735, 911, 736, 738, 72, 640, 1021, 739, 416, 743, 401, 400]. Compliance [927]. Components [79]. Composed [817]. Composite [11, 10]. Composition [308, 870, 9]. Compositional [708]. Comprehensive [788]. Compressed [93, 92]. Compression [451, 567]. Compromise [945]. Computability [407]. Computation [1023, 938, 145, 444, 168]. Computational [986, 412, 768, 1015, 640, 75]. Computationally [869]. Computations [528]. Compute [305]. Computed [107]. Computer [1028, 414, 767, 988, 997, 643, 882, 1034, 988, 993]. Computer-Aided [882]. Computers [1017, 1018]. Computing [143, 201, 834, 929, 62, 926, 251, 840, 1024, 1025, 724, 674, 309, 94, 304, 326, 810, 789]. Concept [564, 531, 524, 537]. Conceptual [303, 8]. Concurrency [716]. Concurrent [171]. Condition [441]. Conference [1028, 986, 1022, 987, 1006, 1029, 999, 1024, 1025, 997, 1032, 998, 993, 1014, 1017, 1018, 1009, 1016, 1013, 1012, 1011, 1010, 991, 1007, 1001]. Conferences [1030]. Confidence [185]. Confidentiality [880]. Confinement [423].

Conflict [776]. Conflicting [78]. Congestion [108, 942]. Congestion-Aware [942]. Congruential [855]. Conjecture [253]. Conjunctive [737]. Connected [139, 65]. Connectivity [98, 104]. Consecutive [86]. Consensus [424, 630]. Conserved [503]. Consistency [73, 219]. Consistent [181]. Constant [729, 81, 947, 184]. Constant-Time [81]. Constants [335]. Constrained [569, 694]. Constraint [154, 293, 903, 410, 339, 192, 365]. Constraints [560, 210, 819, 828, 904, 43, 197]. Constructing [65, 378, 537]. Construction [86, 245]. Constructions [457, 487]. Consumers [929]. Consumption [159, 549]. Contact [882]. Contact-Less [882]. Container [6]. Container-Level [6]. Containment [847, 291]. Content [663, 120, 125]. Content-Based [125]. Context [306, 971, 734, 858, 336, 24, 910]. Context-Free [971, 734, 24, 910]. Contextual [740, 519]. Contextualized [868]. Contingency [558]. Continual [691]. Continuous [974, 584, 275, 610]. Continuum [40]. Contracts [924]. Contrary [864, 863, 211]. Contrary-To-Duties [211]. Contrary-to-Duty [863]. Contrast [229, 218, 220, 216, 217, 192, 219]. Contrast-Enhanced [218, 220, 216, 219]. Contribution [195]. Contributions [587]. Control [753, 421, 503, 752, 420, 631, 512, 322, 422, 554, 213, 950, 949, 610, 886, 321]. Controllable [616]. Controlled [904, 480]. Controller [506, 159, 322]. Controlling [908]. Conundrum [960]. Conversational [570]. Convertible [457]. Convex [822, 149]. Convolutional [367]. Cooled [275]. CoolEmAll [921]. Cooperative [16, 266]. Coordinate [261]. Coordination [22]. Copenhagen [1023, 1030]. Coprocessors [895]. Coq [237]. Corasick [83]. Core [546]. Corecursion [977]. Corpus [118]. Correcting [379]. Correlated [264]. Correlation [532]. Correspondence [838]. Cortex [327, 370]. Cortical [332, 191]. Cost [243, 951, 272]. Cost-Effective [243]. Costs [822]. CoTAGs [21]. Count [392]. Counter [852, 294]. Counterexample [706]. counterfeiting [877]. Counters [66]. Coupled [758]. Courses [656]. Cover [99, 139, 141, 401, 789]. Coverage [830, 146, 133, 262, 703]. Covering [422]. Coverings [58]. Covers [60]. Covert [886]. CPM [989]. CPU [925]. CPU- [925]. Crawler [799]. Creation [109]. Critiquing [601]. Cross [90, 435, 493, 526]. Cross-Document [90]. Cross-Oriented [526]. Cross-Site [435]. Cross-Terms [493]. Crosstalk [465]. Crosswatch [666]. Crowd [611]. Crowdsourced [747]. CRT [495]. CRT-RSA [495]. Cryptanalysis [489, 448, 491, 493]. Crypto [398]. Cryptographic [454, 939, 945]. Cryptography [1033, 376, 1032]. Cryptology [1009]. CsI [226, 223]. CsI-CMOS [226]. CSP [318]. CSP-Based [318]. CSPs [816]. CSR [988]. CT [182]. Cues [332]. Currency [341]. Current [461, 207]. Curry [838]. Curve [389]. Customer [936]. Customer-Chosen [936]. Customers [415]. Cuteforce [895]. CWFM [558]. CWL [882]. Cyber [618]. Cyberworlds [608]. Cycle [283]. Cycle-Restricted [283]. Cyclic [164, 377]. Cycling [99]. Cylindrical [255].

D [180, 570, 609, 229, 775, 878, 571, 190, 577, 582, 261, 584, 468, 567, 180]. D.C [943]. DAC [422]. DAC9 [516]. DAISY [660]. Dance [514]. d’Andratx [1013]. Dark [229, 689]. Data [675, 269, 879, 1033, 559, 938, 850, 538, 47, 164, 46, 447, 929, 430, 965, 692, 967, 1031, 529, 597, 1032, 359, 547, 43, 564, 920, 614, 549, 548, 429, 1012, 1011, 919, 892, 928, 930, 950, 275, 545, 888, 804, 438, 272, 557, 353, 337, 778, 922, 921, 1006]. Data-Based [353]. Data-Minimizing [879]. Database [330, 652, 556, 370]. Databases [423]. Datalog [847]. Dataset [585]. DB2 [434]. DBSec [1006]. DBT [231]. DC [1031, 923]. DC-Benchmarking [923]. DCFS [1021]. De-amortizing [826]. Deaf [642, 655, 677]. Dealing [645]. Debugging [273, 718]. Dec [435]. Decentralized [432, 213]. Deciding [112, 731]. Decimal [470]. Decision [852, 298, 316, 348, 227, 953, 563, 778]. Deck [529]. Decoding [878, 464]. Decomposition [485, 255]. Dedicated [1005, 1000]. Deduction [26]. Defeasible [861]. Deficit [329]. Deficit/Hyperactivity [329]. Defining [971]. Deformable [1013]. Deformation [579, 181, 176]. Degree [846, 831, 64]. Degree-Distributions [64]. Delaunay [135]. Delay [264, 63]. Delay-Tolerant [264]. Delays [349]. delegatable [459]. Deletion [140]. Deletions [79]. Delivery [927]. Denmark [1023, 1030]. Dense [270]. Density [232]. DEON [1028, 210]. Deontic [1028, 767, 864, 862, 861, 870, 868]. Dependability [4]. Dependencies [237, 899, 423, 902]. Dependency [19, 119]. Dependent [332, 832, 183]. Deploying [938]. Deployment [308]. Deposition [581]. Depth [566]. Derivation [911]. Derivative [351]. Derived [380, 180]. Description [284, 287, 535]. Descriptional [730, 738, 72, 739, 1021]. Design [791, 833, 471, 698, 474, 359, 476, 462, 747, 681, 1008, 475, 470, 267, 464, 409, 802, 708]. Designated [459]. Designing [957, 951]. Detail [584]. Detecting [552, 170, 230]. Detection [113, 366, 175, 959, 687, 756, 578, 706, 367, 551, 875, 120, 437, 271, 665, 561, 318, 746, 341, 799]. Detective [226]. Detector [479, 573, 226, 223]. Detector-Driven [573]. Detectors [115]. Determination [219]. Determine [598]. Determining [780, 597]. Deterministic [139, 562, 100, 639, 158]. Determinization [976]. Developing [680]. Development [763, 658, 644, 761, 563]. Deviations [551]. Devices [675, 670, 492, 952]. DFAs [640]. Diagnosing [167, 547]. Diagnosis [795, 550]. Diagnostic [231]. Diagonalization [970]. Diagram [50, 41]. Diagrammatic [41, 51, 987]. Diagrams [38, 35, 52, 987, 36, 44, 136, 43, 49, 51]. Diameter [831]. Diameters [137]. Dictionary [564, 365]. Diffeomorphic [176]. Difference [380]. Differences [315]. Different [784, 345]. Differential [690, 332, 489, 498, 491, 777]. Differentially [691, 692]. Difficulty [364]. Diffusion [373]. Digital [2, 504, 228, 647, 218, 220, 251, 883, 217, 242, 875, 219, 231]. Digraph [628]. Digraphs [65]. DIKE [494]. Dimensional [559, 692, 658, 84, 805]. Dimensionality [541]. Direct [800, 485]. Directed [102, 835]. Directional [148, 262, 234]. Directions [416, 672]. Directly [176]. Disabilities [658, 645]. Disabled [646]. Disclosure [686]. Discontinuity [907].

Discontinuous [775, 782]. Discovering [128]. Discovery [431, 121, 123]. Discrete [827, 251, 339]. Discrete-Time [339]. Discretionary [949]. Discriminant [124, 539, 805]. Discrimination [807, 358]. Discussion [871]. Disease [75]. Disjoint [144]. Disorder [329]. Disorders [678]. Distance [528, 260, 132]. Distinguishers [446]. Distinguishing [906]. Distributed [690, 728, 113, 691, 712, 840, 444, 260, 213, 105]. Distribution [377]. Distributions [47, 827, 64, 595]. Distributive [24]. Distributivity [282]. Ditransitives [913]. Divergence [360]. Diversification [754]. Division [946, 801]. DM [451]. DML [997]. DNA [931]. DNA-Inspired [931]. Do [148]. Document [878, 644, 89, 88, 90, 247]. Documenting [248]. Documents [124, 127]. Domain [146, 834, 477, 461, 760, 166, 118, 754]. Dominating [401]. Domination [70]. Dosimetry [233]. Dot [30]. Dot-types [30]. Double [451, 104]. Double-Length [451]. Doubling [831]. Downey [415]. Drift [524]. Driven [752, 509, 477, 316, 759, 761, 573, 313, 557]. DS [562]. DSL [719]. DSL-Generated [719]. DTI [190]. Dual [220, 466, 608, 381]. Dual- [466]. Dual-Energy [220]. due [760]. Duplicated [799]. During [399]. Duties [211]. Duty [99, 864, 863]. Dynamic [753, 102, 642, 869, 308, 462, 343, 905, 916, 309, 226, 11, 192, 884, 949, 814]. Dynamical [522]. Dynamics [845, 569, 20, 265]. Dysarthria [680]. Dysfunction [329].

e-Services [2]. EA [383]. EA-Equivalence [383]. Early [112, 397]. Earthquake [677]. East [677]. ECDSA [482]. ECG [374]. ECoG [331]. Economy [788]. Ecosystem [697]. Ecosystems [2]. Edge [79, 60]. Edge-Multi-Covers [60]. Editing [243]. Education [651, 647, 653, 645, 649]. Educational [651, 682, 650]. EEG [591]. Effect [47, 217, 265, 327, 605]. Effective [285, 145, 596, 243, 128]. Effectiveness [274]. Effects [884]. Efficiency [328, 924, 887, 226, 919, 921]. Efficient [83, 421, 819, 159, 827, 471, 360, 385, 93, 461, 84, 1031, 940, 927, 452, 953, 952, 548, 475, 472, 945, 922]. Effort [760, 454]. Effort-Release [454]. eHealth [894]. Eigengene [806]. Eigengene-Based [806]. Elastic [505]. Electric [507]. Electricity [549]. Electroactive [509]. Electroencephalogram [329]. Electronic [317, 602]. Elements [513, 528]. Eleven [936]. Eliminating [913]. Elimination [363]. Elliptic [389]. Elucidating [41]. Embedded [462, 309]. Embedding [862]. Embeddings [96]. Embodied [323]. Embodiment [521]. Emergence [517]. Emission [930]. Emotional [570]. Emperor [304]. Empirical [173, 937, 541]. Empty [149]. enabled [428]. Enabling [656, 959]. Encoder [465]. Encoding [283, 463]. Encodings [50]. Encounters [401]. Encrypted [888]. Encryption [496, 452, 453, 455, 454, 885, 880]. End [962, 555]. End-user [962]. Energetic [816]. Energy [924, 931, 220, 920, 549, 927, 919, 930, 475, 922, 921, 1031]. Energy-Efficiency [924, 921]. Energy-Efficient [922]. Enforcement [434].

Enforcing [896]. Engagement [591, 600]. Engine [755]. Engineering [985, 1022, 1, 1000, 761, 747, 886]. Enhanced [763, 36, 218, 220, 216, 217, 602, 950, 310, 192, 219, 493]. Enhancement [119]. Enhancements [657]. Enhancing [1019]. Enlarging [280]. Enriching [46]. Ensemble [792, 550]. Entailment [291]. Enterprise [1]. Entity [122]. Entropic [358]. Entropy [189, 537]. Environment [570, 609, 319, 263, 571]. Environmental [929, 121]. Environments [243, 315]. Enzymatic [520]. Ephemeral [441]. Epistemic [203]. EPR [296]. EPR-Based [296]. Equal [731]. Equality [339]. Equals [630]. Equation [709, 388]. Equations [350, 841]. Equivalence [469, 383]. Error [379, 175, 616]. Error-Controllable [616]. Errors [883]. Escherichia [814]. Essays [995, 1005, 1000]. Essence [249]. ESSLLI [1023]. Establishment [306]. Estimates [47, 591]. Estimation [569, 349, 553, 580, 260, 347]. Estimator [189]. Estonia [1034]. ETAPS [1034]. Ethical [964, 967]. Ethics [966, 965]. Ethnic [232]. Euclidean [801]. Euler [50, 35]. Eulerian [628]. Evading [696]. Evaluating [675, 761, 274]. Evaluation [319, 492, 216, 583, 681, 928, 370, 788]. Event [201, 571, 200, 756, 33]. Event-Based [571]. Every [936]. Evidence [48, 329, 798]. Evoked [320]. Evolution [323, 758, 777]. Evolutionary [790]. Evolving [556]. Exact [633]. Examiners [127]. Example [614]. Exchange [441]. Excluded [89]. Executions [979]. Existence [392]. Expansion [177, 444, 127]. Expectations [648]. Expected [852]. Experience [220, 677, 518]. Experiences [239]. Experiencing [579]. Experimental [303, 664, 223]. Experiments [510, 305, 518, 433, 612]. Expert [807]. Explanation [706]. Explanations [295, 41, 604]. Exploiting [926]. Exploration [507, 456, 123, 804]. Exploring [597, 697, 962]. Exponential [141, 614, 796]. Exponentiation [384]. Exposure [225]. Expression [370]. Expressions [63, 371]. Extended [483, 113, 66]. Extending [241]. Extension [626, 286, 385]. Extensions [978, 300]. External [707]. External-Memory [707]. Extractability [26]. Extraction [237, 908]. Extractor [478]. Extrema [258]. Extreme [342, 343, 356, 337]. Eye [655]. Eyesight [674].

Face [612]. Facial [583, 577, 370]. Facilitate [148]. Facility [829, 101, 991]. Factor [334, 326]. Factorization [592, 365]. Factors [609, 41]. Failure [167, 198, 115]. Fair [170]. Fall [675]. Fall-Related [675]. Familiarity [916]. Families [833, 253, 741]. Family [500, 377]. Farm [267]. Fast [633, 91, 479, 582, 222, 188, 189]. Faster [140, 443, 78]. Fault [948, 831, 97, 695, 550, 428]. Fault-Tolerant [831, 695]. FAuST [718]. FC [1033, 1032, 253]. FC-Families [253]. FCM [336]. FCM-Based [336]. FCSRs [500]. FDES [781]. Feature [811, 362, 374, 876, 227, 795, 779, 371, 364, 358, 357, 363]. Features [128, 371]. Februray [1032]. Federated [930]. Federations [307]. Feedback [506, 835, 117, 464]. Feedforward [335]. Feistel [446]. Fellows [1005, 415, 402, 401, 399]. FEMTO [91]. Few [57]. Fewer [699]. FFDM [231]. FG [1030]. Field [229, 228, 961, 681, 386, 390, 176, 231]. Fields [385, 251, 1004]. File [893]. Filter [785, 573, 324, 485, 896]. Filtering [810]. Final [57]. Finalists [499]. Financial [938, 1033, 1032]. Finding [387, 69, 669]. Fine [716]. FinFET [466]. Fingerprints [882]. Finitary [973]. Finite [285, 382, 57, 736, 1004, 386]. Finite-Field [386]. Finite-Valued [285]. Finland [992, 989]. Finnish [126]. Firewalls [896]. Firing [351]. Firms [2]. First [702, 15, 1031, 1010, 458]. First-Order [702]. Fish [803]. Fix [456]. Fixed [702, 408, 835, 396]. Fixed-Parameter [408, 835, 396]. Flat [294]. Flexible [305, 472]. Flip [467]. Flip-Flop [467]. Floor [672]. Flop [467]. Flow [184, 593, 553, 43, 781, 373]. Flows [791, 176]. Fluidized [553]. Flyby [399]. Flying [503]. FOL [208]. Foraging [516]. Forbidden [67]. Force [812]. Forensic [883]. Forest [366, 536]. Forests [685]. Forgeable [110]. Forgeries [876]. Form [382, 19, 181, 392, 176, 962]. Formal [237, 623, 712, 912, 252, 1021, 531, 156, 207, 246, 718, 248, 1030]. Formalizing [253]. Formally [381]. Formats [241, 648]. Forms [970, 764, 126]. Formulae [387]. Formulas [846]. Formulation [18, 128]. Forward [362]. Foundations [865]. Four [801]. Four-Quadrant [801]. Fourier [286]. FPGAs [472]. FPT [415, 141]. Fractional [100]. Fragment [158]. Fragmentation [438]. Framework [690, 675, 34, 935, 763, 240, 623, 316, 593, 138, 242, 322, 303, 515, 428, 718, 371, 433, 274, 896]. France [986, 1006]. Frankl [253]. Fréchet [132]. Free [824, 971, 290, 732, 186, 734, 144, 24, 181, 962, 49, 176, 910]. Free-Form [176, 962]. Free-Hand [186]. Frequency [461, 479, 372, 277, 364]. Frequent [558]. Fresh [300]. Friendly [661, 428]. Front [13, 32, 502, 523, 544, 565, 586, 607, 620, 641, 662, 683, 701, 723, 744, 765, 773, 794, 815, 836, 857, 873, 874, 881, 890, 897, 918, 933, 53, 954, 969, 984, 74, 95, 116, 130, 151, 172, 193, 215, 236, 257, 278, 299, 312, 333, 354, 375, 394, 395, 404, 418, 439, 460, 481]. Frontal [329]. FSM [463]. FSMD [469]. Full [228, 24, 313, 231]. Fully [496, 113, 182]. Function [819, 500, 828, 775, 61, 451, 336, 359, 608, 351, 783]. Function-Based [608]. Functional [423, 813]. Functioning [623]. Functions [817, 382, 383, 830, 161, 59, 62, 457, 389, 380, 782, 381, 345, 814]. Functors [973, 978]. Fusion [227]. Future [666, 1, 416]. Fuzzy [780, 287, 486, 344, 369, 336, 359, 355, 782, 781].

Gain [512]. Galore [398]. Game [424, 818, 851, 442]. Game-Based [851]. Game-Theoretic [442]. Games [108, 159, 856, 109, 769, 211, 919]. Gap [717]. Gapping [22]. Garbled [952]. Gastric [547]. Gated [463]. Gathering [272]. Gating [467]. Gauge [166]. Gaze [597, 655]. GECKOmmender [589]. Gene [75]. General [56, 540, 564, 527]. Generalised [864, 376]. Generalization [516]. Generalized [780, 17, 138, 907, 487, 458, 526, 909]. Generated [719]. Generating [295, 126]. Generation [450, 749, 952, 718]. Generator [478, 485]. Generic [563]. Genetic [463, 156, 357, 802]. Geodesic [173, 174]. Geometric [133, 260]. Geometrical [219]. Germany [1022, 997, 1004, 1012, 1011].

Gestural [572]. GHT [528]. Giant [839]. Girls [711]. Global [579, 201, 4, 173, 800]. GLUON [500]. GOAL [204]. Goals [919]. Goes [879]. Good [892]. Gossiping [711]. Governance [213]. GPS [667]. GPU [710]. Gradient [188, 373]. Graduate [38, 399]. Grained [716]. Gram [92]. Grammar [775, 903, 92, 912, 11, 535, 1030]. Grammar-Based [775, 92, 11]. Grammars [14, 17, 737, 19, 914, 740, 908]. Granular [581]. Grapes [807]. Graph [135, 845, 80, 67, 145, 69, 100, 981, 11, 841, 409, 541]. Graphic [47]. Graphics [37]. Graphs [45, 552, 143, 833, 64, 432, 140, 144, 70]. Great [677]. Greedy [96]. Green [926]. GreenIT [923]. GreenSLAs [924]. Grid [712]. Grishin [898, 900]. Gröbner [255]. Grocery [674]. Grøstl [499]. Ground [295]. Grounded [869]. Group [293]. Grouped [39]. Groups [820, 727]. Growing [346]. GSM [940]. Guarantee [406]. Guaranteeing [4]. Guards [99]. Guides [602].

H.264 [880]. H.264/SVC [880]. Halting [413]. Hand [186, 366, 54, 617]. Handheld [678]. HANDi [52]. Handwriting [876, 812, 884]. Haptic [763, 608, 617]. Haptic-Enhanced [763]. Hard [671, 677, 443]. Hard-of-Hearing [677]. Harder [934]. Hardness [85, 80]. Harnessing [673]. Harvesting [475]. Hash [500, 455, 498]. Hashing [389]. Hashtag [746]. Hausa [901]. Having [57]. head [742]. Heads [637]. Headway [551]. Health [546]. Healthcare [317]. Hearing [671, 643, 655, 677]. Heart [369]. Heat [350]. Heavy [691]. Height [729]. Held [997]. Helping [1017, 1018]. Helsinki [992, 989]. Hemodynamic [327]. Henstock [782]. Hermitian [378]. Heterogeneous [34, 895, 9]. Heuristic [791, 538]. Hidden [685, 610]. Hidding [428]. Hiding [440]. Hierarchical [560, 124, 573, 179, 338]. Hierarchically [712]. Hierarchy [741, 370]. High [229, 559, 957, 301, 461, 479, 703, 937, 226, 472, 470, 541]. High-Coverage [703]. High-Level [301]. High-Yield [937]. Higher [289, 653, 136, 84, 645]. Higher-Order [289, 84]. Highly [451]. HIGHT [489]. Hilbert [331]. Histogram [360, 692]. Histogram-Based [360]. Histology [187]. Hitters [691]. HNN [801]. Hoc [593, 103, 259, 260, 277, 105, 998]. Holling [813]. Holling-Type [813]. Home [598]. Homology [147]. Homomorphic [496]. Homomorphisms [69]. Homonyms [110]. Homotopy [613]. Honeybee [514]. Honor [995]. Hood [901]. Hop [831]. Hop-Diameter [831]. Horizontal [438]. Hot [529]. Hough [366]. Howard [838]. HPC [931]. Huang [331]. Huangshan [1024, 1025]. Human [579, 956, 571, 585, 568, 574, 575, 522, 611, 576]. Human-Protocol [956]. Humanoids [518]. Humans [570]. Hybrid [790, 644, 344, 796, 726, 435, 272, 223]. Hyper [639]. Hyper-minimization [639]. Hyperedge [42]. Hyperintensional [905].

IAL [358]. IC [468]. IC3 [155]. ICALP [1026, 1027]. ICCHP [1017, 1018]. ICDM [1012]. ICEBERG [491]. Iceland [990]. ICIC [1024, 1025]. ICWE [1022]. IDDM [545]. Ideal [638]. Identification [566, 774, 343, 78].

Identifiers [110, 100]. Identify [127]. Identifying [654]. Identity [209]. If [965]. IFIP [1006, 1029]. Ifrance [1009]. II [1027, 813, 1018, 1003]. IJCAR [999]. Illative [18]. Illicit [965]. Illuminance [578]. Illumination [578]. ILP [480]. Image [184, 183, 994, 177, 585, 174, 178, 187, 188, 877, 189, 179, 180, 176, 125, 458]. Image-Based [877]. Images [218, 875]. Imaging [229, 319, 251, 217, 996, 223]. Imitation [514, 518]. Immediate [115]. Immersion [145]. Immune [802]. Impact [964, 761, 400, 231]. Impaired [676, 647, 37, 643, 669, 667]. Implant [372, 182]. Implementation [34, 504, 478, 492, 384, 325, 467, 785, 928, 472, 464, 30, 1016]. Implemented [520]. Implementing [653, 895]. Implicative [197]. Implicatures [16]. Implicit [768]. Importance [199]. Important [493, 657]. Importing [20]. Impossible [489]. Improve [595, 476]. Improved [829, 379, 190, 680, 43, 259, 749, 446]. Improving [572, 498, 391, 599, 601, 488, 557, 710]. Impulse [361]. Impulsive [270]. Imputing [529]. IMS [889]. IMU [667]. In-Home [598]. In-Network [269]. Inclusion [627, 656]. Inclusive [657]. Incomplete [556]. Incorporate [596]. Incorporating [777]. Increasing [587, 115]. Incremental [155]. Incrementally [161]. IND-CCA2 [452]. Indented [46]. Independent [848, 838]. Index [86, 88]. India [1008]. Indicative [532]. Indicator [777]. Indicator-Based [777]. Indirect [450]. Indistinguishable [688]. Individual [41, 315]. Individuals [332]. Indoor [668]. Indoor/Outdoor [668]. Induced [143, 144]. Induction [298]. Inductive [155]. Industrial [1012]. Industry [549]. Inequality [166]. Inertial [667]. Inference [423, 987, 878, 535]. Inference-Usability [423]. Infinitary [977]. Infinite [849, 741]. Infinity [157]. Influence [555]. Influenced [41]. Information [990, 764, 177, 126, 795, 368, 809, 812, 261, 991, 798, 781, 337, 1007]. Infrastructure [273]. Initial [220]. Initialization [802]. Initiative [173]. Injection [450]. Ink [242]. Innovative [807]. Input [862, 480, 555]. Input-Output [862]. Insight [517]. Insights [518]. Inspired [931, 623, 317, 313, 324, 561, 321, 1001, 508]. Installing [960]. Instance [541, 537]. Instance-Based [541]. Instances [699]. Instantiate [204]. InStep [652]. Institution [653]. Instructions [384]. Integer [286, 946]. Integral [526]. Integrals [782]. Integrated [322, 222]. Integrating [307]. Integration [508, 647, 325, 243]. Integrator [516]. Integrity [197]. Intelligent [322, 1024, 1025, 997]. Intended [39]. Intense [925]. Intensity [186, 222]. Intensity-Based [186, 222]. Intensive [2]. Intention [204]. Intentions [202]. Inter [3]. Inter-organizational [3]. Interacting [33]. Interaction [651, 956, 213, 522, 608, 617]. Interactions [587, 585]. Interactive [970, 597, 810, 947]. Interconnects [465]. Interdisciplinary [400]. Interest [590, 810]. Interface [680]. Interfaces [319]. Interference [270, 103]. Interferences [148]. Interleaver [464]. Internal [740, 982]. International [1028, 986, 1022, 35, 987, 1026, 1027, 994, 1029, 1020, 990, 1019, 999, 988, 1024, 1025, 1031, 997, 1032, 1021, 998, 993, 996, 1014, 1017, 1018, 1009, 1016, 1004, 1034, 1013, 1011, 1010, 1008, 991, 1002, 1003, 1001, 1030]. Internet [943, 925, 273]. Interpolant [169]. Interpolants [702, 162]. Interpolating [720]. Interpolation [816, 290, 84, 610]. Interpretation [671]. Interpreting [47]. Interrogatives [906]. Intersecting [57]. Intersection [666]. Interval [199, 798]. Intra [182]. Intra-operative [182]. Introduction [52, 768, 646, 770]. Invariants [166]. invasive [745]. Inverse [730, 615, 181, 345]. Inverse-Consistent [181]. Investigating [604]. Investigation [49]. Investigations [883]. Investment [937]. Invisible [658, 54]. Involves [351]. IRBs [966]. IRFC [991]. Irrationality [206]. Island [901, 834]. Island-Hood [901]. ISM [277]. ISNN [1002, 1003]. Isomorphism [820]. ISP [941]. Issues [985, 261]. Italian [330]. Itemsets [558]. Iterated [899, 902, 115]. Iteration [902]. Iterative [797]. IVC [480]. IVUS [187]. IVUS-Histology [187]. IWDM [996].

J [995]. Jacobians [376]. Japan [677]. Java [164, 428]. JavaScript [762]. JH [499]. Jitter [479]. Join [423]. Joinery [613]. Joint [999]. JSART [762]. Judgment [770]. July [1028, 986, 1022, 987, 1006, 1026, 1027, 994, 1020, 990, 1019, 992, 988, 1024, 1025, 997, 989, 1021, 998, 993, 996, 1014, 1017, 1018, 1009, 1016, 1004, 1013, 1012, 1011, 1010, 1008, 991, 1007, 1002, 1003, 1001, 893]. Jump [349]. Jumping [511]. June [990, 999]. Justify [604].

Keccak [499]. Keisler [970]. Kernel [337]. Kernelization [406]. Key [441, 455, 454, 446, 888]. Key-Leakage [441]. Keys [891]. Keyterm [119]. Keyword [837, 126]. Kinect [577]. Kinect-Based [577]. Kinetic [135]. Kleene [902]. Knapsacks [443]. Knee [173]. Knowledge [2, 823, 453, 313, 123, 798, 557]. Knowledge-Driven [313, 557]. Knowledge-Intensive [2]. Known [446]. Known-Key [446]. Kolmogorov [824]. Kralendijk [1033, 1032]. Krämer [1000]. Kullback [360]. Kurtosis [361]. Kyoto [927].

l [634]. Labeling [191]. Labelled [285, 972]. LACL [986]. Lambda [977]. Lambek [898, 900, 23, 914, 24, 29]. Landmark [508]. Language [330, 15, 912, 126, 741, 246, 29]. Languages [848, 849, 732, 622, 1026, 1027, 734, 855, 735, 911, 736, 638, 739, 740, 910]. Laplacian [841]. Large [239, 91, 584, 708]. Large-Scale [239]. Latency [268]. Latent [882]. Latest [644]. LaTeX [661]. Lattice [350]. Lawvere [970]. Lax [978]. Layer [784, 351]. Layered [437]. Layout [44, 43]. Leading [345]. Leads [126]. Leaf [805]. Leakage [441, 455]. Leakage-Resilience [455]. Leaker [462]. Learn [518]. Learner [244]. Learning [34, 899, 504, 538, 161, 642, 540, 513, 369, 597, 654, 163, 244, 342, 117, 343, 518, 1011, 539, 310, 866, 806, 797, 365, 541, 657, 537, 910, 356, 337, 340]. Learning-Based [117]. Leasing [101]. Least [686, 555]. Lecture [1023].

Lectures [657, 1023]. LED [445]. Left [64, 371]. Left-Perfect [64]. Legal [869]. Leibler [360]. Length [731, 451, 443, 535]. Lengths [85]. Lens [509]. Lesion [228]. Less [112, 577, 882]. Letter [699]. Level [6, 279, 296, 301, 241, 584, 829, 889]. Leveraging [169]. Libraries [246]. Library [626, 647]. License [367]. Life [923, 33, 399]. Lifetime [259]. Lightweight [500, 489, 445]. Like [22, 517]. Like-Category [22]. Likelihood [334]. Limitations [430]. Limited [114, 741, 743]. Limits [611]. Line [707, 33]. Linear [83, 143, 819, 286, 146, 86, 388, 298, 904, 70, 625, 464, 841, 610, 166]. Lines [45, 134]. Linguistic [955]. Linguistics [986]. Linked [305]. Linz [1017, 1018]. Lip [642]. Lipschitz [335]. Liquid [793]. Listing [89]. Literature [546]. Living [623, 1010]. Ljubljana [1023, 1030]. Load [343]. Local [100, 261, 630, 192, 335, 777, 576, 353, 347]. Locality [911, 805]. Localization [785, 882, 519, 520, 271]. Locating [811]. Location [829, 175, 425, 940, 962, 310]. Location-Based [310]. Lock [171]. Logging [945]. Logic [1028, 284, 167, 1023, 287, 767, 864, 769, 862, 861, 869, 204, 18, 196, 870, 203, 197, 199, 254, 26, 866, 867, 868, 49, 995]. Logic-Automata [769]. Logic-Based [866]. Logical [212, 14, 865, 726, 371, 986]. Logics [766, 632, 199, 982, 419]. Login [958]. Logistic [534]. Logs [127]. Long [143, 891]. Longest [85]. Longitudinal [183]. Look [300]. Looks [934]. looping [297]. Low [465, 96, 447, 463, 478, 692, 461, 479, 467, 476, 462, 882, 951, 272, 788]. Low-Carbon [788]. Low-Cost [951]. Low-Dimensional [692]. Low-Power [478]. Low-Resolution [882]. Low-Stretch [96]. Low-Voltage [478]. Lower [846, 96, 61]. LP [829, 832]. LP-Rounding [829, 832]. LTL [294, 160, 158]. Ludics [15]. LZW [93]. LZW-Compressed [93].

m2n [123]. M5’ [534]. MAC [450, 422, 261]. Machine [848, 795, 342, 343, 527, 356, 1011]. Machine-Independent [848]. Machines [539, 1010, 28, 541, 781, 337]. Madrid [1031]. Magic [606]. Magnitude [175]. Mai [1031]. Mainstream [647]. Makes [205]. Mallorca [1013]. Malware [436, 961]. Mammograms [228]. Mammographic [221, 233]. Mammography [218, 227, 795, 222, 231]. Man [980]. Manage [571, 959]. Management [985, 4, 764, 1, 926, 920, 8]. Managing [239, 307]. Manchester [999]. MANETs [274]. Manifold [173]. Manipulated [176]. mantis [223]. Manual [893]. Many [536]. Maple [252]. MapReduce [693]. Maps [387, 613]. March [1033, 1032, 1034]. MAREC [118]. Marek [995, 195, 194]. Margin [476]. Marker [577]. Marker-Less [577]. Market [54]. Markov [852, 685, 362, 610]. Markovian [349]. Masquerade [437]. Mass [553, 795, 224, 793]. Masses [227]. Match [727]. Matching [81, 642, 82, 633, 93, 84, 989, 90, 132]. Matchings [64]. Material [235, 581]. Math [414, 591, 659]. Mathematical [246, 247]. Mathematics [237, 997, 244, 280, 403, 248, 402, 660]. Matrix [592, 342, 365, 710]. Matrix-Vector [710]. Matroid [828]. Matter [12, 31, 129, 214, 311, 393, 417, 619, 700, 722, 772, 872, 917, 932, 968, 983, 13, 32, 502, 523, 544, 565, 586, 607, 620, 641, 662, 683, 701, 723, 744, 765, 773, 794, 815, 836, 857, 873, 874, 881, 890, 897, 918, 933, 53, 954, 969, 984, 74, 95, 116, 130, 151, 172, 193, 215, 236, 257, 278, 299, 312, 333, 354, 375, 394, 395, 404, 418, 439, 460, 481]. max [355]. Maximal [824]. Maximization [819, 828, 334, 480]. Maximizing [878]. Maximum [146, 86, 77, 149, 259, 537]. May [3]. McEliece [453]. McPatom [715]. mctau [717]. MDC [497]. MDC-4 [497]. MDM [434]. Mean [368, 444, 796]. Means [514, 120]. Measure [183, 364]. Measurement [579, 275, 555, 357, 963, 667]. Measurements [577]. Measures [552, 568, 327]. Measuring [941]. Mechanism [41, 894, 930, 370]. Mechanisms [842]. Media [658]. Median [832]. Meet [490]. Meet-in-the-Middle [490]. Melodramas [403]. Membrane [326]. Memoir [194]. Memoriam [621]. Memory [925, 707, 513, 952, 324, 315]. Memory-Efficient [952]. Memory-Intense [925]. MEMS [474, 475]. Mental [48]. Mesh [567, 180]. Mesh-Derived [180]. MESI [471]. Message [450, 238, 258, 443, 487, 105]. Messages [39]. Metabolic [814]. Metabolomics [793]. Metaplastic [618]. Meters [951]. Method [816, 854, 366, 184, 360, 707, 378, 348, 596, 883, 369, 191, 359, 342, 804, 168, 326, 793]. Methodologies [749]. Methodology [961, 776]. Methods [827, 190, 529, 547, 1034, 804, 347]. Metrics [831, 941]. Michael [1005]. Microblogging [594]. Microcalcifications [230]. Microcontrollers [951]. Middle [980, 490]. Migration [926, 434]. Mike [402, 401, 399]. MIMO [786]. Min [355]. Min-max [355]. Mine [778]. Minicomplexity [725]. Minimal [640, 78]. Minimalism [20]. Minimalist [911, 17]. Minimally [910]. Minimization [532, 708, 83, 639]. Minimizing [879, 852, 146]. Minimum [137, 79, 60, 149, 535, 268, 561, 789]. Mining [562, 122, 92, 547, 532, 549, 531, 551, 429, 558, 1012, 1011, 436, 705, 778]. Minor [833, 67, 141]. Minor-Ancestral [67]. Minor-Closed [833]. Minors [409]. Missing [529, 582]. Mitigate [435]. Mixed [349, 527]. Mizar [237]. MKM [997, 241, 243]. ML [562]. ML-DS [562]. MLDM [1011]. Mobile [675, 879, 670, 593, 109, 998, 515, 664, 677, 952, 272]. Mobile-Type [677]. Mobility [264]. MOD [62]. MOD-Functions [62]. Modal [982, 177, 178]. Modalities [56, 24]. Model [570, 106, 752, 850, 714, 590, 420, 653, 153, 296, 759, 221, 713, 721, 422, 760, 350, 813, 203, 809, 516, 761, 452, 314, 115, 234, 169, 716, 719, 458, 610, 268, 318, 710, 526, 105, 715, 370, 341, 814, 1020]. Model-Driven [752, 759, 761]. Model-to-Model [760]. Modeled [465]. Modeling [559, 505, 463, 593, 595, 3, 1014, 324, 522, 364, 611, 302, 353, 352, 346, 337, 921]. Modelling [301]. Models [753, 469, 899, 588, 685, 787, 3, 600, 516, 575, 233, 584, 29, 982, 758, 705, 345]. Moderate [443]. Modes [226]. Modest [717]. Modified [467, 813, 351]. Modular [384]. Modularity [290]. Modulo [282]. Moments [224]. Monadic [847, 288]. Monitor [809]. Monitoring [201, 691, 598, 267, 275, 353, 352]. Monoidal [980]. Montagovian [898, 20].

Montreal [1014]. Moore [345]. Moral [865]. Morocco [1009]. Morphologically [126]. Most [622]. Motion [609, 579, 574, 582, 1013, 576]. motor [512]. Motzkin [286]. Movement [146, 17, 580, 320]. Movement-Generalized [17]. Movements [588, 320]. Moving [324]. MRI [186, 173, 227, 222, 192]. Multi [938, 955, 754, 60, 177, 790, 242, 771, 178, 574, 785, 367, 889, 573, 188, 742, 440, 842, 777, 803, 370, 346, 356]. Multi-Agent [771]. Multi-class [356]. Multi-database [370]. Multi-domain [754]. Multi-head [742]. Multi-level [889]. Multi-modal [177, 178]. Multi-object [346]. Multi-Objective [790, 777, 803]. Multi-Party [938, 242]. Multi-Prime [440]. Multi-resolution [188]. Multi-robot [785]. Multi-scale [367]. Multi-unit [842]. Multi-view [574, 573]. Multi-word [955]. Multidimensional [73]. Multidisciplinary [991]. Multilayered [233]. Multimedia [1029]. Multimodal [663, 227, 783]. Multimode [353]. Multimodes [352]. Multiparty [817]. Multiple [159, 486, 508, 82, 837, 858, 93, 595, 322, 537, 910, 105, 814]. Multiple-Controller [322]. Multiple-Message [105]. Multiplication [385, 710]. Multiplier [461, 470]. Multipliers [386, 472]. Multiresolution [179]. Multitape [637]. Multithreaded [170]. Multivariate [1005]. Muscles [509, 511]. Musculoskeletal [505]. Museum [588]. Museums [589]. Musibraille [656]. Music [656]. Mutual [177, 785, 795, 337]. MVP [935]. My [287, 962, 401]. Myths [648].

Nantes [986]. Nashville [994]. Natural [332, 15, 367]. NAVCOM [676]. Navigation [508, 513, 515, 673, 668, 521, 667]. Near [86]. Nearest [138]. Nearly [823]. Need [965]. Needed [966]. Needs [651, 682, 1017, 1018]. Negation [198]. negative [197]. Neighbor [138]. Neighborhood [199]. Neighbors [596]. Nested [853, 632]. Nested-Words [853]. Net [300]. Nets [301]. Network [106, 269, 587, 791, 833, 316, 109, 887, 334, 774, 325, 369, 259, 355, 265, 928, 275, 335, 338, 351, 341, 361]. Networked [1, 309]. Networking [266]. Networks [264, 269, 587, 102, 270, 538, 420, 263, 959, 98, 593, 323, 104, 344, 349, 103, 336, 359, 757, 998, 367, 796, 314, 260, 9, 75, 272, 534, 1002, 1003, 105, 814]. Neural [334, 774, 344, 349, 369, 336, 359, 367, 796, 355, 314, 335, 534, 1002, 1003, 358, 338, 351, 341, 361]. Neuro [508]. Neuro-inspired [508]. Neurobiologically [321]. Neurobiologically-Inspired [321]. Neurocomputing [797]. Neuromorphic [504]. News [58, 746]. News-Topic [746]. Next [279, 416]. Niederreiter [452]. NIRS [327]. Nizhny [988]. No [820, 880]. Node [46, 833]. Node-Oriented [46]. Node-Weighted [833]. Nodes [103, 121]. NOF [817]. Noise [476, 606, 555]. Nominal [631]. Non [16, 170, 766, 947, 175, 297, 571, 745, 374, 583, 564, 459, 427]. Non-cardiology [374]. Non-classical [766]. Non-delegatable [459]. Non-interactive [947]. Non-invasive [745]. Non-looping [297]. Non-rigid [175]. Non-tamper [427]. Non-termination [170, 297]. Non-vectorial [564]. Non-Verbal [571, 583]. Nonassociative [24]. Nondeterminism [729, 743]. Nonlinear [350, 800, 485, 526]. Nonnegative [365]. Nonrigid [179]. Nonterminal [737]. Norm [123, 527]. Normal [820, 382, 19]. Normalization [190]. Normative [210, 767, 208, 859]. Norms [995, 871, 866, 860]. Norway [1028]. Notes [1023]. Notions [98]. Nouns [25, 668]. Novel [562, 790, 348, 221, 812, 314, 518, 793, 803]. Novgorod [988]. NP [900]. NP-Complete [900]. NSW [1007]. NTRU [449]. Number [77, 473, 784, 737, 470, 782, 390]. Numerical [614, 520]. NURB [774]. NuSMV [721]. NVC [570].

OAI [173]. Obfuscation [425]. Object [566, 324, 346]. Objective [825, 790, 814, 777, 803]. Objects [507, 811, 1013]. Obligation [871, 867]. Obligations [864, 863]. Oblivious [442, 953]. Obstructions [145]. Occasion [995, 1005, 1000]. occurrence [746]. Offering [678]. Offset [478]. Omega [63]. Omega-Rational [63]. Once [59]. One [852, 348, 644, 351, 356]. One-Against-One [356]. One-Class [348]. One-Counter [852]. One-Layer [351]. Online [828, 420, 633, 101, 603]. Only [945]. Ontologies [283]. Ontology [590, 316, 10]. Ontology-Based [590, 10]. OOXML [893]. Open [985, 205, 849, 415, 276, 618]. Openbox [34]. OpenFst [626]. Opening [751]. OpenRichServices [308]. Operating [226]. Operation [643, 23]. Operational [974]. Operations [730, 734, 736, 72, 563, 168]. operative [182]. Operator [328, 319]. Opportunistic [263, 265]. Optical [319]. Optimal [387, 64, 66, 88, 440]. Optimality [904]. OPTIMIS [922]. Optimization [218, 903, 790, 784, 344, 574, 800, 786, 188, 930, 475, 326, 777, 783, 803]. Optimizations [615]. Optimized [269, 359]. Optimum [277]. Oracle [458]. Orchestration [305]. Order [702, 289, 136, 84]. Ordered [798]. organizational [3]. Orientation [318]. Orientations [80]. Oriented [306, 46, 309, 746, 526, 799]. Origin [965]. Orthogonal [786, 224, 42, 802]. Oscillation [361]. Osteoarthritis [173]. Outcomes [591]. Outlier [561]. Output [862, 474, 555]. Outsourced [950, 949]. Outsourcing [430, 438]. Over-Sampling [474]. Overlay [98]. Owl [325]. Oxford [1020].

P [634]. P2P [674]. PA [996]. Packet [896]. Packing [819]. Pad [812]. Padding [688]. Pages [799]. Pair [748]. Pair-Wise [748]. Pairs [67, 142]. Panel [965]. Paper [871]. Papers [1033, 990, 1031, 1032, 1034, 1030]. Paradigm [525]. Paradigms [651]. Parallel [469, 714, 709]. Parallelism [118]. Parallelizability [856]. Parallelizing [713]. Parameter [408, 835, 184, 396, 448]. Parameterized [407, 413, 139, 397, 405, 140, 410, 416, 401, 716, 400, 409]. Parameters [219, 364]. Pareto [542]. Paris [1006]. Parity [846, 856]. Parsing [903, 313]. Part [1026, 1027, 997, 1017, 1018, 1002, 1003]. Partial [235, 582, 132, 555, 880, 537]. Participation [751]. Participatory [650]. Particle [784, 574, 785, 573, 324]. Particles [784]. Partitioned [429].

Partitioning [79, 463, 87, 477]. Partitions [149]. Partners [751, 650]. Party [938, 946, 242, 111]. Passion [403]. passive [276]. Passphrases [955]. Password [427]. Password-Based [427]. Past [294, 518, 199]. Patch [703, 340]. Patent [117, 119, 123, 127, 125, 118]. Path [80, 942]. Paths [143, 173, 144, 142]. Pathwidth [408]. PatMedia [125]. Pattern [306, 82, 93, 89, 84, 989, 90, 355, 1011, 524, 358]. Pattern-Based [306]. Patterns [552, 165, 320]. Payment [694]. PC [742]. Pebbles [624]. CoAP [273]. Hyperactivity [329]. Outdoor [668]. Stateful [455]. SVC [880]. TC [1029]. Penrose [345]. People [676, 671, 670, 37, 680, 580, 1017, 1018, 677, 605]. Perception [609, 682, 611]. Perfect [64, 727]. Performance [328, 878, 492, 926, 217, 601, 488, 466, 884, 880, 557, 788]. Performances [595]. Performing [965]. Periodic [269]. Periods [320]. Permission [867]. Permissions [960]. Permutation [392]. Permutations [981]. Personal [959, 764]. Personalised [589]. Personalization [751, 1014]. Personalized [587, 602]. Personalizing [233]. Persons [665]. Perspective [442, 75, 797]. Persuasive [205, 206]. Perturbation [821]. Petri [300]. PETS [1019]. Phantom [235, 234]. Phase [479, 331, 786]. Phased [783]. Phenogrammar [906]. Philadelphia [996]. Phones [676]. Physically [617]. Physically-Based [617]. Piccolo [445]. Pie [135]. Pilot [510]. PINs [936]. Pivot [528]. Pixel [46, 188]. Plain [824]. Planar [833, 262]. Planes [380]. Planner [671]. Plans [672]. Plant [788, 805]. Plate [367]. Platform [276, 618]. Platforms [301]. Plays [403]. PLL [479]. Plots [46]. Plural [915, 909]. Point [186, 616, 248, 608]. Point-and-Write [248]. Point-Based [186]. Pointers [292]. Points [45, 702, 582]. Polarized [898]. Policies [212, 896]. Policy [962]. Poly [786]. Polygon [834]. Polygonal [834, 584]. Polymer [509]. Polynomial [820, 177, 344, 133, 910]. Polynomial-Time [820, 133]. Polynomials [702]. Polysynthetic [912]. Polytopes [149]. Ponzi [937]. Popularization [414]. Population [515]. Port [43, 1013]. Porto [1016]. Portugal [1021, 1016]. Pose [569, 576]. Poses [568]. Position [103, 327]. Possible [891, 861, 27]. Post [468]. Post-bond [468]. Postmodern [937]. Power [925, 971, 463, 60, 478, 477, 926, 461, 467, 476, 462, 115, 788]. Power-Aware [925]. Practical [694]. Practice [518]. Pre [216, 182, 458]. Pre- [182]. Pre-clinical [216]. Pre-image [458]. Predator [813]. Predator-Prey [813]. Predicate [892]. Predicate-Tree [892]. Predict [328, 591]. Prediction [792, 555, 810]. Predictive [595, 715]. Predicts [332]. Preference [592]. Prefix [824, 853]. Prefix-Free [824]. Prefrontal [332, 327]. Prehistory [396]. Preliminary [127]. Preparation [320, 557]. Preprocessing [359, 406]. PRES [469]. Presence [315]. Present [936]. Presentation [250]. Presentations [973]. Preserve [438]. Preserving [693, 948, 957, 695, 431, 894, 951, 429, 694]. Presheaves [972]. Presupposition [18]. Pretty [892]. Prey [813]. Prime [440]. Primer [405]. Principle [977, 535]. Print [658]. Prior [119]. Prior-Art-Search [119].

Priority [66, 707]. PRISM [693]. Privacy [424, 690, 693, 948, 1006, 957, 425, 695, 431, 894, 951, 429, 962, 892, 694, 950, 438, 1007, 419, 1019]. Privacy-Preserving [693, 948, 695, 431, 894, 951, 429]. Private [691, 685, 692, 940, 888]. Probabilistic [974, 150, 165, 625, 188, 866, 710, 356]. Probabilities [796]. Problem [791, 832, 97, 413, 903, 591, 138, 133, 141, 776, 521, 400, 111]. Problems [822, 767, 415, 790, 410, 259, 245, 524]. Procedure [298]. Proceedings [1026, 1027, 1017, 1018, 1002, 1003, 1028, 986, 1022, 987, 1006, 994, 1029, 1020, 1019, 992, 999, 988, 1024, 1025, 997, 989, 1021, 998, 993, 996, 1014, 1009, 1016, 1004, 1013, 1012, 1011, 1010, 1008, 991, 1007, 1001]. Process [4, 364, 352]. Processes [843, 974, 852, 623, 301, 353]. Processing [809, 41, 340]. Processors [501]. Production [734]. Productivity [761]. Profile [931]. Profiles [605, 436]. Profiling [685, 549]. Profit [3]. Prognostic [559]. Program [602]. Programming [825, 1026, 1027, 256, 339, 527]. Programs [995, 170, 154, 164, 165, 862, 171, 252, 196, 197, 953, 937, 292]. Progress [1009, 1008]. Project [922, 656]. Projection [805]. Projective [380]. Projects [414, 997]. Proof [238, 766, 947, 49]. Proof-Theoretical [49]. Proofs [239, 164, 455]. Propagation [590, 258]. Properties [955, 4, 113, 67, 323, 197]. Property [590, 903]. Property-Based [590]. Propositional [199]. Protected [958]. Protection [712]. Protein [792, 75]. Protein-Protein [792]. Protocol [956, 471, 927, 267]. Protocols [293, 453, 886, 494]. Provable [498]. Prover [289]. Proving [297, 165, 636]. Proxies [696]. Proximity [940, 541]. Proximity-Graph [541]. Pruning [335]. Pseudo [83, 117]. Pseudo-minimization [83]. Pseudo-Relevance [117]. PSO [774, 783]. Psychological [48]. Public [676, 965, 455, 454, 888]. Public-Key [455, 454, 888]. PUFs [877]. Pull [462]. Pump [479]. Push [462]. Push-Pull [462]. Pushdown [626, 728, 729, 853, 854, 739, 741, 742]. Putting [646]. Puzzles [454]. PVSS [456].

QoS [274]. Quad [189]. Quad-tree [189]. Quadrant [801]. Quadratic [825, 339]. Quality [555]. Quantifications [909]. Quantifier [290]. Quantifier-Free [290]. Quantifiers [846, 909]. Quantifying [185]. Quantitative [56, 46]. Quantization [340]. Quantum [818, 226]. Quasi [628]. Quasi-Eulerian [628]. Queries [89]. Query [291, 122, 88, 119, 128, 246, 127]. Querying [10]. Questions [281]. Queue [707]. Queues [66]. Quicksort [55].

R [1005]. Radar [786]. Radial [336, 359]. Radii [137]. Random [816, 685, 64, 103, 457, 188, 536, 458]. Randomized [842]. Randomizing [436]. Randomness [733]. Range [947, 65, 681, 226]. Range-Vibrotactile [681]. Ranging [660]. Rank [828, 635, 342]. Ranking [560]. Rat [545]. Rate [114]. Rate-Limited [114]. Ratify [54]. Ratio [825]. Rational [63]. Rationals [388]. Raw [47]. ray [229, 222, 226, 223]. Rayleigh [800]. RBAC [422]. RDFS [291]. Reachability [851]. Reactions [578]. Reactive [201, 196]. Read [59, 476]. Read-Once [59]. Reading [669, 48]. Real [569, 775, 923, 553, 574, 669, 274, 810]. Real-Time [569, 553, 574, 669, 810]. Real-Valued [775]. Realistic [588]. Realization [463]. Realizing [471]. Reasoner [51]. Reasoning [210, 34, 546, 202, 767, 999, 554, 280, 314, 798]. Recognition [566, 519, 1011, 524, 665, 576, 612, 370, 361]. Recognizing [39]. Recommendation [746, 810]. Recommendations [589]. Recommender [592, 596, 606]. Recommenders [599, 605]. Reconfiguration [554, 11]. Reconstruction [915]. Reconstructions [582]. Recorded [657]. Records [317]. Recovering [753]. Recursive [627, 292]. Reduced [447, 342, 386, 490, 491]. Reduced-Round [491]. Reducible [377]. Reducing [148, 883]. Reduction [465, 760, 528, 798]. Reductions [953]. Redundant [564, 363]. Reference [904, 3, 248, 970]. Reference-Set [904]. Referrer [697]. Referrer-Anonymizing [697]. Refined [964]. Reflectance [347]. Reflections [207]. Regarding [598]. Regime [291]. Region [262, 329, 346]. Regions [185]. Register [464]. Registration [186, 184, 175, 994, 173, 177, 190, 178, 191, 187, 582, 222, 188, 182, 192, 189, 179, 180, 176, 185]. Registry [957]. Regression [174, 762, 534, 347]. Regular [849, 622, 629, 855, 66, 735, 638]. Regularity [630]. Regularization [184]. Regularizer [787]. Regulation [75]. Regulator [462]. Reinforcer [332]. Related [675, 414, 41]. Relation [592, 344, 457]. Relation-Based [344]. Relational [423, 46, 525, 972, 340]. Relations [780, 119]. Relationship [420]. Relationship-Based [420]. Relationships [331]. Release [454]. Relevance [117]. Relevant [552, 128]. Relief [363]. Remote [687, 674, 677, 427]. Removal [171]. Removing [729]. Renaming [112]. Rendering [584, 617]. Representability [731]. Representation [987, 208, 554, 602, 981]. Representations [41, 575, 33]. Representing [40]. Reproducibility [583]. Requirements [670, 5, 759, 598, 9]. Resampling [574, 875]. Research [966, 546, 36, 965, 547, 650, 343, 963, 799, 788]. Resettable [823]. Resilience [821, 455]. Resilient [55, 441, 271, 945]. Resistance [498]. Resistant [427]. Resolution [229, 776, 882, 281, 188]. Resolving [915]. Resonator [475]. Resource [925, 159, 631, 694, 247]. Resource-Payment [694]. Respect [29]. Response [813, 223]. Responses [332]. Restricted [283, 383]. Result [754]. Resulting [741]. Results [80, 856, 275, 370]. Resumption [901]. Retrieval [513, 88, 126, 117, 991, 125]. Reusable [745]. Reuse [750]. Revealing [433]. Reversal [23]. Reverse [640, 94]. Revised [1033, 990, 1031, 1032, 1034, 1030]. Revisited [944, 302]. Revolution [1005]. Rewriting [853, 298]. Reykjavik [990]. RFID [276]. RGB [566]. RGB-Depth [566]. Rides [49]. rigid [175]. Rigidity [192]. Ring [379]. Rings [382]. Risk [698, 221, 883]. RKHS [787]. RLC [465]. Roadmap [666]. Robobraille [653]. Robot [505, 507, 511, 514, 515, 516, 520, 522, 785]. Robot-Human [522]. Robots [503, 508, 510, 517]. Robust [851, 65, 173, 368, 875, 188, 324, 189]. Rodent [517]. Rodent-Like [517].

Role [753, 421, 752]. Role-Based [753, 421, 752]. Root [390]. Rooted [629]. Rosser [855]. Rotating [262]. Rotation [84]. Rough [808, 779, 778]. Round [487, 491]. Rounding [829, 832]. Rounds [112]. Route [671]. Routes [134]. Routing [265, 261, 42]. RSA [448, 495, 456]. RSA-Based [456]. Rule [348, 531, 939]. Rule-Based [939]. Rules [562, 563, 962]. Run [85]. Run-Lengths [85]. Russia [988].

s [494]. S2N [721]. S4 [24]. S4-Modalities [24]. Saccadic [510]. Safety [778]. Sampling [827, 184, 562, 474, 188]. Sanction [863]. Sanctions [868]. Sanitizable [484]. Sanitizers [484]. SAT [283, 281]. Satallax [289]. Satisfaction [702, 411, 410, 761]. Satisfiability [279, 68]. Saturation [854]. Scalable [166]. Scale [239, 584, 234, 367]. Scaling [84, 748]. Scan [467]. Scandinavian [992]. Scans [882, 219]. Scatter [776]. Scene [367]. Scenes [566]. Scheduling [894, 268]. Scheme [269, 379, 931, 712, 133, 452, 937, 426, 427]. Schemes [483, 887, 456, 458, 945]. Schoolchildren [682]. Science [1028, 414, 767, 1015, 988, 1034, 963]. Sciences [546]. Scientific [195, 923, 305]. Scope [280]. Screen [670]. SCRIBE [653]. Scripting [435]. Search [693, 826, 622, 837, 91, 776, 809, 800, 548, 119, 335, 755, 125, 777]. Searches [888]. Searching [122, 127]. Secret [430]. Secretary [822]. Secure [690, 938, 946, 967, 451, 887, 444, 452, 487, 271, 427, 945]. Security [753, 483, 499, 6, 1033, 944, 936, 1006, 430, 1029, 293, 497, 445, 1032, 876, 426, 437, 877, 1007, 434, 419]. Seen [903]. Segmentation [221, 48, 346]. Segments [134]. Select [184]. Selected [1023, 1033, 990, 1031, 1032, 1034, 1030]. Selecting [533]. Selection [55, 362, 528, 876, 227, 795, 779, 740, 603, 364, 942, 357, 777]. Selenium [217]. Self [970, 478, 959, 114, 381, 777]. Self-adaptive [777]. Self-dual [381]. Self-manage [959]. Self-reference [970]. Self-stabilizing [114]. Semantic [750, 240, 432, 243, 757, 602, 119]. Semantically [317]. Semantics [285, 974, 898, 861, 976, 863, 197, 905, 916, 313, 27]. Semi [560, 542, 348, 276, 197, 543]. Semi-negative [197]. Semi-passive [276]. Semi-supervised [560, 348, 543]. Semiantichains [58]. Sense [507, 314]. Sensitivity [226]. Sensor [269, 270, 588, 506, 146, 65, 104, 928, 275, 272]. Sensor-Based [588]. Sentences [208]. Sentic [314]. Separation [167]. September [1029]. Sequence [86, 91, 33]. Sequences [85, 517, 551, 380, 567, 789]. Sequential [870, 781]. Serbia [998]. Sergot [995, 195, 194]. Series [971, 46, 174, 549]. Service [306, 671, 1000, 308, 927, 7, 309, 9, 11]. Service-Based [11]. Service-Oriented [306, 309]. Services [925, 6, 2, 679, 5, 243, 889, 697, 310, 10]. Session [646, 840]. Set [835, 864, 904, 779, 78, 563, 401, 778]. Sets [731, 849, 754, 808, 380]. Setting [919]. Several [484]. SHA [499, 501]. SHA-3 [499, 501]. Shapes [642]. Sharing [750, 430, 967, 674, 950]. Sheng [621]. Shenyang [1002, 1003, 1001]. Shibpur [1008]. Shift [496, 368, 464]. Shift-Type [496]. Shifting [508].

Shopping [674]. Short [768, 884, 459]. Shortest [80, 142]. Shoulders [839]. Show [602]. Shows [345]. Side [685]. Sieve [390]. Sign [330, 889]. Sign-On [889]. Signage [665]. Signals [331]. Signature [483, 423, 379, 458, 459]. Signature-Based [423]. Signatures [484, 482, 457]. Signers [484]. Signs [669]. Similarities [118]. Similarity [183, 748, 357]. Similarity-Based [748]. Simple [93, 596, 174, 138]. Simpler [78]. Simplex [286]. Simplex-Based [286]. Simplices [147]. Simplification [616]. Simplifies [906]. Simplify [521]. Simulated [319, 593, 800]. Simulating [505]. Simulation [588, 235, 581, 597, 267, 611, 921]. Simultaneous [791, 463, 179]. Simultaneously [823, 62]. Sina [594]. Single [141, 889, 608]. Single-Exponential [141]. SINR [106, 268, 105]. SIROCCO [990]. Site [435]. Size [228, 47, 81]. Skein [499]. Skeletal [576]. Sketches [940]. Skill [518]. SLAs [922]. Slovenia [1023, 1030]. Small [143, 378, 323, 448, 234]. Small-Bias [378]. Small-Scale [234]. Small-World [323]. Smart [676, 712, 951, 426, 427]. Smartcard [428]. Smartphone [960, 436]. Smartphones [663]. SMT [720, 281]. SMTInterpol [720]. Snapshot [115]. Snapshots [113]. Snyder [615]. Social [587, 420, 263, 756, 934]. Soft [506, 808, 321]. Soft-Body [506]. Software [235, 709, 1020, 384, 643, 961, 761, 718, 704, 1000]. SOHAC [548]. SOI [474]. Solution [456, 841, 435]. Solutions [9]. Solver [284, 720]. Solvers [154]. Solving [791, 286, 388, 591, 790, 443, 400, 345]. Some [732, 734, 1, 207, 392, 152, 910]. Sounds [369]. Source [719]. SP [446]. Space [609, 96, 88, 335, 353]. Space-and-Query-Time [88]. Spacer [466]. Spaces [733, 378, 564, 748]. Spain [1019, 1031, 1013, 1010]. Spam [941, 120]. Spamming [963]. Spanners [831]. Spanning [65, 561]. SPARQL [291]. Sparse [831, 710]. Sparsity [365]. Spatial [229, 811, 175, 568, 517, 185]. Spatial-Based [811]. Special [651, 646, 682, 1017, 1018]. Specialized [895]. Specific [225]. Specification [56, 771, 252, 962]. Specifications [71]. Spectral [361, 337]. Spectrometry [793]. Spectrometry-Based [793]. Spectrum [382, 364, 678]. Speech [679, 642, 682]. Speech-Transcription [679]. Speed [479, 472, 470]. Speeding [92, 255]. Speedith [51]. Spider [51]. Spiking [351]. SPIN [1020, 713, 721]. Split [140]. SPN [498]. SPN-Hash [498]. Sponsored [837]. Spontaneous [320, 41]. Spying [689]. Square [735, 796, 390]. Squares [77, 133, 686, 555]. SRAM [476]. SSReflect [249]. Stability [106, 813, 796]. Stabilization [512]. stabilizing [114]. Stable [113]. Stack [468]. Stacked [468]. Stage [793]. Stakes [957]. Standard [452]. Standards [922]. Standards-Based [922]. Standing [839]. Star [732, 735]. Star-Free [732]. State [974, 463, 735, 736, 349, 640, 254, 743, 351]. Stateless [741, 455]. Stateless/ Stateful [455]. Statement [241]. States [57]. Static [733, 512, 476]. Statistical [45, 878, 575, 686, 436]. Statistics [695, 580]. Status [332]. Stealing [965]. Stegun [247]. Still [281]. Storage [548, 949]. Store [739]. Strategies [818, 508, 769, 179]. Strategy [508, 512, 540]. Stratified [124]. Stream [450, 948, 975, 488]. Streaming [844, 242]. Streams [691, 82, 755].

Strength [169]. Stretch [96]. Strict [630]. String [81, 87, 633, 94]. Strip [99]. Strong [290, 632, 104, 459]. Strongly [65]. Structural [974, 990]. Structure [538, 369, 528, 582]. STS [493]. Student [38, 646, 595, 655]. Students [656, 643, 650, 673, 645, 678]. Studies [412, 598, 399]. Study [964, 510, 594, 769, 331, 784, 643, 961, 779, 664, 516, 265, 575, 277, 543, 127, 329, 434, 361]. Style [7, 699, 610]. Sub [86, 476]. Sub-sums [86]. Sub-threshold [476]. Subdirectional [26]. Subexponential [407]. Subexponential-Time [407]. Subgraph [431]. Subgroups [820]. Subject [339]. Submodular [819]. Subregular [740]. Subsequence [85]. Subset [835, 490]. Subspace [539]. Substitutions [73]. Successful [126, 280]. Sudden [809]. Sudden-Target [809]. Sufficient [441]. Suffix [633]. Suite [123]. Sum [137, 485]. sums [86]. Superior [325]. Supervised [542, 560, 348, 543]. Support [676, 647, 263, 316, 643, 305, 795, 527, 539, 541, 678]. Supporting [924, 764, 682, 645]. Supports [677, 548]. Surface [191, 575, 182]. Surface-Based [182]. Surfaces [132]. Surjective [69]. Survey [410]. Survivable [791]. Suspects [415]. Suspension [511]. Suspicious [605]. Sustainable [920]. SVM [787, 348]. SVR [792]. Swarm [784, 574, 803]. SWAT [992]. Sweep [707]. Sweep-Line [707]. Swiftly [518]. Swimming [503]. Switching [796, 321]. Symbolic [203, 703, 371, 168]. Symbols [737]. Symmetric [382, 181]. Symmetrized [360]. Symposium [38, 1019, 992, 988, 997, 989, 1008, 1002, 1003]. Synchronization [843, 63, 634]. Synchronize [637]. Synchronizing [628, 635]. Synchronous [112, 282]. SYND [488]. Syntactic [732]. Synthesis [159, 160, 704, 152, 610]. Synthesizing [154]. System [985, 284, 807, 642, 775, 473, 256, 614, 682, 514, 577, 372, 677, 254, 470, 219, 669, 667, 778, 943]. Systematic [886, 758]. Systems [424, 728, 854, 113, 931, 293, 294, 592, 709, 388, 5, 1, 904, 926, 596, 3, 771, 997, 1021, 598, 114, 309, 894, 213, 742, 520, 726, 1010, 606, 782, 972, 859, 798, 1001].

Table [807, 48]. Tag [749, 276]. Tagging [750, 748]. Tags [119]. Takers [658]. Taking [279]. Tallinn [1034]. Taming [294]. tamper [427]. Tangent [353]. Target [809, 519, 537]. Task [328, 319, 513, 41, 225]. Task-Related [41]. Task-Specific [225]. Tasks [764, 603]. Taxon [508]. TC [1029]. TCP [689]. TEA [489]. Teacher [910]. Teaching [643, 310]. Teams [761]. Technique [467, 561]. Techniques [258, 280, 436, 567, 534]. Technologies [1019, 673]. Technology [36, 1024, 650, 305, 664, 310]. Tectogrammar [906]. Teleo [196]. Teleo-Reactive [196]. Temperature [275]. Template [324, 884]. Temporal [780, 864, 757, 203, 515, 199, 33]. Temporal-Epistemic [203]. Temporally [183]. Temporally-Dependent [183]. Ten [767]. Tensor [190, 599]. Term [884]. Termination [852, 164, 165, 163, 170, 297]. Terms [127, 493]. Ternary [473]. Terrorist [893]. Test [675, 820, 471, 658, 1008, 718]. Test-Takers [658]. Testability [630]. Testbed [273, 274]. Testing [830, 658, 467, 533, 940, 703, 762, 468].

Tests [59]. Text [698, 779, 41]. Texts [92]. Texture [205, 360]. Texturing [581]. Their [973, 730, 301, 638, 380, 30]. Thematic [646]. Theme [589]. Themes [546]. Theorem [636, 26]. Theorem-Proving [636]. Theoretic [442]. Theoretical [1012, 49]. Theories [56, 14, 293, 1025]. Theory [424, 850, 250, 838, 992, 376, 988, 923, 613, 841, 54, 798, 781]. Thermal [788]. Thickness [217]. Things [273]. Thought [40]. Threat [432]. Three [235, 784, 490]. Three-Layer [784]. Three-Subset [490]. Threshold [379, 478, 476]. Thresholds [68]. Tick [548]. Tight [410]. Time [143, 820, 569, 852, 81, 46, 407, 86, 122, 856, 142, 70, 88, 174, 349, 133, 553, 339, 574, 549, 372, 669, 797, 351, 910, 810, 345]. Time-Frequency [372]. Time-Series [46, 174]. Time-Varying [339, 797, 345]. Timed [848, 851]. Tissue [229, 234]. TN [994]. Toggle [480]. Tolerance [948]. Tolerant [264, 831, 97, 695]. Tomosynthesis [220, 216, 217, 226, 219, 231, 225]. Tool [964, 807, 588, 652, 923, 660, 715]. Toolkit [649]. Tools [244, 518, 704]. Top [88, 532]. Top- [88, 532]. Topic [746, 799]. Topic-Oriented [799]. Topological [71]. Topologies [808, 268]. Topology [305]. Tor [689, 942]. Touch [670]. Tough [415]. Tour [589]. Tower [846]. Tower-Type [846]. Trace [976, 171]. Tracing [980]. Tracking [574, 368, 573, 324]. Tractability [408, 396]. Tractable [835, 458]. Tradeoff [880]. Trading [939]. Traffic [666, 593, 689, 687, 688]. training [525]. Transaction [8]. Transactions [1015]. Transcription [679]. Transducer [626]. Transducers [844]. Transductions [904]. Transductive [525]. Transfer [442, 953, 111]. Transform [94]. Transformation [721, 222, 371]. Transformations [760, 758]. Transforms [331]. Transient [361]. Transition [254, 972]. Transitions [979, 610]. Translating [672]. Translation [469, 49]. Transmission [487]. Transport [676]. Treatment [332, 781]. Tree [844, 46, 77, 911, 904, 639, 892, 268, 561, 189]. Tree-Inspired [561]. Trees [843, 849, 826, 629, 76, 536]. Treewidth [408, 856]. Trends [207]. Triangle [245]. Tripartite [441]. Trusted [596]. Trustworthiness [593]. Trustworthy [7]. Truthful [288]. TSP [801]. TTS [493]. Tumor [806]. Tumour [216]. Tunable [509, 511]. Turbines [550]. Turbo [464]. TV [602]. TV-Show [602]. Twinder [755]. Twitter [594, 756, 755, 746]. Two [899, 946, 84, 456, 207, 793, 805]. Two-Dimensional [84, 805]. Two-Party [946]. Two-Stage [793]. Type [496, 846, 915, 677, 433, 813]. Typed [625]. Types [627, 159, 840, 25, 30].

UAS [319]. Ubiquitous [506]. UDGs [65]. UEL [284]. UK [987, 1026, 1027, 1029, 1020, 999]. Ultra [476]. Ultra-Low [476]. Ultrasound [186]. UMAP [1014]. Unambiguous [737]. Unary [736]. Unbounded [157]. Uncapacitated [829]. Uncertain [796]. Uncertainty [178, 185]. Unconscious [578]. Undeniable [457]. Underlap [466]. Understanding [956, 568, 244, 245, 686]. Underwater [507, 277]. Unicast [791]. Unichain [58]. Unification [282, 283, 284]. Unified [422]. Uniform [728, 425]. Unimodal [227]. Uninterrupted [262]. Union [735]. Unions [145]. Unique [100, 133]. Unit [133, 914, 601, 667, 842]. Units [305]. Universal [733]. Unlinkable [949]. Unplugged [414]. Unsatisfiability [816]. Unsupervised [535]. Updates [587]. Updating [324]. UPPAAL [717]. USA [994, 993, 996]. Usability [423, 680]. Usage [631, 930]. Use [34, 707, 41, 699, 518]. USEC [1033]. Used [575]. Useful [127]. User [588, 590, 420, 670, 566, 594, 680, 597, 598, 600, 426, 427, 810, 896, 962, 1014]. User-Based [896]. User-to-User [420]. Users [486, 959, 593, 764, 606, 669]. Using [450, 675, 328, 569, 228, 714, 463, 430, 878, 709, 965, 165, 177, 301, 687, 384, 221, 331, 614, 574, 785, 227, 305, 244, 515, 516, 531, 224, 613, 563, 426, 233, 324, 550, 535, 267, 433, 132, 335, 272, 657, 357, 537, 715, 353, 347, 356, 814]. USPTO [127].

Validation [469, 593]. Value [614]. Valued [285, 775, 782]. Values [850]. Variable [225]. Variable-Exposure [225]. Variables [381]. Variant [126, 452]. Variants [493]. Variation [232]. Variations [386]. Varying [339, 797, 345]. Vault [486]. VCC [157]. VDAT [1008]. Vector [61, 384, 251, 795, 527, 541, 480, 176, 710, 373, 340]. Vectorial [383, 564]. Vectors [539]. Vedic [470]. Vehicles [676, 322, 321]. Vehicular [593, 266]. Venn [49]. Verbal [571, 583]. Verifiable [457]. Verification [155, 383, 709, 712, 771, 482, 252, 993, 157, 156, 718, 704, 708, 292]. Verifier [459]. Verifying [251]. versus [824, 512, 742]. Vertex [552, 835, 139, 69, 401]. Vertex-Surjective [69]. Vertically [429]. Very [229]. via [424, 469, 66, 904, 976, 213, 527, 176]. Vibrotactile [681]. Video [652, 698, 567]. Vienna [991]. View [521, 574, 573]. View-Based [521]. Vigo [1019]. Violation [715]. Virtual [570, 609, 572, 571, 315, 304, 617]. Virtualized [11]. virus [428]. Visibility [609, 834, 131]. Vision [37, 664]. Visitor [588]. Visitors [673]. Visual [609, 642, 325, 809, 668, 678]. Visualization [229, 613]. Visualizing [859, 40, 185]. Visually [676, 647, 669, 667]. Visuo [512]. Visuo-motor [512]. VLSI [1008, 325, 467]. VM [926]. Voice [680]. Voltage [478, 462]. Volume [228, 235]. Volumetric [232]. Voronoi [136]. Voting [412, 943]. vs [787, 698, 179]. Vulnerabilities [426].

w.r.t [283]. WAIFI [1004]. Walk [610]. Walking [503]. Wallets [936]. Walsh [382]. Warwick [1026, 1027]. Washington [943]. Watchman [134]. Watermarking [891, 885]. Watermarking-Encryption [885]. Waveform [786]. Wavelet [76]. Waves [800]. WBIR [994]. Weak [627, 916]. Weakest [867]. Weaknesses [956]. Wearable [681]. Weaving [402]. Web [753, 752, 750, 1022, 935, 763, 122, 764, 124, 759, 745, 663, 760, 688, 779, 761, 121, 747, 120, 11, 402, 963, 799]. Web-Based [935]. WebML [758]. Website [687]. Websites [751]. WECSR [1033]. Weibo [594]. Weight [377]. Weighted [828, 833, 360, 632, 624, 558, 789]. Weigthed [625]. Weimar [451]. Weimar-DM [451]. WG [1006]. Wheeler [94]. Who [643]. Wide [475]. Widgets [745]. Wind [550]. Wired [928]. Wireless [106, 270, 959, 103, 259, 998, 673, 275, 277, 105]. Wiring [44]. Wise [748]. without [107, 143]. WLAN [676]. Wollongong [1007]. Word [81, 632, 296, 746, 955]. Word-Size [81]. Words [731, 853, 636]. Workbench [303]. Working [315]. Workload [477, 591]. Workshop [35, 36, 994, 1020, 1031, 997, 1021, 996, 1004, 1034]. Workshops [1033, 992]. World [861, 323]. Worlds [27]. Worst [66]. Worst-Case [66]. Write [248]. Writing [699]. WSNs [261, 271].

X [229, 222, 226, 223]. X-ray [229, 222, 226, 223]. XML [627]. XOR [818]. XORSAT [68]. XSS [435]. XSS-Dec [435]. XTEA [489, 445, 490]. XVI [1015]. XXVI [1006].

Years [397]. Yield [937]. Yu [621]. YZ [494].

Z3 [279]. Zero [823, 453]. Zero-Knowledge [453]. Zhang [345]. ZNN [339, 345].

References

Friesen:2012:SCB

[1] Andreas Friesen, Wolfgang Theilmann, Markus Heller, Jens Lemcke, and Christof Momm. On some challenges in business systems management and engineering for the networked enterprise of the future. Lecture Notes in Computer Science, 7350:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

      chapter/10.1007/978-3-642-32439-0_1/.

Badr:2012:DEB

[2] Youakim Badr, Yong Peng, and Frédérique Biennier. Digital ecosystems for business e-services in knowledge-intensive firms. Lecture Notes in Computer Science, 7350:16–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

      chapter/10.1007/978-3-642-32439-0_2/.

Hofreiter:2012:IOR

[3] Birgit Hofreiter, Christian Huemer, Gerti Kappel, Dieter Mayrhofer, and Jan vom Brocke. Inter-organizational reference models — may interorganizational systems profit from reference modeling? Lecture Notes in Computer Science, 7350:32–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

      chapter/10.1007/978-3-642-32439-0_3/.

Cardellini:2012:GGD

[4] Valeria Cardellini, Emiliano Casalicchio, Vincenzo Grassi, and Francesco Lo Presti. On guaranteeing global dependability properties in collaborative business process management. Lecture Notes in Computer Science, 7350:48–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32439-0_4/.

Dubois:2012:CAA

[5] Eric Dubois, Sylvain Kubicki, Sophie Ramel, and André Rifaut. Capturing and aligning assurance requirements for business services systems. Lecture Notes in Computer Science, 7350:71–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32439-0_5/.

Anisetti:2012:CLS

[6] Marco Anisetti, Claudio A. Ardagna, and Ernesto Damiani. Containerlevel security certification of services. Lecture Notes in Computer Science, 7350:93–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32439-0_6/.

Maciaszek:2012:AST

[7] Leszek A. Maciaszek. An architectural style for trustworthy adaptive service based applications. Lecture Notes in Computer Science, 7350:109– 121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32439-0_7/.

Parkin:2012:CAB

[8] Michael Parkin and Mike P. Papazoglou. A conceptual architecture for business-aware transaction management. Lecture Notes in Computer Science, 7350:122–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32439-0_8/.

Niemoller:2012:CHS

[9] Jörg Niemöller, Eugen Freiter, Konstantinos Vandikas, Raphaël Quinet, and Roman Levenshteyn. Composition in heterogeneous service networks: Requirements and solutions. Lecture Notes in Computer Science, 7350:138–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32439-0_9/.

Smith:2012:OBQ

[10] Fabrizio Smith, Michele Missikoff, and Maurizio Proietti. Ontologybased querying of composite services. Lecture Notes in Computer Science, 7350:159–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32439-0_10/.

Rodriguez:2012:GGB

[11] Ismael Bouassida Rodriguez, Riadh Ben Halima, Khalil Drira, and Christophe Chassot. A graph grammar-based dynamic reconfiguration for virtualized Web service-based composite architectures. Lecture Notes in Computer Science, 7350:181–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32439-0_11/.

Anonymous:2012:BMa

[12] Anonymous. Back matter. Lecture Notes in Computer Science, 7350: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32439-0/1.

Anonymous:2012:FMa

[13] Anonymous. Front matter. Lecture Notes in Computer Science, 7350: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32439-0/1.

Clark:2012:LGL

[14] Alexander Clark. Logical grammars, logical theories. Lecture Notes in Computer Science, 7351:1–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31262-5_1/.

Fouquere:2012:LNL

[15] Christophe Fouqueré and Myriam Quatrini. Ludics and natural language: First approaches. Lecture Notes in Computer Science, 7351:21– 44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31262-5_2/.

Asher:2012:NCB

[16] Nicholas Asher. The non cooperative basis of implicatures. Lecture Notes in Computer Science, 7351:45–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31262-5_3/.

Graf:2012:MGM

[17] Thomas Graf. Movement-generalized minimalist grammars. Lecture Notes in Computer Science, 7351:58–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31262-5_4/.

Ishishita:2012:TFP

[18] Yuri Ishishita and Daisuke Bekki. Toward the formulation of presupposition by illative combinatory logic. Lecture Notes in Computer Science, 7351:74–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31262-5_5/.

Karlov:2012:AAN

[19] Boris Karlov. Abstract automata and a normal form for categorial dependency grammars. Lecture Notes in Computer Science, 7351:86–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31262-5_6/.

Kobele:2012:IMD

[20] Gregory M. Kobele. Importing Montagovian dynamics into minimalism. Lecture Notes in Computer Science, 7351:103–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31262-5_7/.

Kobele:2012:CA

[21] Gregory M. Kobele and Jens Michaelis. CoTAGs and ACGs. Lecture Notes in Computer Science, 7351:119–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31262-5_8/.

Kubota:2012:GLC

[22] Yusuke Kubota and Robert Levine. Gapping as like-category coordination. Lecture Notes in Computer Science, 7351:135–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31262-5_9/.

Kuznetsov:2012:CLC

[23] Stepan Kuznetsov. L-completeness of the Lambek calculus with the reversal operation. Lecture Notes in Computer Science, 7351:151–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31262-5_10/.

Lin:2012:DFN

[24] Zhe Lin. Distributive full nonassociative Lambek calculus with S4modalities is context-free. Lecture Notes in Computer Science, 7351:161– 172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31262-5_11/.

Luo:2012:CNT

[25] Zhaohui Luo. Common nouns as types. Lecture Notes in Computer Science, 7351:173–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31262-5_12/.

Ozaki:2012:EDT

[26] Hiroko Ozaki and Daisuke Bekki. Extractability as the deduction theorem in subdirectional combinatory logic. Lecture Notes in Computer Science, 7351:186–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31262-5_13/.

Plummer:2012:APW

[27] Andrew Plummer and Carl Pollard. Agnostic possible worlds semantics. Lecture Notes in Computer Science, 7351:201–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31262-5_14/.

Ranalter:2012:AMA

[28] Kurt Ranalter. Abstract machines for argumentation. Lecture Notes in Computer Science, 7351:213–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31262-5_15/.

Sorokin:2012:CLC

[29] Alexey Sorokin. On the completeness of Lambek calculus with respect to cofinite language models. Lecture Notes in Computer Science, 7351:229– 233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31262-5_16/.

Xue:2012:DTT

[30] Tao Xue and Zhaohui Luo. Dot-types and their implementation. Lecture Notes in Computer Science, 7351:234–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31262-5_17/.

Anonymous:2012:BMb

[31] Anonymous. Back matter. Lecture Notes in Computer Science, 7351: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31262-5/1.

Anonymous:2012:FMb

[32] Anonymous. Front matter. Lecture Notes in Computer Science, 7351: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31262-5/1.

Plaisant:2012:LLI

[33] Catherine Plaisant. Life on the line: Interacting with temporal event sequence representations. Lecture Notes in Computer Science, 7352:1– 2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-31223-6_1.

Barker-Plummer:2012:LUO

[34] Dave Barker-Plummer, John Etchemendy, Michael Murray, Emma Pease, and Nik Swoboda. Learning to use the openbox: a framework for the implementation of heterogeneous reasoning. Lecture Notes in Computer Science, 7352:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-31223-6_2.

Chapman:2012:IWE

[35] Peter Chapman and Luana Micallef. 3rd international workshop on Euler diagrams. Lecture Notes in Computer Science, 7352:4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-31223-6_3.

Cox:2012:TED

[36] Richard Cox and Jonathan San Diego. Technology enhanced diagrams research workshop. Lecture Notes in Computer Science, 7352: 5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-31223-6_4.

Goncu:2012:AGG

[37] Cagatay Goncu and Kim Marriott. Accessible graphics: Graphics for vision impaired people. Lecture Notes in Computer Science, 7352:6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-31223-6_5.

Best:2012:GSS

[38] Lisa A. Best. Graduate student symposium of diagrams 2012. Lecture Notes in Computer Science, 7352:7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-31223-6_6.

Burns:2012:ARI

[39] Richard Burns, Sandra Carberry, Stephanie Elzer, and Daniel Chester. Automatically recognizing intended messages in grouped bar charts. Lecture Notes in Computer Science, 7352:8–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31223-6_7/.

Tversky:2012:RCC

[40] Barbara Tversky, James E. Corter, Lixiu Yu, David L. Mason, and Jeffrey V. Nickerson. Representing category and continuum: Visualizing thought. Lecture Notes in Computer Science, 7352:23–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31223-6_8/.

Manalo:2012:EMS

[41] Emmanuel Manalo and Yuri Uesaka. Elucidating the mechanism of spontaneous diagram use in explanations: How cognitive processing of text and diagrammatic representations are influenced by individual and task-related factors. Lecture Notes in Computer Science, 7352:35–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31223-6_9/.

Wybrow:2012:OHR

[42] Michael Wybrow, Kim Marriott, and Peter J. Stuckey. Orthogonal hyperedge routing. Lecture Notes in Computer Science, 7352:51–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31223-6_10/.

Klauske:2012:ILD

[43] Lars Kristian Klauske, Christoph Daniel Schulze, Miro Spönemann, and Reinhard von Hanxleden. Improved layout for data flow diagrams with port constraints. Lecture Notes in Computer Science, 7352:65– 79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31223-6_11/.

Ernstbrunner:2012:ALW

[44] Christian Ernstbrunner and Josef Pichler. Aesthetic layout of wiring diagrams. Lecture Notes in Computer Science, 7352:80–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31223-6_12/.

Acarturk:2012:PLA

[45] Cengiz Acartürk. Points, lines and arrows in statistical graphs. Lecture Notes in Computer Science, 7352:95–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31223-6_13/.

Burch:2012:EIP

[46] Michael Burch, Michael Raschke, Miriam Greis, and Daniel Weiskopf. Enriching indented pixel tree plots with node-oriented quantitative, categorical, relational, and time-series data. Lecture Notes in Computer Science, 7352:102–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31223-6_14/.

Bradley:2012:IES

[47] Michael T. Bradley, Andrew Brand, and A. Luke MacNeill. Interpreting effect size estimates through graphic analysis of raw data distributions. Lecture Notes in Computer Science, 7352:117–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31223-6_15/.

Sugio:2012:PEM

[48] Takeshi Sugio, Atsushi Shimojima, and Yasuhiro Katagiri. Psychological evidence of mental segmentation in table reading. Lecture Notes in Computer Science, 7352:124–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31223-6_16/.

Takemura:2012:PTI

[49] Ryo Takemura. Proof-theoretical investigation of Venn diagrams: a logic translation and free rides. Lecture Notes in Computer Science, 7352:132– 147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31223-6_17/.

Bottoni:2012:EDE

[50] Paolo Bottoni, Gennaro Costagliola, and Andrew Fish. Euler diagram encodings. Lecture Notes in Computer Science, 7352:148–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31223-6_18/.

Urbas:2012:SDR

[51] Matej Urbas, Mateja Jamnik, Gem Stapleton, and Jean Flower. Speedith: a diagrammatic reasoner for spider diagrams. Lecture Notes in Computer Science, 7352:163–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31223-6_19/.

Cheng:2012:ADH

[52] Peter C.-H. Cheng. Algebra diagrams: a HANDi introduction. Lecture Notes in Computer Science, 7352:178–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31223-6_20/.

Anonymous:2012:FMc

[53] Anonymous. Front matter. Lecture Notes in Computer Science, 7352: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31223-6/1.

Vazirani:2012:CTA

[54] Vijay V. Vazirani. Can the theory of algorithms ratify the “invisible hand of the market”? Lecture Notes in Computer Science, 7353:1– 5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30642-6_1/.

Babenko:2012:RQS

[55] Maxim Babenko and Ivan Pouzyrevsky. Resilient Quicksort and selection. Lecture Notes in Computer Science, 7353:6–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30642-6_2/.

Bauer:2012:GQS

[56] Sebastian S. Bauer, Uli Fahrenberg, Axel Legay, and Claus Thrane. General quantitative specification theories with modalities. Lecture Notes in Computer Science, 7353:18–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30642-6_3/.

Blondin:2012:CIF

[57] Michael Blondin and Pierre McKenzie. The complexity of intersecting finite automata having few final states. Lecture Notes in Computer Science, 7353:31–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30642-6_4/.

Bosek:2012:NAS

[58] Bartłomiej Bosek, Stefan Felsner, Kolja Knauer, and Grzegorz Matecki. News about semiantichains and unichain coverings. Lecture Notes in Computer Science, 7353:43–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30642-6_5/.

Chistikov:2012:CTR

[59] Dmitry V. Chistikov. Checking tests for read-once functions over arbitrary bases. Lecture Notes in Computer Science, 7353:52–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30642-6_6/.

Cohen:2012:AMP

[60] Nachshon Cohen and Zeev Nutov. Approximating minimum power edge-multi-covers. Lecture Notes in Computer Science, 7353:64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30642-6_7/.

Demenkov:2012:LBC

[61] Evgeny Demenkov. A lower bound on circuit complexity of vector function in U_{2}. Lecture Notes in Computer Science, 7353:76–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30642-6_8/.

Demenkov:2012:CAM

[62] Evgeny Demenkov, Alexander S. Kulikov, Ivan Mihajlin, and Hiroki Morizumi. Computing all MOD-functions simultaneously. Lecture Notes in Computer Science, 7353:81–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30642-6_9/.

Diekert:2012:BSD

[63] Volker Diekert and Manfred Kufleitner. Bounded synchronization delay in omega-rational expressions. Lecture Notes in Computer Science, 7353: 89–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30642-6_10/.

Dietzfelbinger:2012:TOD

[64] Martin Dietzfelbinger and Michael Rink. Towards optimal degreedistributions for left-perfect matchings in random bipartite graphs. Lecture Notes in Computer Science, 7353:99–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30642-6_11/.

Dobrev:2012:RSR

[65] Stefan Dobrev, Evangelos Kranakis, Oscar Morales Ponce, and Milan Plžík. Robust sensor range for constructing strongly connected spanning digraphs in UDGs. Lecture Notes in Computer Science, 7353:112– 124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30642-6_12/.

Elmasry:2012:WCO

[66] Amr Elmasry and Jyrki Katajainen. Worst-case optimal priority queues via extended regular counters. Lecture Notes in Computer Science, 7353:125–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30642-6_13/.

Fox-Epstein:2012:CMA

[67] Eli Fox-Epstein and Danny Krizanc. The complexity of minor-ancestral graph properties with forbidden pairs. Lecture Notes in Computer Science, 7353:138–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30642-6_14/.

Goerdt:2012:STB

[68] Andreas Goerdt and Lutz Falke. Satisfiability thresholds beyond k-XORSAT. Lecture Notes in Computer Science, 7353:148–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30642-6_15/.

Golovach:2012:FVS

[69] Petr A. Golovach, Bernard Lidický, Barnaby Martin, and Daniël Paulusma. Finding vertex-surjective graph homomorphisms. Lecture Notes in Computer Science, 7353:160–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30642-6_16/.

Heggernes:2012:BDB

[70] Pinar Heggernes and Sigve H. Sæther. Broadcast domination on block graphs in linear time. Lecture Notes in Computer Science, 7353:172– 183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30642-6_17/.

Heinemann:2012:CCT

[71] Bernhard Heinemann. Characterizing certain topological specifications. Lecture Notes in Computer Science, 7353:184–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30642-6_18/.

Jiraskova:2012:DCO

[72] Galina Jirásková. Descriptional complexity of operations on alternating and Boolean automata. Lecture Notes in Computer Science, 7353:196– 204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30642-6_19/.

Jolivet:2012:CMC

[73] Timo Jolivet and Jarkko Kari. Consistency of multidimensional combinatorial substitutions. Lecture Notes in Computer Science, 7353:205– 216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30642-6_20/.

Anonymous:2012:FMd

[74] Anonymous. Front matter. Lecture Notes in Computer Science, 7353: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30642-6/1.

Shamir:2012:GRP

[75] Ron Shamir. Gene regulation, protein networks and disease: a computational perspective. Lecture Notes in Computer Science, 7354:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-31265-6_1.

Navarro:2012:WTA

[76] Gonzalo Navarro. Wavelet trees for all. Lecture Notes in Computer Science, 7354:2–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31265-6_2/.

Crochemore:2012:MNS

[77] Maxime Crochemore, Costas S. Iliopoulos, Tomasz Kociumaka, and Marcin Kubica. The maximum number of squares in a tree. Lecture Notes in Computer Science, 7354:27–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31265-6_3/.

Ouangraoua:2012:FSM

[78] Aïda Ouangraoua and Mathieu Raffinot. Faster and simpler minimal conflicting set identification. Lecture Notes in Computer Science, 7354: 41–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31265-6_4/.

Bruckner:2012:PCC

[79] Sharon Bruckner, Falk Hüffner, Christian Komusiewicz, Rolf Niedermeier, and Sven Thiel. Partitioning into colorful components by minimum edge deletions. Lecture Notes in Computer Science, 7354:56–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31265-6_5/.

Blokh:2012:AAH

[80] Dima Blokh, Danny Segev, and Roded Sharan. Approximation algorithms and hardness results for shortest path based graph orientations. Lecture Notes in Computer Science, 7354:70–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31265-6_6/.

Breslauer:2012:CTW

[81] Dany Breslauer, Leszek Gasieniec, and Roberto Grossi. Constant-time word-size string matching. Lecture Notes in Computer Science, 7354:83– 96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31265-6_7/.

Clifford:2012:PMM

[82] Raphaël Clifford, Markus Jalsenius, Ely Porat, and Benjamin Sach. Pattern matching in multiple streams. Lecture Notes in Computer Science, 7354:97–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31265-6_8/.

AitMous:2012:ELP

[83] Omar AitMous, Frédérique Bassino, and Cyril Nicaud. An efficient linear pseudo-minimization algorithm for Aho–Corasick automata. Lecture Notes in Computer Science, 7354:110–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31265-6_9/.

Hundt:2012:ETD

[84] Christian Hundt and Florian Wendland. Efficient two-dimensional pattern matching with scaling and rotation and higher-order interpolation. Lecture Notes in Computer Science, 7354:124–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31265-6_10/.

Blin:2012:HLC

[85] Guillaume Blin, Laurent Bulteau, Minghui Jiang, Pedro J. Tejada, and Stéphane Vialette. Hardness of longest common subsequence for sequences with bounded run-lengths. Lecture Notes in Computer Science, 7354:138–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31265-6_11/.

Cicalese:2012:NLT

[86] Ferdinando Cicalese, Eduardo Laber, Oren Weimann, and Raphael Yuster. Near linear time construction of an approximate index for all maximum consecutive sub-sums of a sequence. Lecture Notes in Computer Science, 7354:149–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31265-6_12/.

Condon:2012:CSP

[87] Anne Condon, Ján Maňuch, and Chris Thachuk. The complexity of string partitioning. Lecture Notes in Computer Science, 7354:159–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31265-6_13/.

Hon:2012:TOS

[88] Wing-Kai Hon, Rahul Shah, and Sharma V. Thankachan. Towards an optimal space-and-query-time index for top- k document retrieval. Lecture Notes in Computer Science, 7354:173–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31265-6_14/.

Hon:2012:DLQ

[89] Wing-Kai Hon, Rahul Shah, Sharma V. Thankachan, and Jeffrey Scott Vitter. Document listing for queries with excluded pattern. Lecture Notes in Computer Science, 7354:185–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31265-6_15/.

Kucherov:2012:CDP

[90] Gregory Kucherov, Yakov Nekrich, and Tatiana Starikovskaya. Crossdocument pattern matching. Lecture Notes in Computer Science, 7354:196–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31265-6_16/.

Ferguson:2012:FFS

[91] Michael P. Ferguson. FEMTO: Fast search of large sequence collections. Lecture Notes in Computer Science, 7354:208–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31265-6_17/.

Goto:2012:SGM

[92] Keisuke Goto, Hideo Bannai, Shunsuke Inenaga, and Masayuki Takeda. Speeding up q-gram mining on grammar-based compressed texts. Lecture Notes in Computer Science, 7354:220–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31265-6_18/.

Gawrychowski:2012:SEL

[93] Paweł Gawrychowski. Simple and efficient LZW-compressed multiple pattern matching. Lecture Notes in Computer Science, 7354:232–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31265-6_19/.

Ohlebusch:2012:CBW

[94] Enno Ohlebusch, Timo Beller, and Mohamed I. Abouelhoda. Computing the Burrows–Wheeler transform of a string and its reverse. Lecture Notes in Computer Science, 7354:243–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31265-6_20/.

Anonymous:2012:FMe

[95] Anonymous. Front matter. Lecture Notes in Computer Science, 7354: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31265-6/1.

Caragiannis:2012:SLB

[96] Ioannis Caragiannis and Christos Kalaitzis. Space lower bounds for lowstretch greedy embeddings. Lecture Notes in Computer Science, 7355: 1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31104-8_1/.

Chechik:2012:FTC

[97] Shiri Chechik and David Peleg. The fault tolerant capacitated k-center problem. Lecture Notes in Computer Science, 7355:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31104-8_2/.

Emek:2012:NCO

[98] Yuval Emek, Pierre Fraigniaud, and Amos Korman. Notions of connectivity in overlay networks. Lecture Notes in Computer Science, 7355: 25–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31104-8_3/.

Bar-Noy:2012:CGS

[99] Amotz Bar-Noy, Ben Baumer, and Dror Rawitz. Changing of the guards: Strip cover with duty cycling. Lecture Notes in Computer Science, 7355: 36–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31104-8_4/.

Hasemann:2012:DLA

[100] Henning Hasemann and Juho Hirvonen. Deterministic local algorithms, unique identifiers, and fractional graph colouring. Lecture Notes in Computer Science, 7355:48–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31104-8_5/.

Kling:2012:AOF

[101] Peter Kling and Friedhelm Meyer auf der Heide. An algorithm for online facility leasing. Lecture Notes in Computer Science, 7355:61–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31104-8_6/.

Biely:2012:ADD

[102] Martin Biely, Peter Robinson, and Ulrich Schmid. Agreement in directed dynamic networks. Lecture Notes in Computer Science, 7355:73– 84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31104-8_7/.

Khabbazian:2012:BIW

[103] Majid Khabbazian and Stephane Durocher. Bounding interference in wireless ad hoc networks with nodes in random position. Lecture Notes in Computer Science, 7355:85–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31104-8_8/.

Hesari:2012:SCS

[104] Mohsen Eftekhari Hesari and Evangelos Kranakis. Strong connectivity of sensor networks with double antennae. Lecture Notes in Computer Science, 7355:99–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31104-8_9/.

Yu:2012:DMM

[105] Dongxiao Yu, Qiang-Sheng Hua, and Yuexuan Wang. Distributed multiple-message broadcast in wireless ad-hoc networks under the SINR model. Lecture Notes in Computer Science, 7355:111–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31104-8_10/.

Asgeirsson:2012:WNS

[106] Eyjólfur Ingi Ásgeirsson. Wireless network stability in the SINR model. Lecture Notes in Computer Science, 7355:123–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31104-8_11/.

Arfaoui:2012:WCC

[107] Heger Arfaoui and Pierre Fraigniaud. What can be computed without communications? Lecture Notes in Computer Science, 7355:135– 146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31104-8_12/.

Bilo:2012:BCG

[108] Vittorio Bilò and Michele Flammini. On bidimensional congestion games. Lecture Notes in Computer Science, 7355:147–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31104-8_13/.

Flammini:2012:MNC

[109] Michele Flammini and Vasco Gallotti. Mobile network creation games. Lecture Notes in Computer Science, 7355:159–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31104-8_14/.

Delporte-Gallet:2012:HFI

[110] Carole Delporte-Gallet and Hugues Fauconnier. Homonyms with forgeable identifiers. Lecture Notes in Computer Science, 7355:171–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31104-8_15/.

Vilaca:2012:ACP

[111] Xavier Vilaça and Oksana Denysyuk. Asynchrony and collusion in the N-party BAR transfer problem. Lecture Notes in Computer Science, 7355:183–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31104-8_16/.

Alistarh:2012:EDS

[112] Dan Alistarh, Hagit Attiya, and Rachid Guerraoui. Early deciding synchronous renaming in O(log f) rounds or less. Lecture Notes in Computer Science, 7355:195–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31104-8_17/.

Chalopin:2012:SSP

[113] Jérémie Chalopin and Yves Métivier. On snapshots and stable properties detection in anonymous fully distributed systems (extended abstract). Lecture Notes in Computer Science, 7355:207–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31104-8_18/.

Larsson:2012:SSC

[114] Andreas Larsson and Philippas Tsigas. Self-stabilizing (k, r)-clustering in clock rate-limited systems. Lecture Notes in Computer Science, 7355:219–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31104-8_19/.

Raynal:2012:IPI

[115] Michel Raynal and Julien Stainer. Increasing the power of the iterated immediate snapshot model with failure detectors. Lecture Notes in Computer Science, 7355:231–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31104-8_20/.

Anonymous:2012:FMf

[116] Anonymous. Front matter. Lecture Notes in Computer Science, 7355: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31104-8/1.

Mahdabi:2012:LBP

[117] Parvaz Mahdabi and Fabio Crestani. Learning-based pseudo-relevance feedback for patent retrieval. Lecture Notes in Computer Science, 7356: 1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31274-8_1/.

Waschle:2012:APD

[118] Katharina Wäschle and Stefan Riezler. Analyzing parallelism and domain similarities in the MAREC patent corpus. Lecture Notes in Computer Science, 7356:12–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31274-8_2/.

Nguyen:2012:QEP

[119] Khanh-Ly Nguyen and Sung-Hyon Myaeng. Query enhancement for patent prior-art-search based on keyterm dependency relations and semantic tags. Lecture Notes in Computer Science, 7356:28–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31274-8_3/.

Prieto:2012:ADW

[120] Víctor M. Prieto, Manuel Álvarez, Rafael López-García, and Fidel Cacheda. Analysis and detection of Web spam by means of Web content. Lecture Notes in Computer Science, 7356:43–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31274-8_4/.

Moumtzidou:2012:DEN

[121] Anastasia Moumtzidou, Stefanos Vrochidis, Sara Tonelli, and Ioannis Kompatsiaris. Discovery of environmental nodes in the Web. Lecture Notes in Computer Science, 7356:58–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31274-8_5/.

Fafalios:2012:WSE

[122] Pavlos Fafalios, Ioannis Kitsos, Yannis Marketakis, and Claudio Baldassarre. Web searching with entity mining at query time. Lecture Notes in Computer Science, 7356:73–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31274-8_6/.

Schleser:2012:PNE

[123] Tobias Schleser, Doris Ipsmiller, and Roland Mörzinger. Patent and norm exploration with the m2n knowledge discovery suite. Lecture Notes in Computer Science, 7356:89–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31274-8_7/.

Gomez:2012:HCW

[124] Juan Carlos Gomez and Marie-Francine Moens. Hierarchical classification of Web documents by stratified discriminant analysis. Lecture Notes in Computer Science, 7356:94–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31274-8_8/.

Vrochidis:2012:PAP

[125] Stefanos Vrochidis, Anastasia Moumtzidou, Gerard Ypma, and Ioannis Kompatsiaris. PatMedia: Augmenting patent search with contentbased image retrieval. Lecture Notes in Computer Science, 7356:109– 112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31274-8_9/.

Kettunen:2012:GVK

[126] Kimmo Kettunen and Paavo Arvola. Generating variant keyword forms for a morphologically complex language leads to successful information retrieval with Finnish. Lecture Notes in Computer Science, 7356:113– 126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31274-8_10/.

Tannebaum:2012:AQL

[127] Wolfgang Tannebaum and Andreas Rauber. Analyzing query logs of USPTO examiners to identify useful query terms in patent documents for query expansion in patent searching: a preliminary study. Lecture Notes in Computer Science, 7356:127–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31274-8_11/.

Pipanmaekaporn:2012:DRF

[128] Luepol Pipanmaekaporn and Yuefeng Li. Discovering relevant features for effective query formulation. Lecture Notes in Computer Science, 7356:137–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31274-8_12/.

Anonymous:2012:BMc

[129] Anonymous. Back matter. Lecture Notes in Computer Science, 7356: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31274-8/1.

Anonymous:2012:FMg

[130] Anonymous. Front matter. Lecture Notes in Computer Science, 7356: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31274-8/1.

Ghodsi:2012:V

[131] Mohammad Ghodsi, Anil Maheshwari, and Mostafa Nouri. α-visibility. Lecture Notes in Computer Science, 7357:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31155-0_1/.

Sherette:2012:PMB

[132] Jessica Sherette and Carola Wenk. Partial matching between surfaces using Fréchet distance. Lecture Notes in Computer Science, 7357:13– 23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31155-0_2/.

Ito:2012:PTA

[133] Takehiro Ito, Shin-Ichi Nakano, Yoshio Okamoto, and Yota Otachi. A polynomial-time approximation scheme for the geometric unique coverage problem on unit squares. Lecture Notes in Computer Science, 7357: 24–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31155-0_3/.

Dumitrescu:2012:WRL

[134] Adrian Dumitrescu, Joseph S. B. Mitchell, and Paweł Żyliński. Watchman routes for lines and segments. Lecture Notes in Computer Science, 7357:36–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31155-0_4/.

Abam:2012:KPD

[135] Mohammad Ali Abam, Zahed Rahmati, and Alireza Zarei. Kinetic Pie Delaunay graph and its applications. Lecture Notes in Computer Science, 7357:48–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31155-0_5/.

Gemsa:2012:HOC

[136] Andreas Gemsa, D. T. Lee, Chih-Hung Liu, and Dorothea Wagner. Higher order city Voronoi diagrams. Lecture Notes in Computer Science, 7357:59–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31155-0_6/.

Behsaz:2012:MSR

[137] Babak Behsaz and Mohammad R. Salavatipour. On minimum sum of radii and diameters clustering. Lecture Notes in Computer Science, 7357: 71–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31155-0_7/.

Hruz:2012:SFG

[138] Tomas Hruz and Marcel Schöngens. A simple framework for the generalized nearest neighbor problem. Lecture Notes in Computer Science, 7357: 83–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31155-0_8/.

Cygan:2012:DPC

[139] Marek Cygan. Deterministic parameterized connected vertex cover. Lecture Notes in Computer Science, 7357:95–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31155-0_9/.

Ghosh:2012:FPA

[140] Esha Ghosh, Sudeshna Kolay, Mrinal Kumar, and Pranabendu Misra. Faster parameterized algorithms for deletion to split graphs. Lecture Notes in Computer Science, 7357:107–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31155-0_10/.

Kim:2012:SEF

[141] Eun Jung Kim, Christophe Paul, and Geevarghese Philip. A singleexponential FPT algorithm for the K_{4}-minor cover problem. Lecture Notes in Computer Science, 7357:119–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31155-0_11/.

Han:2012:TAA

[142] Yijie Han and Tadao Takaoka. An O(n^{3} log log n/ log^{2} n) time algorithm for all pairs shortest paths. Lecture Notes in Computer Science, 7357:131–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31155-0_12/.

Atminas:2012:LTA

[143] Aistis Atminas, Vadim V. Lozin, and Igor Razgon. Linear time algorithm for computing a small biclique in graphs without Long induced paths. Lecture Notes in Computer Science, 7357:142–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31155-0_13/.

Golovach:2012:IDP

[144] Petr A. Golovach, Daniël Paulusma, and Erik Jan van Leeuwen. Induced disjoint paths in AT-free graphs. Lecture Notes in Computer Science, 7357:153–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31155-0_14/.

Giannopoulou:2012:ECI

[145] Archontia C. Giannopoulou, Iosif Salem, and Dimitris Zoros. Effective computation of immersion obstructions for unions of graph classes. Lecture Notes in Computer Science, 7357:165–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31155-0_15/.

Chen:2012:AMM

[146] Danny Z. Chen, Yan Gu, Jian Li, and Haitao Wang. Algorithms on minimizing the maximum sensor movement for barrier coverage of a linear domain. Lecture Notes in Computer Science, 7357:177–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31155-0_16/.

Busaryev:2012:ASH

[147] Oleksiy Busaryev, Sergio Cabello, Chao Chen, and Tamal K. Dey. Annotating simplices with a homology basis and its applications. Lecture Notes in Computer Science, 7357:189–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31155-0_17/.

Aschner:2012:DDA

[148] Rom Aschner, Matthew J. Katz, and Gila Morgenstern. Do directional antennas facilitate in reducing interferences? Lecture Notes in Computer Science, 7357:201–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31155-0_18/.

Dumitrescu:2012:MCP

[149] Adrian Dumitrescu, Sariel Har-Peled, and Csaba D. Tóth. Minimum convex partitions and maximum empty polytopes. Lecture Notes in Computer Science, 7357:213–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31155-0_19/.

Blaser:2012:PAC

[150] Markus Bläser and Konstatinos Panagiotou. A probabilistic analysis of Christofides’ Algorithm. Lecture Notes in Computer Science, 7357:225– 236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31155-0_20/.

Anonymous:2012:FMh

[151] Anonymous. Front matter. Lecture Notes in Computer Science, 7357: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31155-0/1.

Thomas:2012:SSC

[152] Wolfgang Thomas. Synthesis and some of its challenges. Lecture Notes in Computer Science, 7358:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-31424-7_1.

Dill:2012:MCC

[153] David L. Dill. Model checking cell biology. Lecture Notes in Computer Science, 7358:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-31424-7_2.

Bodik:2012:SPC

[154] Rastislav Bodik and Emina Torlak. Synthesizing programs with constraint solvers. Lecture Notes in Computer Science, 7358:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-31424-7_3.

Bradley:2012:IBI

[155] Aaron R. Bradley. IC3 and beyond: Incremental, inductive verification. Lecture Notes in Computer Science, 7358:4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-31424-7_4.

Myers:2012:FVG

[156] Chris J. Myers. Formal verification of genetic circuits. Lecture Notes in Computer Science, 7358:5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-31424-7_5.

Moskal:2012:CIB

[157] Michał Moskal. From C to infinity and back: Unbounded auto-active verification with VCC. Lecture Notes in Computer Science, 7358:6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-31424-7_6.

Kretinsky:2012:DAF

[158] Jan Křetínský and Javier Esparza. Deterministic automata for the (F, G)-fragment of LTL. Lecture Notes in Computer Science, 7358:7– 22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31424-7_7/.

Brazdil:2012:ECS

[159] Tomáš Brázdil, Krishnendu Chatterjee, Antonín Kučera, and Petr Novotný. Efficient controller synthesis for consumption games with multiple resource types. Lecture Notes in Computer Science, 7358:23–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31424-7_8/.

Ehlers:2012:ALS

[160] Rüdiger Ehlers. ACTL ∩ LTL synthesis. Lecture Notes in Computer Science, 7358:39–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31424-7_9/.

Chen:2012:LBF

[161] Yu-Fang Chen and Bow-Yaw Wang. Learning Boolean functions incrementally. Lecture Notes in Computer Science, 7358:55–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31424-7_10/.

Sharma:2012:IC

[162] Rahul Sharma, Aditya V. Nori, and Alex Aiken. Interpolants as classifiers. Lecture Notes in Computer Science, 7358:71–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31424-7_11/.

Lee:2012:TAA

[163] Wonchan Lee, Bow-Yaw Wang, and Kwangkeun Yi. Termination analysis with algorithmic learning. Lecture Notes in Computer Science, 7358:88– 104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31424-7_12/.

Brockschmidt:2012:ATP

[164] Marc Brockschmidt, Richard Musiol, Carsten Otto, and Jürgen Giesl. Automated termination proofs for Java programs with cyclic data. Lecture Notes in Computer Science, 7358:105–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31424-7_13/.

Esparza:2012:PTP

[165] Javier Esparza, Andreas Gaiser, and Stefan Kiefer. Proving termination of probabilistic programs using patterns. Lecture Notes in Computer Science, 7358:123–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31424-7_14/.

Venet:2012:GDS

[166] Arnaud J. Venet. The gauge domain: Scalable analysis of linear inequality invariants. Lecture Notes in Computer Science, 7358:139–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31424-7_15/.

Berdine:2012:DAF

[167] Josh Berdine, Arlen Cox, Samin Ishtiaq, and Christoph M. Wintersteiger. Diagnosing abstraction failure for separation logic–based analyses. Lecture Notes in Computer Science, 7358:155–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31424-7_16/.

Thakur:2012:MSC

[168] Aditya Thakur and Thomas Reps. A method for symbolic computation of abstract operations. Lecture Notes in Computer Science, 7358:174– 192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31424-7_17/.

Rollini:2012:LIS

[169] Simone Fulvio Rollini, Ondrej Sery, and Natasha Sharygina. Leveraging interpolant strength in model checking. Lecture Notes in Computer Science, 7358:193–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31424-7_18/.

Atig:2012:DFN

[170] Mohamed Faouzi Atig, Ahmed Bouajjani, Michael Emmi, and Akash Lal. Detecting fair non-termination in multithreaded programs. Lecture Notes in Computer Science, 7358:210–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31424-7_19/.

Kahlon:2012:LRC

[171] Vineet Kahlon and Chao Wang. Lock removal for concurrent trace programs. Lecture Notes in Computer Science, 7358:227–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31424-7_20/.

Anonymous:2012:FMi

[172] Anonymous. Front matter. Lecture Notes in Computer Science, 7358: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31424-7/1.

Donoghue:2012:RGR

[173] Claire R. Donoghue, Anil Rao, Anthony M. J. Bull, and Daniel Rueckert. Robust global registration through geodesic paths on an empirical manifold with knee MRI from the osteoarthritis initiative (OAI). Lecture Notes in Computer Science, 7359:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31340-0_1/.

Hong:2012:SGR

[174] Yi Hong, Yundi Shi, Martin Styner, Mar Sanchez, and Marc Niethammer. Simple geodesic regression for image time-series. Lecture Notes in Computer Science, 7359:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31340-0_2/.

Datteri:2012:ADM

[175] Ryan D. Datteri and Benoît M. Dawant. Automatic detection of the magnitude and spatial location of error in non-rigid registration. Lecture Notes in Computer Science, 7359:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31340-0_3/.

Tustison:2012:DDM

[176] Nicholas J. Tustison and Brian B. Avants. Diffeomorphic directly manipulated free-form deformation image registration via vector field flows. Lecture Notes in Computer Science, 7359:31–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31340-0_4/.

Forsberg:2012:MMI

[177] Daniel Forsberg, Gunnar Farnebäck, Hans Knutsson, and Carl-Fredrik Westin. Multi-modal image registration using polynomial expansion and mutual information. Lecture Notes in Computer Science, 7359:40–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31340-0_5/.

Janoos:2012:BCU

[178] Firdaus Janoos, Petter Risholm, and William Wells III. Bayesian characterization of uncertainty in multi-modal image registration. Lecture Notes in Computer Science, 7359:50–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31340-0_6/.

Sun:2012:HVS

[179] Wei Sun, Wiro J. Niessen, and Stefan Klein. Hierarchical vs. simultaneous multiresolution strategies for nonrigid image registration. Lecture Notes in Computer Science, 7359:60–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31340-0_7/.

Thivierge-Gaulin:2012:RBM

[180] David Thivierge-Gaulin, Chen-Rui Chou, Atilla P. Kiraly, and Christophe Chefd’Hotel. 3D–2D registration based on mesh-derived image bisection. Lecture Notes in Computer Science, 7359:70–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-31340-0_
8/.

Modat:2012:ICS

[181] Marc Modat, M. Jorge Cardoso, Pankaj Daga, David Cash, Nick C. Fox, and Sébastien Ourselin. Inverse-consistent symmetric free form deformation. Lecture Notes in Computer Science, 7359:79–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31340-0_9/.

Reda:2012:FAS

[182] Fitsum A. Reda, Jack H. Noble, Robert F. Labadie, and Benoît M. Dawant. Fully automatic surface-based pre- to intra-operative CT registration for cochlear implant. Lecture Notes in Computer Science, 7359: 89–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31340-0_10/.

Csapo:2012:TDI

[183] Istvan Csapo, Brad Davis, Yundi Shi, Mar Sanchez, Martin Styner, and Marc Niethammer. Temporally-dependent image similarity measure for longitudinal analysis. Lecture Notes in Computer Science, 7359:99–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31340-0_11/.

Compte:2012:CFS

[184] Benoît Compte, Adrien Bartoli, and Daniel Pizarro. Constant flow sampling: a method to automatically select the regularization parameter in image registration. Lecture Notes in Computer Science, 7359:110– 119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31340-0_12/.

Watanabe:2012:SCR

[185] Takanori Watanabe and Clayton Scott. Spatial confidence regions for quantifying and visualizing registration uncertainty. Lecture Notes in Computer Science, 7359:120–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31340-0_13/.

Carvalho:2012:RFH

[186] Diego D. B. Carvalho, Stefan Klein, Zeynettin Akkus, Gerrit L. ten Kate, and Hui Tang. Registration of free-hand ultrasound and MRI of carotid arteries through combination of point-based and intensitybased algorithms. Lecture Notes in Computer Science, 7359:131–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31340-0_14/.

Katouzian:2012:IHI

[187] Amin Katouzian, Athanasios Karamalis, Jennifer Lisauskas, Abouzar Eslami, and Nassir Navab. IVUS-histology image registration. Lecture Notes in Computer Science, 7359:141–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31340-0_15/.

Oreshkin:2012:ORG

[188] Boris N. Oreshkin and Tal Arbel. Optimization over random and gradient probabilistic pixel sampling for fast, robust multi-resolution image registration. Lecture Notes in Computer Science, 7359:150–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31340-0_16/.

Spiclin:2012:QTB

[189] Žiga Špiclin, Boštjan Likar, and Franjo Pernuš. Quad-tree based entropy estimator for fast and robust brain image registration. Lecture Notes in Computer Science, 7359:160–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31340-0_17/.

Gupta:2012:TNI

[190] Aditya Gupta, Maria Escolar, Cheryl Dietrich, John Gilmore, Guido Gerig, and Martin Styner. 3D tensor normalization for improved accuracy in DTI tensor registration methods. Lecture Notes in Computer Science, 7359:170–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31340-0_18/.

Joshi:2012:MAC

[191] Anand A. Joshi, David W. Shattuck, and Richard M. Leahy. A method for automated cortical surface registration and labeling. Lecture Notes in Computer Science, 7359:180–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31340-0_19/.

Ruthotto:2012:RDC

[192] Lars Ruthotto, Erlend Hodneland, and Jan Modersitzki. Registration of dynamic contrast enhanced MRI with local rigidity constraint. Lecture Notes in Computer Science, 7359:190–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31340-0_20/.

Anonymous:2012:FMj

[193] Anonymous. Front matter. Lecture Notes in Computer Science, 7359: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31340-0/1.

Bench-Capon:2012:MSM

[194] Trevor Bench-Capon. Marek sergot: a memoir. Lecture Notes in Computer Science, 7360:1–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29414-3_1/.

Barker:2012:SCM

[195] Steve Barker, Andrew J. I. Jones, Antonis Kakas, Robert A. Kowalski, and Alessio Lomuscio. The scientific contribution of marek sergot. Lecture Notes in Computer Science, 7360:4–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29414-3_2/.

Kowalski:2012:TRA

[196] Robert A. Kowalski and Fariba Sadri. Teleo-reactive abductive logic programs. Lecture Notes in Computer Science, 7360:12–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29414-3_3/.

Mancarella:2012:SNA

[197] Paolo Mancarella and Francesca Toni. Semi-negative abductive logic programs with implicative integrity constraints: Semantics and properties. Lecture Notes in Computer Science, 7360:33–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29414-3_4/.

Gabbay:2012:WNF

[198] Dov M. Gabbay. What is negation as failure? Lecture Notes in Computer Science, 7360:52–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29414-3_5/.

Monica:2012:IPI

[199] Dario Della Monica, Angelo Montanari, and Pietro Sala. The importance of the past in interval temporal logics: The case of propositional neighborhood logic. Lecture Notes in Computer Science, 7360:79–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29414-3_6/.

Hadjisoteriou:2012:AEC

[200] Evgenios Hadjisoteriou and Antonis Kakas. Argumentation and the event calculus. Lecture Notes in Computer Science, 7360:103–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29414-3_7/.

Bragaglia:2012:REC

[201] Stefano Bragaglia, Federico Chesani, Paola Mello, Marco Montali, and Paolo Torroni. Reactive event calculus for monitoring global computing applications. Lecture Notes in Computer Science, 7360:123–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29414-3_8/.

Blount:2012:RAI

[202] Justin Blount and Michael Gelfond. Reasoning about the intentions of agents. Lecture Notes in Computer Science, 7360:147–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29414-3_9/.

Lomuscio:2012:SMC

[203] Alessio Lomuscio and Wojciech Penczek. Symbolic model checking for temporal-epistemic logic. Lecture Notes in Computer Science, 7360:172– 195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29414-3_10/.

Hindriks:2012:GAI

[204] Koen V. Hindriks, Wiebe van der Hoek, and John-Jules Ch. Meyer. GOAL agents instantiate intention logic. Lecture Notes in Computer Science, 7360:196–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29414-3_11/.

Bench-Capon:2012:OTA

[205] Trevor Bench-Capon. Open texture and argumentation: What makes an argument persuasive? Lecture Notes in Computer Science, 7360:220– 233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29414-3_12/.

Dunne:2012:IPA

[206] Paul E. Dunne. Irrationality in persuasive argumentation. Lecture Notes in Computer Science, 7360:234–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29414-3_13/.

Prakken:2012:SRT

[207] Henry Prakken. Some reflections on two current trends in formal argumentation. Lecture Notes in Computer Science, 7360:249–272, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29414-3_14/.

Jones:2012:RNS

[208] Andrew J. I. Jones and Steven O. Kimbrough. On the representation of normative sentences in FOL. Lecture Notes in Computer Science, 7360:273–294, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29414-3_15/.

Parent:2012:WAI

[209] Xavier Parent. Why be afraid of identity? Lecture Notes in Computer Science, 7360:295–307, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29414-3_16/.

Alberti:2012:DAC

[210] Marco Alberti, Marco Gavanelli, and Evelina Lamma. Deon^{+}: Abduction and constraints for normative reasoning. Lecture Notes in Computer Science, 7360:308–328, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29414-3_17/.

Turrini:2012:CDG

[211] Paolo Turrini, Xavier Parent, Leendert van der Torre, and Silvano Colombo Tosatto. Contrary-to-duties in games. Lecture Notes in Computer Science, 7360:329–348, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29414-3_18/.

Barker:2012:LAA

[212] Steve Barker. Logical approaches to authorization policies. Lecture Notes in Computer Science, 7360:349–373, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29414-3_19/.

Minsky:2012:DGD

[213] Naftaly H. Minsky. Decentralized governance of distributed systems via interaction control. Lecture Notes in Computer Science, 7360:374–400, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29414-3_20/.

Anonymous:2012:BMd

[214] Anonymous. Back matter. Lecture Notes in Computer Science, 7360: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29414-3/1.

Anonymous:2012:FMk

[215] Anonymous. Front matter. Lecture Notes in Computer Science, 7360: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29414-3/1.

Hill:2012:PCE

[216] Melissa L. Hill, Kela Liu, James G. Mainprize, Ronald B. Levitin, and Rushin Shojaii. Pre-clinical evaluation of tumour angiogenesis with contrast-enhanced breast tomosynthesis. Lecture Notes in Computer Science, 7361:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31271-7_1/.

Hu:2012:EAS

[217] Yue-Houng Hu, David A. Scaduto, and Wei Zhao. The effect of amorphous selenium thickness on imaging performance of contrast enhanced digital breast tomosynthesis. Lecture Notes in Computer Science, 7361: 9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31271-7_2/.

Cruz-Bastida:2012:COC

[218] Juan-Pablo Cruz-Bastida, Iván Rosado-Méndez, Héctor Pérez-Ponce, and Yolanda Villaseñor. Contrast optimization in clinical contrastenhanced digital mammography images. Lecture Notes in Computer Science, 7361:17–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31271-7_3/.

Scaduto:2012:DSG

[219] David A. Scaduto and Wei Zhao. Determination of system geometrical parameters and consistency between scans for contrast-enhanced digital breast tomosynthesis. Lecture Notes in Computer Science, 7361:24– 31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31271-7_4/.

Gavenonis:2012:IED

[220] Sara Gavenonis, Kristen Lau, Roshan Karunamuni, Yiheng Zhang, Baorui Ren, and Chris Ruth. Initial experience with dual-energy contrast-enhanced digital breast tomosynthesis in the characterization of breast cancer. Lecture Notes in Computer Science, 7361:32–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31271-7_5/.

He:2012:MSR

[221] Wenda He, Erika R. E. Denton, and Reyer Zwiggelaar. Mammographic segmentation and risk classification using a novel binary model based Bayes classifier. Lecture Notes in Computer Science, 7361:40– 47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31271-7_6/.

Mertzanidou:2012:IBM

[222] Thomy Mertzanidou, John H. Hipwell, Lianghao Han, Zeike Taylor, and Henkjan Huisman. Intensity-based MRI to X-ray mammography registration with an integrated fast biomechanical transformation. Lecture Notes in Computer Science, 7361:48–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31271-7_7/.

Sharma:2012:CEM

[223] Diksha Sharma and Aldo Badano. Comparison of experimental, mantis, and hybrid mantis X-ray response for a breast imaging CsI detector. Lecture Notes in Computer Science, 7361:56–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31271-7_8/.

Narvaez:2012:BMC

[224] Fabián Narváez and Eduardo Romero. Breast mass classification using orthogonal moments. Lecture Notes in Computer Science, 7361:64– 71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31271-7_9/.

Young:2012:TSA

[225] Stefano Young, Andreu Badal, Kyle J. Myers, and Subok Park. A taskspecific argument for variable-exposure breast tomosynthesis. Lecture Notes in Computer Science, 7361:72–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31271-7_10/.

Patel:2012:DQE

[226] Tushita Patel, Kelly Klanian, Zongyi Gong, and Mark B. Williams. Detective quantum efficiency of a CsI-CMOS X-ray detector for breast tomosynthesis operating in high dynamic range and high sensitivity modes. Lecture Notes in Computer Science, 7361:80–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31271-7_11/.

Lesniak:2012:MCB

[227] Jan M. Lesniak, Guido van Schie, Christine Tanner, Bram Platel, and Henkjan Huisman. Multimodal classification of breast masses in mammography and MRI using unimodal feature selection and decision fusion. Lecture Notes in Computer Science, 7361:88–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31271-7_12/.

Bozek:2012:CLS

[228] Jelena Bozek, Michiel Kallenberg, Mislav Grgic, and Nico Karssemeijer. Comparison of lesion size using area and volume in full field digital mammograms. Lecture Notes in Computer Science, 7361:96–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31271-7_13/.

Ando:2012:VHC

[229] Masami Ando, Qingkai Huo, Shu Ichihara, Tokiko Endo, Tetsuya Yuasa, and Naoki Sunaguchi. Very high contrast and very high spatial resolution 2-D, 2.5-D and 3-D breast tissue visualization under X-ray dark field imaging. Lecture Notes in Computer Science, 7361:104–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31271-7_14/.

Bria:2012:DCM

[230] Alessandro Bria, Claudio Marrocco, Mario Molinara, and Francesco Tortorella. Detecting clusters of microcalcifications with a cascadebased approach. Lecture Notes in Computer Science, 7361:111–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31271-7_15/.

Uchiyama:2012:DIA

[231] Nachiko Uchiyama, Takayuki Kinoshita, Takashi Hojo, Sota Asaga, and Junko Suzuki. Diagnostic impact of adjunction of digital breast tomosynthesis (DBT) to full field digital mammography (FFDM) and in comparison with full field digital mammography (FFDM). Lecture Notes in Computer Science, 7361:119–126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31271-7_16/.

Hashmi:2012:EVV

[232] Sadaf Hashmi, Jamie C. Sergeant, Julie Morris, Sigrid Whiteside, and Paula Stavrinos. Ethnic variation in volumetric breast density. Lecture Notes in Computer Science, 7361:127–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31271-7_17/.

Porras-Chaverri:2012:PMD

[233] Mariela A. Porras-Chaverri, John R. Vetter, and Ralph Highnam. Personalizing mammographic dosimetry using multilayered anatomy-based breast models. Lecture Notes in Computer Science, 7361:134–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31271-7_18/.

Reiser:2012:DSS

[234] Ingrid Reiser, Beverly A. Lau, Robert M. Nishikawa, and Predrag R. Bakic. A directional small-scale tissue model for an anthropomorphic breast phantom. Lecture Notes in Computer Science, 7361:141–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31271-7_19/.

Chen:2012:STM

[235] Feiyu Chen, David D. Pokrajac, Xiquan Shi, Fengshan Liu, and Andrew D. A. Maidment. Simulation of three material partial volume averaging in a software breast phantom. Lecture Notes in Computer Science, 7361:149–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31271-7_20/.

Anonymous:2012:FMl

[236] Anonymous. Front matter. Lecture Notes in Computer Science, 7361: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31271-7/1.

Alama:2012:DFM

[237] Jesse Alama, Lionel Mamane, and Josef Urban. Dependencies in formal mathematics: Applications and extraction for Coq and Mizar. Lecture Notes in Computer Science, 7362:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31374-5_1/.

Asperti:2012:PMC

[238] Andrea Asperti. Proof, message and certificate. Lecture Notes in Computer Science, 7362:17–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31374-5_2/.

Bourke:2012:CEM

[239] Timothy Bourke, Matthias Daum, Gerwin Klein, and Rafal Kolanski. Challenges and experiences in managing large-scale proofs. Lecture Notes in Computer Science, 7362:32–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31374-5_3/.

David:2012:SAF

[240] Catalin David, Constantin Jucovschi, Andrea Kohlhase, and Michael Kohlhase. Semantic alliance: a framework for semantic allies. Lecture Notes in Computer Science, 7362:49–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31374-5_4/.

Horozal:2012:EMF

[241] Fulya Horozal, Michael Kohlhase, and Florian Rabe. Extending MKM formats at the statement level. Lecture Notes in Computer Science, 7362: 65–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31374-5_5/.

Hu:2012:SDI

[242] Rui Hu, Vadim Mazalov, and Stephen M. Watt. A streaming digital ink framework for multi-party collaboration. Lecture Notes in Computer Science, 7362:81–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31374-5_6/.

Jucovschi:2012:CEI

[243] Constantin Jucovschi. Cost-effective integration of MKM semantic services into editing environments. Lecture Notes in Computer Science, 7362:96–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31374-5_7/.

Libbrecht:2012:ULA

[244] Paul Libbrecht, Sandra Rebholz, Daniel Herding, Wolfgang Müller, and Felix Tscheulin. Understanding the learners’ actions when using mathematics learning tools. Lecture Notes in Computer Science, 7362:111– 126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31374-5_8/.

Marinkovic:2012:TUT

[245] Vesna Marinković and Predrag Janičić. Towards understanding triangle construction problems. Lecture Notes in Computer Science, 7362:127– 142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31374-5_9/.

Rabe:2012:QLF

[246] Florian Rabe. A query language for formal mathematical libraries. Lecture Notes in Computer Science, 7362:143–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31374-5_10/.

Sexton:2012:ASR

[247] Alan P. Sexton. Abramowitz and Stegun — a resource for mathematical document analysis. Lecture Notes in Computer Science, 7362:159– 168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31374-5_11/.

Tankink:2012:PWD

[248] Carst Tankink, Christoph Lange, and Josef Urban. Point-and-write — documenting formal mathematics by reference. Lecture Notes in Computer Science, 7362:169–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31374-5_12/.

Whiteside:2012:ES

[249] Iain Whiteside, David Aspinall, and Gudmund Grov. An essence of SSReflect. Lecture Notes in Computer Science, 7362:186–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31374-5_13/.

Carette:2012:TPC

[250] Jacques Carette and Russell O’Connor. Theory presentation combinators. Lecture Notes in Computer Science, 7362:202–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31374-5_14/.

Heras:2012:VAC

[251] Jónathan Heras, María Poza, and Julio Rubio. Verifying an algorithm computing discrete vector fields for digital imaging. Lecture Notes in Computer Science, 7362:216–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31374-5_15/.

Khan:2012:TFS

[252] Muhammad Taimoor Khan and Wolfgang Schreiner. Towards the formal specification and verification of Maple programs. Lecture Notes in Computer Science, 7362:231–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31374-5_16/.

Maric:2012:FFC

[253] Filip Marić, Miodrag Živković, and Bojan Vučković. Formalizing Frankl’s conjecture: FC-families. Lecture Notes in Computer Science, 7362:248– 263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31374-5_17/.

Nikolic:2012:CBA

[254] Mladen Nikolić and Predrag Janičić. CDCL-based abstract state transition system for coherent logic. Lecture Notes in Computer Science, 7362:264–279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31374-5_18/.

Wilson:2012:SCA

[255] David J. Wilson, Russell J. Bradford, and James H. Davenport. Speeding up cylindrical algebraic decomposition by Gröbner bases. Lecture Notes in Computer Science, 7362:280–294, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31374-5_19/.

DosReis:2012:SAP

[256] Gabriel Dos Reis. A system for axiomatic programming. Lecture Notes in Computer Science, 7362:295–309, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31374-5_20/.

Anonymous:2012:FMm

[257] Anonymous. Front matter. Lecture Notes in Computer Science, 7362: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31374-5/1.

Cichon:2012:MCE

[258] Jacek Cichoń, Jakub Lemiesz, and Marcin Zawada. On message complexity of extrema propagation techniques. Lecture Notes in Computer Science, 7363:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31638-8_1/.

Lee:2012:IAB

[259] Sang Hyuk Lee and Tomasz Radzik. Improved approximation bounds for maximum lifetime problems in wireless ad-hoc network. Lecture Notes in Computer Science, 7363:14–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31638-8_2/.

Merkel:2012:DGD

[260] Sabrina Merkel, Sanaz Mostaghim, and Hartmut Schmeck. Distributed geometric distance estimation in ad hoc networks. Lecture Notes in Computer Science, 7363:28–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31638-8_3/.

Mouradian:2012:DCB

[261] Alexandre Mouradian and Isabelle Augé-Blum. 1-D coordinate based on local information for MAC and routing issues in WSNs. Lecture Notes in Computer Science, 7363:42–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31638-8_4/.

Kranakis:2012:UCP

[262] Evangelos Kranakis, Fraser MacQuarie, Oscar Morales-Ponce, and Jorge Urrutia. Uninterrupted coverage of a planar region with rotating directional antennae. Lecture Notes in Computer Science, 7363:56–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31638-8_5/.

Ciobanu:2012:SAS

[263] Radu Ioan Ciobanu, Ciprian Dobre, and Valentin Cristea. Social aspects to support opportunistic networks in an academic environment. Lecture Notes in Computer Science, 7363:69–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31638-8_6/.

Asplund:2012:ADT

[264] Mikael Asplund and Simin Nadjm-Tehrani. Analysing delay-tolerant networks with correlated mobility. Lecture Notes in Computer Science, 7363: 83–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31638-8_7/.

Moreira:2012:SEN

[265] Waldir Moreira, Manuel de Souza, Paulo Mendes, and Susana Sargento. Study on the effect of network dynamics on opportunistic routing. Lecture Notes in Computer Science, 7363:98–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31638-8_8/.

Wodczak:2012:ACN

[266] Michał Wódczak. Autonomic cooperative networking for vehicular communications. Lecture Notes in Computer Science, 7363:112–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31638-8_9/.

Sarkar:2012:PDF

[267] Shikha Sarkar, Lina Stankovic, and Ivan Andonovic. Protocol design for farm animal monitoring using simulation. Lecture Notes in Computer Science, 7363:126–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31638-8_10/.

Wang:2012:MLA

[268] Guanyu Wang, Qiang-Sheng Hua, and Yuexuan Wang. Minimum latency aggregation scheduling for arbitrary tree topologies under the SINR model. Lecture Notes in Computer Science, 7363:139–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31638-8_11/.

Bahi:2012:ONA

[269] Jacques M. Bahi, Abdallah Makhoul, and Maguy Medlej. An optimized in-network aggregation scheme for data collection in periodic sensor networks. Lecture Notes in Computer Science, 7363:153–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31638-8_12/.

Boers:2012:IIA

[270] Nicholas M. Boers, Ioanis Nikolaidis, and Pawel Gburzynski. Impulsive interference avoidance in dense wireless sensor networks. Lecture Notes in Computer Science, 7363:167–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31638-8_13/.

Shi:2012:RSL

[271] Wei Shi, Meng Yao, and Jean-Pierre Corriveau. Resilient secure localization and detection of colluding attackers in WSNs. Lecture Notes in Computer Science, 7363:181–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31638-8_14/.

Tao:2012:LCD

[272] Dan Tao, Shaojie Tang, and Huadong Ma. Low cost data gathering using mobile hybrid sensor networks. Lecture Notes in Computer Science, 7363:193–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31638-8_15/.

Bimschas:2012:DIT

[273] Daniel Bimschas, Oliver Kleine, and Dennis Pfisterer. Debugging the Internet of Things: a 6LoWPAN/CoAP testbed infrastructure. Lecture Notes in Computer Science, 7363:207–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31638-8_16/.

Torres:2012:EEQ

[274] Álvaro Torres, Carlos T. Calafate, Juan-Carlos Cano, and Pietro Manzoni. Evaluating the effectiveness of a QoS framework for MANETs in a real testbed. Lecture Notes in Computer Science, 7363:221–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31638-8_17/.

Scherer:2012:WSN

[275] Thomas Scherer, Clemens Lombriser, Wolfgang Schott, Hong Linh Truong, and Beat Weiss. Wireless sensor network for continuous temperature monitoring in air-cooled data centers: Applications and measurement results. Lecture Notes in Computer Science, 7363:235–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31638-8_18/.

Li:2012:OPS

[276] Tzu Hao Li, Alexey Borisenko, and Miodrag Bolic. Open platform semipassive RFID tag. Lecture Notes in Computer Science, 7363:249–259, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31638-8_19/.

Sendra:2012:SOF

[277] Sandra Sendra, Jose V. Lamparero, Jaime Lloret, and Miguel Ardid. Study of the optimum frequency at 2.4GHz ISM band for underwater wireless ad hoc communications. Lecture Notes in Computer Science, 7363:260–273, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31638-8_20/.

Anonymous:2012:FMn

[278] Anonymous. Front matter. Lecture Notes in Computer Science, 7363: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31638-8/1.

Bjorner:2012:TSN

[279] Nikolaj Bjørner. Taking satisfiability to the next level with Z3. Lecture Notes in Computer Science, 7364:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31365-3_1/.

Matiyasevich:2012:ESA

[280] Yuri Matiyasevich. Enlarging the Scope of applicability of successful techniques for automated reasoning in mathematics. Lecture Notes in Computer Science, 7364:9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-31365-3_2.

Nieuwenhuis:2012:SSS

[281] Robert Nieuwenhuis. SAT and SMT are still resolution: Questions and challenges. Lecture Notes in Computer Science, 7364:10–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31365-3_3/.

Anantharaman:2012:UMS

[282] Siva Anantharaman, Serdar Erbatur, Christopher Lynch, and Paliath Narendran. Unification modulo synchronous distributivity. Lecture Notes in Computer Science, 7364:14–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31365-3_4/.

Baader:2012:SEU

[283] Franz Baader, Stefan Borgwardt, and Barbara Morawska. SAT encoding of unification in ELH_{R^{+}} w.r.t. cycle-restricted ontologies. Lecture Notes in Computer Science, 7364:30–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31365-3_5/.

Baader:2012:UUS

[284] Franz Baader, Julian Mendez, and Barbara Morawska. UEL: Unification solver for the description logic EL — system description. Lecture Notes in Computer Science, 7364:45–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31365-3_6/.

Baaz:2012:EFV

[285] Matthias Baaz, Ori Lahav, and Anna Zamansky. Effective finite-valued semantics for labelled calculi. Lecture Notes in Computer Science, 7364: 52–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31365-3_7/.

Bobot:2012:SBE

[286] François Bobot, Sylvain Conchon, Evelyne Contejean, Mohamed Iguernelala, and Assia Mahboubi. A simplex-based extension of Fourier– Motzkin for solving linear integer arithmetic. Lecture Notes in Computer Science, 7364:67–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31365-3_8/.

Borgwardt:2012:HFM

[287] Stefan Borgwardt, Felix Distel, and Rafael Peñaloza. How fuzzy is my fuzzy description logic? Lecture Notes in Computer Science, 7364:82– 96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31365-3_9/.

Brock-Nannestad:2012:TMA

[288] Taus Brock-Nannestad and Carsten Schürmann. Truthful monadic abstractions. Lecture Notes in Computer Science, 7364:97–110, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-31365-3_
10/.

Brown:2012:SAH

[289] Chad E. Brown. Satallax: An automatic higher-order prover. Lecture Notes in Computer Science, 7364:111–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31365-3_11/.

Bruttomesso:2012:SAM

[290] Roberto Bruttomesso, Silvio Ghilardi, and Silvio Ranise. From strong amalgamability to modularity of quantifier-free interpolation. Lecture Notes in Computer Science, 7364:118–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31365-3_12/.

Chekol:2012:SQC

[291] Melisachew Wudage Chekol, Jérôme Euzenat, Pierre Genevès, and Nabil Layaïda. SPARQL query containment under RDFS entailment regime. Lecture Notes in Computer Science, 7364:134–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31365-3_13/.

deBoer:2012:AVR

[292] Frank de Boer, Marcello Bonsangue, and Jurriaan Rot. Automated verification of recursive programs with pointers. Lecture Notes in Computer Science, 7364:149–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31365-3_14/.

Delaune:2012:SPC

[293] Stéphanie Delaune, Steve Kremer, and Daniel Pasaila. Security protocols, constraint systems, and group theories. Lecture Notes in Computer Science, 7364:164–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31365-3_15/.

Demri:2012:TPL

[294] Stéphane Demri, Amit Kumar Dhar, and Arnaud Sangnier. Taming past LTL and flat counter systems. Lecture Notes in Computer Science, 7364:179–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31365-3_16/.

Echenim:2012:CGG

[295] Mnacho Echenim and Nicolas Peltier. A calculus for generating ground explanations. Lecture Notes in Computer Science, 7364:194– 209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31365-3_17/.

Emmer:2012:EBB

[296] Moshe Emmer, Zurab Khasidashvili, Konstantin Korovin, and Christoph Sticksel. EPR-based bounded model checking at word level. Lecture Notes in Computer Science, 7364:210–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31365-3_18/.

Emmes:2012:PNL

[297] Fabian Emmes, Tim Enger, and Jürgen Giesl. Proving non-looping non-termination automatically. Lecture Notes in Computer Science, 7364:225–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31365-3_19/.

Falke:2012:RIL

[298] Stephan Falke and Deepak Kapur. Rewriting induction + linear arithmetic = decision procedure. Lecture Notes in Computer Science, 7364:241–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31365-3_20/.

Anonymous:2012:FMo

[299] Anonymous. Front matter. Lecture Notes in Computer Science, 7364: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31365-3/1.

Reisig:2012:FLP

[300] Wolfgang Reisig. A fresh look at Petri net extensions. Lecture Notes in Computer Science, 7365:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30835-2_1/.

Gabriel:2012:MCP

[301] Karsten Gabriel and Hartmut Ehrig. Modelling of communication platforms using algebraic high-level nets and their processes. Lecture Notes in Computer Science, 7365:10–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30835-2_2/.

Weber:2012:MR

[302] Herbert Weber. Modeling revisited. Lecture Notes in Computer Science, 7365:26–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30835-2_3/.

Konersmann:2012:CFE

[303] Marco Konersmann and Michael Goedicke. A conceptual framework and experimental workbench for architectures. Lecture Notes in Computer Science, 7365:36–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30835-2_4/.

Tavangarian:2012:VCE

[304] Djamshid Tavangarian. Virtual computing: The Emperor’s new clothes? Lecture Notes in Computer Science, 7365:53–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30835-2_5/.

Leymann:2012:LCU

[305] Frank Leymann. Linked compute units and linked experiments: Using topology and orchestration technology for flexible support of scientific applications. Lecture Notes in Computer Science, 7365:71–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30835-2_6/.

Beckers:2012:PBC

[306] Kristian Beckers and Stephan Faßbender. Pattern-based context establishment for service-oriented architectures. Lecture Notes in Computer Science, 7365:81–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30835-2_7/.

Papazoglou:2012:CBI

[307] Michael P. Papazoglou. Cloud blueprints for integrating and managing cloud federations. Lecture Notes in Computer Science, 7365:102– 119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30835-2_8/.

Kruger:2012:DSC

[308] Ingolf Krüger and Barry Demchak. Dynamic service composition and deployment with OpenRichServices. Lecture Notes in Computer Science, 7365:120–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30835-2_9/.

Mechitov:2012:ADS

[309] Kirill Mechitov and Gul Agha. An architecture for dynamic serviceoriented computing in networked embedded systems. Lecture Notes in Computer Science, 7365:147–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30835-2_10/.

Rensing:2012:LBS

[310] Christoph Rensing and Stephan Tittel. Location-based services for technology enhanced learning and teaching. Lecture Notes in Computer Science, 7365:165–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30835-2_11/.

Anonymous:2012:BMe

[311] Anonymous. Back matter. Lecture Notes in Computer Science, 7365: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30835-2/1.

Anonymous:2012:FMp

[312] Anonymous. Front matter. Lecture Notes in Computer Science, 7365: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30835-2/1.

Olsher:2012:CBI

[313] Daniel J. Olsher. COGPARSE: Brain-inspired knowledge-driven full semantics parsing. Lecture Notes in Computer Science, 7366:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31561-9_1/.

Mazzocco:2012:SNN

[314] Thomas Mazzocco, Erik Cambria, Amir Hussain, and Qiu-Feng Wang. Sentic neural networks: a novel cognitive model for affective common sense reasoning. Lecture Notes in Computer Science, 7366:12–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31561-9_2/.

Rawlinson:2012:IDW

[315] Terry G. Rawlinson, Shulan Lu, and Patrick Coleman. Individual differences in working memory capacity and presence in virtual environments. Lecture Notes in Computer Science, 7366:22–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31561-9_3/.

Farooq:2012:ODB

[316] Kamran Farooq, Amir Hussain, Stephen Leslie, Chris Eckl, Calum MacRae, and Warner Slack. An ontology driven and Bayesian network based cardiovascular decision support framework. Lecture Notes in Computer Science, 7366:31–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31561-9_4/.

Farooq:2012:SIE

[317] Kamran Farooq, Amir Hussain, Stephen Leslie, Chris Eckl, Calum MacRae, and Warner Slack. Semantically inspired electronic healthcare records. Lecture Notes in Computer Science, 7366:42–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31561-9_5/.

Wei:2012:CBO

[318] Hui Wei and Zheng Dong. A CSP-based orientation detection model. Lecture Notes in Computer Science, 7366:52–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31561-9_6/.

Cakir:2012:EUC

[319] Murat Perit Çakır, Abdullah Murat Şenyiğit, Daryal Murat Akay, Hasan Ayaz, and Veysi İşler. Evaluation of UAS camera operator interfaces in a simulated task environment: An optical brain imaging approach. Lecture Notes in Computer Science, 7366:62–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31561-9_7/.

Li:2012:CAP

[320] Chunguang Li and Lining Sun. Cerebral activation patterns in the preparation and movement periods of spontaneous and evoked movements. Lecture Notes in Computer Science, 7366:72–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31561-9_8/.

Yang:2012:NIS

[321] Erfu Yang, Amir Hussain, and Kevin Gurney. Neurobiologically-inspired soft switching control of autonomous vehicles. Lecture Notes in Computer Science, 7366:82–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31561-9_9/.

Hussain:2012:IMC

[322] Amir Hussain, Rudwan Abdullah, Erfu Yang, and Kevin Gurney. An intelligent multiple-controller framework for the integrated control of autonomous vehicles. Lecture Notes in Computer Science, 7366:92–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31561-9_10/.

Harter:2012:ESW

[323] Derek Harter. Evolution of small-world properties in embodied networks. Lecture Notes in Computer Science, 7366:102–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31561-9_11/.

Qi:2012:BMI

[324] Yujuan Qi, Yanjiang Wang, and Tingting Xue. Brain memory inspired template updating modeling for robust moving object tracking using particle filter. Lecture Notes in Computer Science, 7366:112–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31561-9_12/.

Huo:2012:VIB

[325] Juan Huo and Alan Murray. VLSI implementation of barn owl superior colliculus network for visual and auditory integration. Lecture Notes in Computer Science, 7366:120–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31561-9_13/.

Wang:2012:MCO

[326] Fuluo Wang, Yourui Huang, Ming Shi, and Shanshan Wu. Membrane computing optimization method based on catalytic factor. Lecture Notes in Computer Science, 7366:129–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31561-9_14/.

Ozgoren:2012:EBP

[327] Murat Ozgoren, Merve Tetik, Kurtulus Izzetoglu, Adile Oniz, and Banu Onaral. Effect of body position on NIRS based hemodynamic measures from prefrontal cortex. Lecture Notes in Computer Science, 7366:138– 146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31561-9_15/.

Ayaz:2012:UBA

[328] Hasan Ayaz, Scott Bunce, Patricia Shewokis, Kurtulus Izzetoglu, and Ben Willems. Using brain activity to predict task performance and operator efficiency. Lecture Notes in Computer Science, 7366:147–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31561-9_16/.

Wang:2012:AAD

[329] Ligang Wang, Jie Kong, Jing Luo, Wenbin Gao, and Xianju Guo. “arousal” or “activation” dysfunction in the frontal region of children with attention–deficit/hyperactivity disorder: Evidence from an electroencephalogram study. Lecture Notes in Computer Science, 7366:156– 163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31561-9_17/.

Fagiani:2012:NIS

[330] Marco Fagiani, Emanuele Principi, Stefano Squartini, and Francesco Piazza. A new Italian sign language database. Lecture Notes in Computer Science, 7366:164–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31561-9_18/.

Hossain:2012:SPR

[331] Gahangir Hossain, Mark H. Myers, and Robert Kozma. Study of phase relationships in ECoG signals using Hilbert–Huang transforms. Lecture Notes in Computer Science, 7366:174–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31561-9_19/.

Bunce:2012:TSP

[332] Scott C. Bunce, Kurtulus Izzetoglu, Meltem Izzetoglu, Hasan Ayaz, and Kambiz Pourrezaei. Treatment status predicts differential prefrontal cortical responses to alcohol and natural reinforcer cues among alcohol dependent individuals. Lecture Notes in Computer Science, 7366:183– 191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31561-9_20/.

Anonymous:2012:FMq

[333] Anonymous. Front matter. Lecture Notes in Computer Science, 7366: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31561-9/1.

Frolov:2012:ANN

[334] Alexander A. Frolov, Dušan Húsek, and Pavel Yu. Polyakov. Attractor neural network combined with likelihood maximization algorithm for Boolean factor analysis. Lecture Notes in Computer Science, 7367:1– 10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31346-2_1/.

Tang:2012:PFN

[335] Zaiyong Tang, Kallol Bagchi, Youqin Pan, and Gary J. Koehler. Pruning feedforward neural network search space using local Lipschitz constants. Lecture Notes in Computer Science, 7367:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31346-2_2/.

Kim:2012:CFB

[336] Wook-Dong Kim, Sung-Kwun Oh, and Hyun-Ki Kim. Context FCMbased radial basis function neural networks with the aid of fuzzy clustering. Lecture Notes in Computer Science, 7367:21–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31346-2_3/.

Zhao:2012:MSD

[337] Li-Jie Zhao, Jian Tang, and Tian you Chai. Modeling spectral data based on mutual information and kernel extreme learning machines. Lecture Notes in Computer Science, 7367:29–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31346-2_4/.

Wang:2012:HNN

[338] Jing Wang, Haibo He, Yuan Cao, Jin Xu, and Dongbin Zhao. A hierarchical neural network architecture for classification. Lecture Notes in Computer Science, 7367:37–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31346-2_5/.

Ke:2012:DTZ

[339] Zhende Ke, Yiwen Yang, and Yunong Zhang. Discrete-time ZNN algorithms for time-varying quadratic programming subject to time-varying equality constraint. Lecture Notes in Computer Science, 7367:47–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31346-2_6/.

Zhu:2012:PPR

[340] Xibin Zhu, Frank-Michael Schleif, and Barbara Hammer. Patch processing for relational learning vector quantization. Lecture Notes in Computer Science, 7367:55–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31346-2_7/.

Zhang:2012:NNM

[341] Zheng Zhang. A neural network model for currency arbitrage detection. Lecture Notes in Computer Science, 7367:64–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31346-2_8/.

Lu:2012:RRM

[342] Shuxia Lu, Guiqiang Zhang, and Xizhao Wang. A rank reduced matrix method in extreme learning machine. Lecture Notes in Computer Science, 7367:72–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31346-2_9/.

Mao:2012:RDL

[343] Wentao Mao, Mei Tian, Guirong Yan, and Xianfang Wang. Research of dynamic load identification based on extreme learning machine. Lecture Notes in Computer Science, 7367:80–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31346-2_10/.

Huang:2012:FRB

[344] Wei Huang and Sung-Kwun Oh. Fuzzy relation-based polynomial neural networks based on hybrid optimization. Lecture Notes in Computer Science, 7367:90–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31346-2_11/.

Zhang:2012:TVM

[345] Yunong Zhang, Yunjia Xie, and Hongzhou Tan. Time-varying Moore– Penrose inverse solving shows different Zhang functions leading to different ZNN models. Lecture Notes in Computer Science, 7367:98–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31346-2_12/.

Zhang:2012:MOS

[346] Kun Zhang, Cuirong Wang, and Baoyan Wang. A multi-object segmentation algorithm based on background modeling and region growing. Lecture Notes in Computer Science, 7367:106–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31346-2_13/.

Zhang:2012:REU

[347] Wei-Feng Zhang, Peng Yang, Dao-Qing Dai, and Arye Nehorai. Reflectance estimation using local regression methods. Lecture Notes in Computer Science, 7367:116–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31346-2_14/.

Gu:2012:AND

[348] Lei Gu. Applying a novel decision rule to the semi-supervised clustering method based on one-class SVM. Lecture Notes in Computer Science, 7367:123–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31346-2_15/.

Huang:2012:SEM

[349] He Huang and Xiaoping Chen. State estimation of Markovian jump neural networks with mixed time delays. Lecture Notes in Computer Science, 7367:132–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31346-2_16/.

Li:2012:LBM

[350] Qiaojie Li, Zhoushun Zheng, Shuang Wang, and Jiankang Liu. Lattice Boltzmann model for nonlinear heat equations. Lecture Notes in Computer Science, 7367:140–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31346-2_17/.

Yang:2012:MOL

[351] Wenyu Yang, Jie Yang, and Wei Wu. A modified one-layer spiking neural network involves derivative of the state function at firing time. Lecture Notes in Computer Science, 7367:149–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31346-2_18/.

Zhang:2012:MMM

[352] Yingwei Zhang and Chuang Wang. Modeling and monitoring of multimodes process. Lecture Notes in Computer Science, 7367:159–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31346-2_19/.

Zhang:2012:DBM

[353] Yingwei Zhang and Hailong Zhang. Data-based modeling and monitoring for multimode processes using local tangent space alignment. Lecture Notes in Computer Science, 7367:169–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31346-2_20/.

Anonymous:2012:FMr

[354] Anonymous. Front matter. Lecture Notes in Computer Science, 7367: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31346-2/1.

Ma:2012:PCB

[355] Dazhong Ma, Jinhai Liu, and Zhanshan Wang. The pattern classification based on fuzzy min-max neural network with new algorithm. Lecture Notes in Computer Science, 7368:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31362-2_1/.

Zhao:2012:MCC

[356] Li jie Zhao, Tian you Chai, Xiao kun Diao, and De cheng Yuan. Multi-class classification with one-against-one using probabilistic extreme learning machine. Lecture Notes in Computer Science, 7368:10– 19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31362-2_2/.

Wang:2012:SMF

[357] Shangfei Wang, Shan He, and Hua Zhu. Similarity measurement and feature selection using genetic algorithm. Lecture Notes in Computer Science, 7368:20–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31362-2_3/.

Wang:2012:EFD

[358] Ting Wang, Sheng-Uei Guan, and Fei Liu. Entropic feature discrimination ability for pattern classification based on neural IAL. Lecture Notes in Computer Science, 7368:30–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31362-2_4/.

Kim:2012:DOR

[359] Wook-Dong Kim, Sung-Kwun Oh, and Jeong-Tae Kim. Design of optimized radial basis function neural networks classifier with the aid of fuzzy clustering and data preprocessing method. Lecture Notes in Computer Science, 7368:38–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31362-2_5/.

Dong:2012:EHB

[360] Yongsheng Dong and Jinwen Ma. An efficient histogram-based texture classification method with weighted symmetrized Kullback–Leibler divergence. Lecture Notes in Computer Science, 7368:46–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31362-2_6/.

Zhang:2012:RSI

[361] Qiaoge Zhang, Zhigang Liu, and Gang Chen. The recognition study of impulse and oscillation transient based on spectral kurtosis and neural network. Lecture Notes in Computer Science, 7368:56–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31362-2_7/.

Han:2012:FFS

[362] Min Han and Xiaoxin Liu. Forward feature selection based on approximate Markov blanket. Lecture Notes in Computer Science, 7368:64– 72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31362-2_8/.

Wu:2012:ARR

[363] Tianshu Wu, Kunqing Xie, Chengkai Nie, and Guojie Song. An adaption of relief for redundant feature elimination. Lecture Notes in Computer Science, 7368:73–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31362-2_9/.

Tang:2012:FSF

[364] Jian Tang, Li-Jie Zhao, Yi miao Li, Tian you Chai, and S. Joe Qin. Feature selection of frequency spectrum for modeling difficulty to measure process parameters. Lecture Notes in Computer Science, 7368:82– 91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31362-2_10/.

Tang:2012:NDL

[365] Zunyi Tang and Shuxue Ding. Nonnegative dictionary learning by nonnegative matrix factorization with a sparsity constraint. Lecture Notes in Computer Science, 7368:92–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31362-2_11/.

Chen:2012:NMH

[366] Dongyue Chen, Zongwen Chen, and Xiaosheng Yu. A new method for hand detection based on Hough forest. Lecture Notes in Computer Science, 7368:102–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31362-2_12/.

Li:2012:MSC

[367] Jia Li, Changyong Niu, and Ming Fan. Multi-scale convolutional neural networks for natural scene license plate detection. Lecture Notes in Computer Science, 7368:110–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31362-2_13/.

Liu:2012:RMS

[368] Zhao Liu, Guiyu Feng, and Dewen Hu. Robust mean shift tracking with background information. Lecture Notes in Computer Science, 7368:120– 129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31362-2_14/.

Jia:2012:HSC

[369] Lijuan Jia, Dandan Song, Linmi Tao, and Yao Lu. Heart sounds classification with a fuzzy neural network method with structure learning. Lecture Notes in Computer Science, 7368:130–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31362-2_15/.

Zhang:2012:CMH

[370] Ting Zhang, Guosheng Yang, and Xinkai Kuai. On cortex mechanism hierarchy model for facial expression recognition: Multi-database evaluation results. Lecture Notes in Computer Science, 7368:141–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31362-2_16/.

Saeed:2012:LLE

[371] Mehreen Saeed. LEFT–logical expressions feature transformation: a framework for transformation of symbolic features. Lecture Notes in Computer Science, 7368:149–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31362-2_17/.

Mai:2012:TFA

[372] Songping Mai, Yixin Zhao, Chun Zhang, and Zhihua Wang. A timefrequency aware cochlear implant: Algorithm and system. Lecture Notes in Computer Science, 7368:159–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31362-2_18/.

Yu:2012:GVF

[373] Xiaosheng Yu, Chengdong Wu, Dongyue Chen, Ting Zhou, and Tong Jia. Gradient vector flow based on anisotropic diffusion. Lecture Notes in Computer Science, 7368:169–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31362-2_19/.

Huang:2012:ECB

[374] Kai Huang, Liqing Zhang, and Yang Wu. ECG classification based on non-cardiology feature. Lecture Notes in Computer Science, 7368:179– 186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31362-2_20/.

Anonymous:2012:FMs

[375] Anonymous. Front matter. Lecture Notes in Computer Science, 7368: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31362-2/1.

Hess:2012:GJC

[376] Florian Hess. Generalised Jacobians in cryptography and coding theory. Lecture Notes in Computer Science, 7369:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31662-3_1/.

Vega:2012:WDF

[377] Gerardo Vega and Carlos A. Vázquez. The weight distribution of a family of reducible cyclic codes. Lecture Notes in Computer Science, 7369:

16–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31662-3_2/.

Geil:2012:NMC

[378] Olav Geil, Stefano Martin, and Ryutaroh Matsumoto. A new method for constructing small-bias spaces from Hermitian codes. Lecture Notes in Computer Science, 7369:29–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31662-3_3/.

Cayrel:2012:ITR

[379] Pierre-Louis Cayrel, Sidi Mohamed El Yousfi Alaoui, Gerhrad Hoffmann, and Pascal Véron. An improved threshold ring signature scheme based on error correcting codes. Lecture Notes in Computer Science, 7369:45– 63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31662-3_4/.

Pott:2012:SFD

[380] Alexander Pott, Qi Wang, and Yue Zhou. Sequences and functions derived from projective planes and their difference sets. Lecture Notes in Computer Science, 7369:64–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31662-3_5/.

Sok:2012:FSD

[381] Lin Sok and Patrick Solé. On formally self-dual Boolean functions in 2,4 and 6 variables. Lecture Notes in Computer Science, 7369:81–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31662-3_6/.

Batteux:2012:ANF

[382] Boris Batteux. On the algebraic normal form and Walsh spectrum of symmetric functions over finite rings. Lecture Notes in Computer Science, 7369:92–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31662-3_7/.

Budaghyan:2012:VRE

[383] Lilya Budaghyan and Oleksandr Kazymyrov. Verification of restricted EA-equivalence for vectorial Boolean functions. Lecture Notes in Computer Science, 7369:108–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31662-3_8/.

Gueron:2012:SIM

[384] Shay Gueron and Vlad Krasnov. Software implementation of modular exponentiation, using advanced vector instructions architectures. Lecture Notes in Computer Science, 7369:119–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31662-3_9/.

ElMrabet:2012:EME

[385] Nadia El Mrabet and Nicolas Gama. Efficient multiplication over extension fields. Lecture Notes in Computer Science, 7369:136–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31662-3_10/.

Pamula:2012:FFM

[386] Danuta Pamula and Arnaud Tisserand. gf(2^{m}) finite-field multipliers with reduced activity variations. Lecture Notes in Computer Science, 7369:152–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31662-3_11/.

Barbulescu:2012:FOF

[387] Razvan Barbulescu, Jérémie Detrey, Nicolas Estibals, and Paul Zimmermann. Finding optimal formulae for bilinear maps. Lecture Notes in Computer Science, 7369:168–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31662-3_12/.

Driessen:2012:SBL

[388] Benedikt Driessen and Christof Paar. Solving binary linear equation systems over the rationals and binaries. Lecture Notes in Computer Science, 7369:187–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31662-3_13/.

Omar:2012:HEC

[389] Sami Omar, Raouf Ouni, and Saber Bouanani. Hashing with elliptic curve L-functions. Lecture Notes in Computer Science, 7369:196–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31662-3_14/.

Thome:2012:SRA

[390] Emmanuel Thomé. Square root algorithms for the number field sieve. Lecture Notes in Computer Science, 7369:208–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31662-3_15/.

Harasawa:2012:IBA

[391] Ryuichi Harasawa, Yutaka Sueyoshi, and Aichi Kudo. Improving the Berlekamp algorithm for binomials x^{n} − a. Lecture Notes in Computer Science, 7369:225–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31662-3_16/.

Sarkar:2012:SPB

[392] Sumanta Sarkar, Srimanta Bhattacharya, and Ayça Çeşmelioğlu. On some permutation binomials of the form x^{\frac{2^{n}−1}{k} +1} + ax over F_{2^{n}}: Existence and count. Lecture Notes in Computer Science, 7369:236–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31662-3_17/.

Anonymous:2012:BMf

[393] Anonymous. Back matter. Lecture Notes in Computer Science, 7369: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31662-3/1.

Anonymous:2012:FMt

[394] Anonymous. Front matter. Lecture Notes in Computer Science, 7369: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31662-3/1.

Anonymous:2012:FMu

[395] Anonymous. Front matter. Lecture Notes in Computer Science, 7370: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30891-8/1/1.

Langston:2012:FPT

[396] Michael A. Langston. Fixed-parameter tractability, A prehistory. Lecture Notes in Computer Science, 7370:3–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30891-8_1/.

Downey:2012:BEY

[397] Rod Downey. The birth and early years of parameterized complexity. Lecture Notes in Computer Science, 7370:17–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30891-8_2/.

Koblitz:2012:CG

[398] Neal Koblitz. Crypto galore! Lecture Notes in Computer Science, 7370: 39–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30891-8_3/.

Wareham:2012:FLB

[399] Todd Wareham. Flyby: Life before, during, and after graduate studies with Mike Fellows. Lecture Notes in Computer Science, 7370:51– 55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30891-8_4/.

Stege:2012:IPC

[400] Ulrike Stege. The impact of parameterized complexity to interdisciplinary problem solving. Lecture Notes in Computer Science, 7370:56– 68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30891-8_5/.

Raman:2012:VCD

[401] Venkatesh Raman. Vertex cover, dominating set and my encounters with parameterized complexity and Mike Fellows. Lecture Notes in Computer Science, 7370:69–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30891-8_6/.

Telle:2012:MFW

[402] Jan Arne Telle. Mike fellows: Weaving the Web of mathematics and adventure. Lecture Notes in Computer Science, 7370:74–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30891-8_7/.

Rosamond:2012:PPM

[403] Frances Rosamond. Passion plays: Melodramas about mathematics. Lecture Notes in Computer Science, 7370:80–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30891-8_8/.

Anonymous:2012:FMv

[404] Anonymous. Front matter. Lecture Notes in Computer Science, 7370: 89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30891-8/2/1.

Downey:2012:BPC

[405] Rod Downey. A basic parameterized complexity primer. Lecture Notes in Computer Science, 7370:91–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30891-8_9/.

Lokshtanov:2012:KPG

[406] Daniel Lokshtanov, Neeldhara Misra, and Saket Saurabh. Kernelization — preprocessing with a guarantee. Lecture Notes in Computer Science, 7370:129–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30891-8_10/.

Chen:2012:PCS

[407] Jianer Chen and Iyad A. Kanj. Parameterized complexity and subexponential-time computability. Lecture Notes in Computer Science, 7370:162–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30891-8_11/.

Bodlaender:2012:FPT

[408] Hans L. Bodlaender. Fixed-parameter tractability of treewidth and pathwidth. Lecture Notes in Computer Science, 7370:196–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30891-8_12/.

Thilikos:2012:GMP

[409] Dimitrios M. Thilikos. Graph minors and parameterized algorithm design. Lecture Notes in Computer Science, 7370:228–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30891-8_13/.

Gutin:2012:CSP

[410] Gregory Gutin and Anders Yeo. Constraint satisfaction problems parameterized above or below tight bounds: a survey. Lecture Notes in Computer Science, 7370:257–286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30891-8_14/.

Gaspers:2012:BS

[411] Serge Gaspers and Stefan Szeider. Backdoors to satisfaction. Lecture Notes in Computer Science, 7370:287–317, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30891-8_15/.

Betzler:2012:SCA

[412] Nadja Betzler, Robert Bredereck, Jiehua Chen, and Rolf Niedermeier. Studies in computational aspects of voting. Lecture Notes in Computer Science, 7370:318–363, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30891-8_16/.

Chen:2012:PHP

[413] Yijia Chen and Jörg Flum. A parameterized halting problem. Lecture Notes in Computer Science, 7370:364–397, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30891-8_17/.

Bell:2012:CSU

[414] Tim Bell, Frances Rosamond, and Nancy Casey. Computer science unplugged and related projects in math and computer science popularization. Lecture Notes in Computer Science, 7370:398–456, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30891-8_18/.

Fomin:2012:FST

[415] Fedor V. Fomin and Dániel Marx. FPT suspects and tough customers: Open problems of downey and fellows. Lecture Notes in Computer Science, 7370:457–468, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30891-8_19/.

Marx:2012:WNF

[416] Dániel Marx. What’s next? future directions in parameterized complexity. Lecture Notes in Computer Science, 7370:469–496, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30891-8_20/.

Anonymous:2012:BMg

[417] Anonymous. Back matter. Lecture Notes in Computer Science, 7370: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30891-8/1.

Anonymous:2012:FMw

[418] Anonymous. Front matter. Lecture Notes in Computer Science, 7370: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30891-8/1.

vanderTorre:2012:LSP

[419] Leendert van der Torre. Logics for security and privacy. Lecture Notes in Computer Science, 7371:1–7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31540-4_1/.

Cheng:2012:UUR

[420] Yuan Cheng, Jaehong Park, and Ravi Sandhu. A user-to-user relationship-based access control model for online social networks. Lecture Notes in Computer Science, 7371:8–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31540-4_2/.

Armando:2012:AEA

[421] Alessandro Armando and Silvio Ranise. Automated and efficient analysis of role-based access control with attributes. Lecture Notes in Computer Science, 7371:25–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31540-4_3/.

Jin:2012:UAB

[422] Xin Jin, Ram Krishnan, and Ravi Sandhu. A unified attribute-based access control model covering DAC, MAC and RBAC. Lecture Notes in Computer Science, 7371:41–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31540-4_4/.

Biskup:2012:SBI

[423] Joachim Biskup, Sven Hartmann, Sebastian Link, Jan-Hendrik Lochner, and Torsten Schlotmann. Signature-based inference-usability confinement for relational databases under functional and join dependencies. Lecture Notes in Computer Science, 7371:56–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31540-4_5/.

Adl:2012:PCA

[424] Rosa Karimi Adl, Mina Askari, Ken Barker, and Reihaneh Safavi-Naini. Privacy consensus in anonymization systems via game theory. Lecture Notes in Computer Science, 7371:74–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31540-4_6/.

Dini:2012:UOL

[425] Gianluca Dini and Pericle Perazzo. Uniform obfuscation for location privacy. Lecture Notes in Computer Science, 7371:90–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31540-4_7/.

Pippal:2012:SVU

[426] Ravi Singh Pippal, Jaidhar C. D., and Shashikala Tapaswi. Security vulnerabilities of user authentication scheme using Smart Card. Lecture Notes in Computer Science, 7371:106–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31540-4_8/.

Wang:2012:SPB

[427] Ding Wang, Chun guang Ma, and Peng Wu. Secure password-based remote user authentication scheme with non-tamper resistant Smart Cards. Lecture Notes in Computer Science, 7371:114–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31540-4_9/.

Razafindralambo:2012:FFH

[428] Tiana Razafindralambo, Guillaume Bouffard, and Jean-Louis Lanet. A friendly framework for hidding fault enabled virus for Java based Smartcard. Lecture Notes in Computer Science, 7371:122–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31540-4_10/.

Nix:2012:APP

[429] Robert Nix, Murat Kantarcioglu, and Keesook J. Han. Approximate privacy-preserving data mining on vertically partitioned data. Lecture Notes in Computer Science, 7371:129–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31540-4_11/.

Dautrich:2012:SLU

[430] Jonathan L. Dautrich and Chinya V. Ravishankar. Security limitations of using secret sharing for data outsourcing. Lecture Notes in Computer Science, 7371:145–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31540-4_12/.

Mehmood:2012:PPS

[431] Danish Mehmood, Basit Shafiq, Jaideep Vaidya, Yuan Hong, and Nabil Adam. Privacy-preserving subgraph discovery. Lecture Notes in Computer Science, 7371:161–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31540-4_13/.

Foley:2012:DST

[432] Simon N. Foley and William M. Fitzgerald. Decentralized semantic threat graphs. Lecture Notes in Computer Science, 7371:177–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31540-4_14/.

Sharon:2012:CTR

[433] Rami Sharon and Ehud Gudes. Code type revealing using experiments framework. Lecture Notes in Computer Science, 7371:193–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31540-4_15/.

Yakovets:2012:MDC

[434] Nikolay Yakovets, Jarek Gryz, Stephanie Hazlewood, and Paul van Run. From MDM to DB2: a case study of security enforcement migration. Lecture Notes in Computer Science, 7371:207–222, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31540-4_16/.

Sundareswaran:2012:XDH

[435] Smitha Sundareswaran and Anna Cinzia Squicciarini. XSS-Dec: a hybrid solution to mitigate cross-site scripting attacks. Lecture Notes in Computer Science, 7371:223–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31540-4_17/.

Shastry:2012:RSM

[436] Abhijith Shastry, Murat Kantarcioglu, Yan Zhou, and Bhavani Thuraisingham. Randomizing Smartphone malware profiles against statistical mining techniques. Lecture Notes in Computer Science, 7371:239– 254, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31540-4_18/.

Saljooghinejad:2012:LSA

[437] Hamed Saljooghinejad and Wilson Naik Bhukya. Layered security architecture for masquerade attack detection. Lecture Notes in Computer Science, 7371:255–262, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31540-4_19/.

Soodejani:2012:ABH

[438] Abbas Taheri Soodejani, Mohammad Ali Hadavi, and Rasool Jalili. k-anonymity-based horizontal fragmentation to preserve privacy in data outsourcing. Lecture Notes in Computer Science, 7371:263–273, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31540-4_20/.

Anonymous:2012:FMx

[439] Anonymous. Front matter. Lecture Notes in Computer Science, 7371: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31540-4/1.

Tosu:2012:OBM

[440] Kaori Tosu and Noboru Kunihiro. Optimal bounds for multi-prime Φ-hiding assumption. Lecture Notes in Computer Science, 7372:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31448-3_1/.

Fujioka:2012:SCE

[441] Atsushi Fujioka, Mark Manulis, Koutarou Suzuki, and Berkant Ustaoğlu. Sufficient condition for ephemeral key-leakage resilient tripartite key exchange. Lecture Notes in Computer Science, 7372:15–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31448-3_2/.

Higo:2012:GTP

[442] Haruna Higo, Keisuke Tanaka, Akihiro Yamada, and Kenji Yasunaga. A game-theoretic perspective on oblivious transfer. Lecture Notes in Computer Science, 7372:29–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31448-3_3/.

Nagashima:2012:FAS

[443] Yuji Nagashima and Noboru Kunihiro. Faster algorithm for solving hard knapsacks for moderate message length. Lecture Notes in Computer Science, 7372:43–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31448-3_4/.

Lory:2012:ASD

[444] Peter Lory and Manuel Liedel. Accelerating the secure distributed computation of the mean by a Chebyshev expansion. Lecture Notes in Computer Science, 7372:57–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31448-3_5/.

Isobe:2012:SAL

[445] Takanori Isobe and Kyoji Shibutani. Security analysis of the lightweight block ciphers XTEA, LED and Piccolo. Lecture Notes in Computer Science, 7372:71–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31448-3_6/.

Sasaki:2012:IKK

[446] Yu Sasaki, Sareh Emami, Deukjo Hong, and Ashish Kumar. Improved known-key distinguishers on Feistel–SP ciphers and application to Camellia. Lecture Notes in Computer Science, 7372:87–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31448-3_7/.

Chen:2012:LDC

[447] Jiazhe Chen and Leibo Li. Low data complexity attack on reduced Camellia-256. Lecture Notes in Computer Science, 7372:101–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31448-3_8/.

Meng:2012:CRS

[448] Xianmeng Meng and Xuexin Zheng. Cryptanalysis of RSA with a small parameter. Lecture Notes in Computer Science, 7372:115–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31448-3_9/.

Ding:2012:ABA

[449] Jintai Ding, Yanbin Pan, and Yingpu Deng. An algebraic broadcast attack against NTRU. Lecture Notes in Computer Science, 7372:124– 137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31448-3_10/.

ALMashrafi:2012:AIM

[450] Mufeed ALMashrafi, Harry Bartlett, Leonie Simpson, Ed Dawson, and Kenneth Koon-Ho Wong. Analysis of indirect message injection for MAC generation using stream ciphers. Lecture Notes in Computer Science, 7372:138–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31448-3_11/.

Fleischmann:2012:WDH

[451] Ewan Fleischmann, Christian Forler, Stefan Lucks, and Jakob Wenzel. Weimar-DM: a highly secure double-length compression function. Lecture Notes in Computer Science, 7372:152–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31448-3_12/.

Mathew:2012:EIC

[452] K. Preetha Mathew, Sachin Vasant, Sridhar Venkatesan, and C. Pandu Rangan. An efficient IND-CCA2 secure variant of the Niederreiter encryption scheme in the standard model. Lecture Notes in Computer Science, 7372:166–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31448-3_13/.

Morozov:2012:ZKP

[453] Kirill Morozov and Tsuyoshi Takagi. Zero-knowledge protocols for the McEliece encryption. Lecture Notes in Computer Science, 7372:180– 193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31448-3_14/.

Rangasamy:2012:ERP

[454] Jothi Rangasamy, Douglas Stebila, Colin Boyd, and Juan Manuel González-Nieto. Effort-release public-key encryption from cryptographic puzzles. Lecture Notes in Computer Science, 7372:194–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31448-3_15/.

Nguyen:2012:LRS

[455] Manh Ha Nguyen, Keisuke Tanaka, and Kenji Yasunaga. Leakageresilience of stateless/stateful public-key encryption from hash proofs. Lecture Notes in Computer Science, 7372:208–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31448-3_16/.

Peng:2012:HFT

[456] Kun Peng and Matt Henricksen. How to fix two RSA-based PVSS schemes— exploration and solution. Lecture Notes in Computer Science, 7372:223–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31448-3_17/.

Kurosawa:2012:RBV

[457] Kaoru Kurosawa, Ryo Nojima, and Le Trieu Phong. Relation between verifiable random functions and convertible undeniable signatures, and new constructions. Lecture Notes in Computer Science, 7372:235–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31448-3_18/.

Tan:2012:GFP

[458] Xiao Tan and Duncan S. Wong. Generalized first pre-image tractable random oracle model and signature schemes. Lecture Notes in Computer Science, 7372:247–260, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31448-3_19/.

Tian:2012:SND

[459] Haibo Tian, Xiaofeng Chen, and Jin Li. A short non-delegatable strong designated verifier signature. Lecture Notes in Computer Science, 7372:261–279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31448-3_20/.

Anonymous:2012:FMy

[460] Anonymous. Front matter. Lecture Notes in Computer Science, 7372: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31448-3/1.

Gupta:2012:EHF

[461] Anu Gupta and Subhrojyoti Sarkar. An efficient high frequency and low power analog multiplier in current domain. Lecture Notes in Computer Science, 7373:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31494-0_1/.

Maity:2012:DPP

[462] Biswajit Maity and Pradip Mandal. Design of push-pull dynamic leaker circuit for a low power embedded voltage regulator. Lecture Notes in Computer Science, 7373:10–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31494-0_2/.

Choudhury:2012:PMP

[463] Priyanka Choudhury and Sambhu Nath Pradhan. Power modeling of power gated FSM and its low power realization by simultaneous partitioning and state encoding using genetic algorithm. Lecture Notes in Computer Science, 7373:19–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31494-0_3/.

Shrestha:2012:DIL

[464] Rahul Shrestha and Roy Paily. Design and implementation of a linear feedback shift register interleaver for turbo decoding. Lecture Notes in Computer Science, 7373:30–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31494-0_4/.

Babu:2012:LCE

[465] Gunti Nagendra Babu, Brajesh Kumar Kaushik, Anand Bulusu, and Manoj Kumar Majumder. Low complexity encoder for crosstalk reduction in RLC modeled interconnects. Lecture Notes in Computer Science, 7373:40–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31494-0_5/.

Nandi:2012:APA

[466] Ashutosh Nandi, Ashok K. Saxena, and Sudeb Dasgupta. Analog performance analysis of dual-k spacer based underlap FinFET. Lecture Notes in Computer Science, 7373:46–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31494-0_6/.

Jayagowri:2012:IGT

[467] R. Jayagowri and K. S. Gurumurthy. Implementation of gating technique with modified scan flip-flop for low power testing of VLSI chips. Lecture Notes in Computer Science, 7373:52–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31494-0_7/.

Roy:2012:PBS

[468] Surajit Kumar Roy, Dona Roy, Chandan Giri, and Hafizur Rahaman. Post-bond stack testing for 3D stacked IC. Lecture Notes in Computer Science, 7373:59–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31494-0_8/.

Bandyopadhyay:2012:TVP

[469] Soumyadip Bandyopadhyay, Kunal Banerjee, Dipankar Sarkar, and Chittaranjan R. Mandal. Translation validation for PRES+ models of parallel behaviours via an FSMD equivalence checker. Lecture Notes in Computer Science, 7373:69–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31494-0_9/.

Saha:2012:DHS

[470] Prabir Saha, Arindam Banerjee, Anup Dandapat, and Partha Bhattacharyya. Design of high speed Vedic multiplier for decimal number system. Lecture Notes in Computer Science, 7373:79–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31494-0_10/.

Dalui:2012:ETD

[471] Mamata Dalui and Biplab K. Sikdar. An efficient test design for CMPs cache coherence realizing MESI protocol. Lecture Notes in Computer Science, 7373:89–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31494-0_11/.

Roy:2012:EHS

[472] Debapriya Basu Roy and Debdeep Mukhopadhyay. An efficient high speed implementation of flexible characteristic-2 multipliers on FP-GAs. Lecture Notes in Computer Science, 7373:99–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31494-0_12/.

Das:2012:AAT

[473] Subrata Das, Partha Sarathi Dasgupta, and Samar Sensarma. Arithmetic algorithms for ternary number system. Lecture Notes in Computer Science, 7373:111–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31494-0_13/.

Juneja:2012:SMB

[474] Dushyant Juneja, Sougata Kar, Procheta Chatterjee, and Siddhartha Sen. SOI MEMS based over-sampling accelerometer design with ΔΣ output. Lecture Notes in Computer Science, 7373:121–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31494-0_14/.

Rana:2012:DOW

[475] Goutam Rana, Samir Kumar Lahiri, and Chirasree Roy Chaudhuri. Design optimization of a wide band MEMS resonator for efficient energy harvesting. Lecture Notes in Computer Science, 7373:129–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31494-0_15/.

Kushwah:2012:ULP

[476] Chandrabhan Kushwah and Santosh K. Vishvakarma. Ultra-low power sub-threshold SRAM cell design to improve read static noise margin. Lecture Notes in Computer Science, 7373:139–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31494-0_16/.

Dobriyal:2012:WDP

[477] Arun Dobriyal, Rahul Gonnabattula, Pallab Dasgupta, and Chittaranjan R. Mandal. Workload driven power domain partitioning. Lecture Notes in Computer Science, 7373:147–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31494-0_17/.

Dasgupta:2012:INO

[478] Rituparna Dasgupta, Dipankar Saha, Jagannath Samanta, and Sayan Chatterjee. Implementation of a new offset generator block for the lowvoltage, low-power self biased threshold voltage extractor circuit. Lecture Notes in Computer Science, 7373:156–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31494-0_18/.

Hati:2012:HSL

[479] Manas Kumar Hati and Tarun Kanti Bhattacharyya. A high speed, low jitter and fast acquisition CMOS phase frequency detector for charge pump PLL. Lecture Notes in Computer Science, 7373:166–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31494-0_19/.

Tudu:2012:IBA

[480] Jaynarayan T. Tudu, Deepak Malani, and Virendra Singh. ILP based approach for input vector controlled (IVC) toggle maximization in combinational circuits. Lecture Notes in Computer Science, 7373:172–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31494-0_20/.

Anonymous:2012:FMz

[481] Anonymous. Front matter. Lecture Notes in Computer Science, 7373: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31494-0/1.

Karati:2012:BVE

[482] Sabyasachi Karati, Abhijit Das, Dipanwita Roychowdhury, and Bhargav Bellur. Batch verification of ECDSA signatures. Lecture Notes in Computer Science, 7374:1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31410-0_1/.

Alaoui:2012:ESA

[483] Sidi Mohamed El Yousfi Alaoui, Özgür Dagdelen, Pascal Véron, and David Galindo. Extended security arguments for signature schemes. Lecture Notes in Computer Science, 7374:19–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31410-0_2/.

Canard:2012:SSS

[484] Sébastien Canard, Amandine Jambert, and Roch Lescuyer. Sanitizable signatures with several signers and sanitizers. Lecture Notes in Computer Science, 7374:35–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31410-0_3/.

Wang:2012:ABD

[485] Jingjing Wang, Xiangxue Li, Kefei Chen, and Wenzheng Zhang. Attack based on direct sum decomposition against the nonlinear filter generator. Lecture Notes in Computer Science, 7374:53–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31410-0_4/.

Bringer:2012:FVM

[486] Julien Bringer, Hervé Chabanne, and Mélanie Favre. Fuzzy vault for multiple users. Lecture Notes in Computer Science, 7374:67–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31410-0_5/.

Safavi-Naini:2012:BCR

[487] Reihaneh Safavi-Naini and Mohammed Ashraful Alam Tuhin. Bounds and constructions for 1-round (0, δ)-secure message transmission against generalized adversary. Lecture Notes in Computer Science, 7374:82–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31410-0_6/.

Meziani:2012:IPS

[488] Mohammed Meziani, Gerhard Hoffmann, and Pierre-Louis Cayrel. Improving the performance of the SYND stream cipher. Lecture Notes in Computer Science, 7374:99–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31410-0_7/.

Chen:2012:IDC

[489] Jiazhe Chen, Meiqin Wang, and Bart Preneel. Impossible differential cryptanalysis of the lightweight block ciphers TEA, XTEA and HIGHT. Lecture Notes in Computer Science, 7374:117–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31410-0_8/.

Sasaki:2012:TSM

[490] Yu Sasaki, Lei Wang, Yasuhide Sakai, Kazuo Sakiyama, and Kazuo Ohta. Three-subset meet-in-the-middle attack on reduced XTEA. Lecture Notes in Computer Science, 7374:138–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31410-0_9/.

Sun:2012:DCR

[491] Yue Sun, Meiqin Wang, Shujia Jiang, and Qiumei Sun. Differential cryptanalysis of reduced-round ICEBERG. Lecture Notes in Computer Science, 7374:155–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31410-0_10/.

Eisenbarth:2012:CIP

[492] Thomas Eisenbarth, Zheng Gong, Tim Güneysu, Stefan Heyse, and Sebastiaan Indesteege. Compact implementation and performance evaluation of block ciphers in ATtiny devices. Lecture Notes in Computer Science, 7374:172–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31410-0_11/.

Thomae:2012:CET

[493] Enrico Thomae and Christopher Wolf. Cryptanalysis of enhanced TTS, STS and all its variants, or: Why cross-terms are important. Lecture Notes in Computer Science, 7374:188–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31410-0_12/.

Sarr:2012:CAY

[494] Augustin P. Sarr and Philippe Elbaz-Vincent. A complementary analysis of the (s)YZ and DIKE Protocols. Lecture Notes in Computer Science, 7374:203–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31410-0_13/.

Nitaj:2012:NAR

[495] Abderrahmane Nitaj. A new attack on RSA and CRT-RSA. Lecture Notes in Computer Science, 7374:221–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31410-0_14/.

Armknecht:2012:STH

[496] Frederik Armknecht, Stefan Katzenbeisser, and Andreas Peter. Shifttype homomorphic encryption and its application to fully homomorphic encryption. Lecture Notes in Computer Science, 7374:234–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31410-0_15/.

Fleischmann:2012:CSM

[497] Ewan Fleischmann, Christian Forler, and Stefan Lucks. The collision security of MDC-4. Lecture Notes in Computer Science, 7374:252–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31410-0_16/.

Choy:2012:SHI

[498] Jiali Choy, Huihui Yap, Khoongming Khoo, Jian Guo, Thomas Peyrin, and Axel Poschmann. SPN-hash: Improving the provable resistance against differential collision attacks. Lecture Notes in Computer Science, 7374:270–286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31410-0_17/.

Andreeva:2012:SAC

[499] Elena Andreeva, Bart Mennink, Bart Preneel, and Marjan Škrobot. Security analysis and comparison of the SHA-3 finalists BLAKE, Grøstl, JH, Keccak, and Skein. Lecture Notes in Computer Science, 7374:287– 305, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31410-0_18/.

Berger:2012:GFL

[500] Thierry P. Berger, Joffrey D’Hayer, Kevin Marquet, Marine Minier, and Gaël Thomas. The GLUON family: a lightweight hash function family based on FCSRs. Lecture Notes in Computer Science, 7374:306– 323, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31410-0_19/.

Schwabe:2012:SAP

[501] Peter Schwabe, Bo-Yin Yang, and Shang-Yi Yang. SHA-3 on ARM11 processors. Lecture Notes in Computer Science, 7374:324–341, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31410-0_20/.

Anonymous:2012:FMba

[502] Anonymous. Front matter. Lecture Notes in Computer Science, 7374: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31410-0/1.

Ayers:2012:CBC

[503] Joseph Ayers, Daniel Blustein, and Anthony Westphal. A conserved biomimetic control architecture for walking, swimming and flying robots. Lecture Notes in Computer Science, 7375:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31525-1_1/.

Bobo:2012:DNI

[504] Luis Bobo, Ivan Herreros, and Paul F. M. J. Verschure. A digital neuromorphic implementation of cerebellar associative learning. Lecture Notes in Computer Science, 7375:13–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31525-1_2/.

Bortoletto:2012:SEB

[505] Roberto Bortoletto, Massimo Sartori, Fuben He, and Enrico Pagello. Simulating an elastic bipedal robot based on musculoskeletal modeling. Lecture Notes in Computer Science, 7375:26–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31525-1_3/.

Boxerbaum:2012:SBC

[506] Alexander S. Boxerbaum, Kathryn A. Daltorio, Hillel J. Chiel, and Roger D. Quinn. A soft-body controller with ubiquitous sensor feedback. Lecture Notes in Computer Science, 7375:38–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31525-1_4/.

Boyer:2012:EOU

[507] Frédéric Boyer and Vincent Lebastard. Exploration of objects by an underwater robot with electric sense. Lecture Notes in Computer Science, 7375:50–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31525-1_5/.

Caluwaerts:2012:NIN

[508] Ken Caluwaerts, Antoine Favre-Félix, Mariacarla Staffa, Steve N’Guyen, and Christophe Grand. Neuro-inspired navigation strategies Shifting for robots: Integration of a multiple landmark taxon strategy. Lecture Notes in Computer Science, 7375:62–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31525-1_6/.

Carpi:2012:BTL

[509] Federico Carpi, Gabriele Frediani, and Danilo De Rossi. Bioinspired tunable lens driven by electroactive polymer artificial muscles. Lecture Notes in Computer Science, 7375:74–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31525-1_7/.

Chinellato:2012:PSS

[510] Eris Chinellato, Marco Antontelli, and Angel P. del Pobil. A pilot study on saccadic adaptation experiments with robots. Lecture Notes in Computer Science, 7375:83–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31525-1_8/.

Dastoor:2012:JRT

[511] Sanjay Dastoor, Sam Weiss, Hannah Stuart, and Mark Cutkosky. Jumping robot with a tunable suspension based on artificial muscles. Lecture Notes in Computer Science, 7375:95–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31525-1_9/.

Ejaz:2012:SVA

[512] Naveed Ejaz, Reiko J. Tanaka, and Holger G. Krapp. Static versus adaptive gain control strategy for visuo-motor stabilization. Lecture Notes in Computer Science, 7375:107–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31525-1_10/.

Hoinville:2012:LRM

[513] Thierry Hoinville, Rüdiger Wehner, and Holk Cruse. Learning and retrieval of memory elements in a navigation task. Lecture Notes in Computer Science, 7375:120–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31525-1_11/.

Landgraf:2012:IHD

[514] Tim Landgraf, Michael Oertel, Andreas Kirbach, Randolf Menzel, and Raúl Rojas. Imitation of the honeybee dance communication system by means of a biomimetic robot. Lecture Notes in Computer Science, 7375:132–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31525-1_12/.

Luvizotto:2012:FMR

[515] André Luvizotto, César Rennó-Costa, and Paul Verschure. A framework for mobile robot navigation using a temporal population code. Lecture Notes in Computer Science, 7375:144–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31525-1_13/.

Marcos:2012:GIM

[516] Encarni Marcos, Armin Duff, Martí Sánchez-Fibla, and Paul F. M. J. Verschure. Generalization of integrator models to foraging: a robot study using the DAC9 model. Lecture Notes in Computer Science, 7375:156– 167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31525-1_14/.

Mitchinson:2012:EAS

[517] Ben Mitchinson, Martin J. Pearson, Anthony G. Pipe, and Tony J. Prescott. The emergence of action sequences from spatial attention: Insight from rodent-like robots. Lecture Notes in Computer Science, 7375:168–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31525-1_15/.

Mohan:2012:HPE

[518] Vishwanathan Mohan and Pietro Morasso. How past experience, imitation and practice can be combined to swiftly learn to use novel “tools”: Insights from skill learning experiments with baby humanoids. Lecture Notes in Computer Science, 7375:180–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31525-1_16/.

Ognibene:2012:TCA

[519] Dimitri Ognibene, Eris Chinellato, Miguel Sarabia, and Yiannis Demiris. Towards contextual action recognition and target localization with active allocation of attention. Lecture Notes in Computer Science, 7375:192– 203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31525-1_17/.

Pavel:2012:RLI

[520] Ana Brânduşa Pavel, Cristian Ioan Vasile, and Ioan Dumitrache. Robot localization implemented with enzymatic numerical P systems. Lecture Notes in Computer Science, 7375:204–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31525-1_18/.

Philippides:2012:HCE

[521] Andrew Philippides, Bart Baddeley, Philip Husbands, and Paul Graham. How can embodiment simplify the problem of view-based navigation? Lecture Notes in Computer Science, 7375:216–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31525-1_19/.

Rabinovich:2012:DMC

[522] Mikhail I. Rabinovich and Pablo Varona. The dynamical modeling of cognitive robot-human centered interaction. Lecture Notes in Computer Science, 7375:228–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31525-1_20/.

Anonymous:2012:FMbb

[523] Anonymous. Front matter. Lecture Notes in Computer Science, 7375: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31525-1/1.

Turkov:2012:BAC

[524] Pavel Turkov, Olga Krasotkina, and Vadim Mottl. Bayesian approach to the concept drift in the pattern recognition problems. Lecture Notes in Computer Science, 7376:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31537-4_1/.

Ceci:2012:TRC

[525] Michelangelo Ceci, Annalisa Appice, Herna L. Viktor, Donato Malerba, and Eric Paquet. Transductive relational classification in the cotraining paradigm. Lecture Notes in Computer Science, 7376:11–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31537-4_2/.

Yang:2012:GNC

[526] Rong Yang and Zhenyuan Wang. Generalized nonlinear classification model based on cross-oriented Choquet integral. Lecture Notes in Computer Science, 7376:26–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31537-4_3/.

Nguyen:2012:GNS

[527] Hai Thanh Nguyen and Katrin Franke. A general L_{p}-norm support vector machine via mixed 0–1 programming. Lecture Notes in Computer Science, 7376:40–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31537-4_4/.

Kovacs:2012:RDC

[528] László Kovács. Reduction of distance computations in selection of pivot elements for balanced GHT structure. Lecture Notes in Computer Science, 7376:50–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31537-4_5/.

Joenssen:2012:HDM

[529] Dieter William Joenssen and Udo Bankhofer. Hot deck methods for imputing missing data. Lecture Notes in Computer Science, 7376:63– 75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31537-4_6/.

Bharambe:2012:B

[530] Saket Bharambe, Harshit Dubey, and Vikram Pudi. BINER. Lecture Notes in Computer Science, 7376:76–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31537-4_7/.

Mondal:2012:NAA

[531] Kartick Chandra Mondal, Nicolas Pasquier, Anirban Mukhopadhyay, and Ujjwal Maulik. A new approach for association rule mining and bi-clustering using formal concept analysis. Lecture Notes in Computer Science, 7376:86–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31537-4_8/.

Li:2012:TMA

[532] Aixiang Li, Makoto Haraguchi, and Yoshiaki Okubo. Top-N minimization approach for indicative correlation change mining. Lecture Notes in Computer Science, 7376:102–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31537-4_9/.

Leite:2012:SCA

[533] Rui Leite, Pavel Brazdil, and Joaquin Vanschoren. Selecting classification algorithms with active testing. Lecture Notes in Computer Science, 7376:117–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31537-4_10/.

Thombre:2012:CLR

[534] Amit Thombre. Comparing logistic regression, neural networks, C5.0 and M5’ classification techniques. Lecture Notes in Computer Science, 7376:132–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31537-4_11/.

Sapkota:2012:UGI

[535] Upendra Sapkota, Barrett R. Bryant, and Alan Sprague. Unsupervised grammar inference using the minimum description length principle. Lecture Notes in Computer Science, 7376:141–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31537-4_12/.

Oshiro:2012:HMT

[536] Thais Mayumi Oshiro, Pedro Santoro Perez, and José Augusto Baranauskas. How many trees in a random forest? Lecture Notes in Computer Science, 7376:154–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31537-4_13/.

Xu:2012:CTC

[537] Tao Xu, David Chiu, and Iker Gondra. Constructing target concept in multiple instance learning using maximum partial entropy. Lecture Notes in Computer Science, 7376:169–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31537-4_14/.

Bouhamed:2012:NLS

[538] Heni Bouhamed, Afif Masmoudi, Thierry Lecroq, and Ahmed Rebaï. A new learning structure heuristic of Bayesian networks from data. Lecture Notes in Computer Science, 7376:183–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31537-4_15/.

Pitelis:2012:DSL

[539] Nikolaos Pitelis and Anastasios Tefas. Discriminant subspace learning based on support vectors machines. Lecture Notes in Computer Science, 7376:198–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31537-4_16/.

Hoa:2012:NLS

[540] Nong Thi Hoa and Bui The Duy. A new learning strategy of general BAMs. Lecture Notes in Computer Science, 7376:213–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31537-4_17/.

Toussaint:2012:PGI

[541] Godfried T. Toussaint and Constantin Berzan. Proximity-graph instancebased learning, support vector machines, and high dimensionality: An empirical comparison. Lecture Notes in Computer Science, 7376:222– 236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31537-4_18/.

Ebrahimi:2012:SSC

[542] Javid Ebrahimi and Mohammad Saniee Abadeh. Semi supervised clustering: a Pareto approach. Lecture Notes in Computer Science, 7376:237– 251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31537-4_19/.

Silva:2012:SSC

[543] Andreia Silva and Cláudia Antunes. Semi-supervised clustering: a case study. Lecture Notes in Computer Science, 7376:252–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31537-4_20/.

Anonymous:2012:FMbc

[544] Anonymous. Front matter. Lecture Notes in Computer Science, 7376: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31537-4/1.

Schmidt:2012:ACA

[545] Rainer Schmidt, Heike Weiss, and Georg Fuellen. Application of classification algorithms on IDDM rat data. Lecture Notes in Computer Science, 7377:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31488-9_1/.

Bichindaritz:2012:RTC

[546] Isabelle Bichindaritz. Research themes in the case-based reasoning in health sciences core literature. Lecture Notes in Computer Science, 7377: 9–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31488-9_2/.

Kirshners:2012:RAD

[547] Arnis Kirshners, Serge Parshutin, and Marcis Leja. Research on application of data mining methods to diagnosing gastric cancer. Lecture Notes in Computer Science, 7377:24–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31488-9_3/.

Nagy:2012:SES

[548] Gabor I. Nagy and Krisztian Buza. SOHAC: Efficient storage of tick data that supports search and analysis. Lecture Notes in Computer Science, 7377:38–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31488-9_4/.

Liu:2012:ECT

[549] Hongyan Liu, Zhiyuan Yao, Tomas Eklund, and Barbro Back. Electricity consumption time series profiling: a data mining application in energy industry. Lecture Notes in Computer Science, 7377:52–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31488-9_5/.

Santos:2012:WTF

[550] Pedro Santos, Luisa F. Villa, Aníbal Reñones, Andrés Bustillo, and Jesús Maudes. Wind turbines fault diagnosis using ensemble classifiers. Lecture Notes in Computer Science, 7377:67–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31488-9_6/.

Moreira-Matias:2012:BBD

[551] Luís Moreira-Matias, Carlos Ferreira, João Gama, and João Mendes-Moreira. Bus bunching detection by mining sequences of headway deviations. Lecture Notes in Computer Science, 7377:77–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31488-9_7/.

Alves:2012:DAP

[552] Ronnie Alves, Pedro Ferreira, Joel Ribeiro, and Orlando Belo. Detecting abnormal patterns in call graphs based on the aggregation of relevant vertex measures. Lecture Notes in Computer Science, 7377:92– 102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31488-9_8/.

Ivannikov:2012:RTM

[553] Andriy Ivannikov, Mikko Jegoroff, and Tommi Kärkkäinen. Real-time mass flow estimation in circulating fluidized bed. Lecture Notes in Computer Science, 7377:103–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31488-9_9/.

Lejri:2012:RCB

[554] Ons Lejri and Moncef Tagina. Representation in case-based reasoning applied to control reconfiguration. Lecture Notes in Computer Science, 7377:113–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31488-9_10/.

Vanlaer:2012:IIO

[555] Jef Vanlaer, Pieter Van den Kerkhof, Geert Gins, and Jan F. M. Van Impe. The influence of input and output measurement noise on batch-end quality prediction with partial least squares. Lecture Notes in Computer Science, 7377:121–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31488-9_11/.

Shimada:2012:EAC

[556] Kaoru Shimada. An evolving associative classifier for incomplete database. Lecture Notes in Computer Science, 7377:136–150, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-31488-9_
12/.

Welcker:2012:ICP

[557] Laura Welcker, Stephan Koch, and Frank Dellmann. Improving classifier performance by knowledge-driven data preparation. Lecture Notes in Computer Science, 7377:151–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31488-9_13/.

Park:2012:CCC

[558] Eunkyoung Park, Younghee Kim, Ieejoon Kim, Jaeyeol Yoon, Jiyeon Lim, and Ungmo Kim. CWFM: Closed contingency weighted frequent itemsets mining. Lecture Notes in Computer Science, 7377:166–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31488-9_14/.

Bobrowski:2012:PMH

[559] Leon Bobrowski and Tomasz Łukaszuk. Prognostic modeling with high dimensional and censored data. Lecture Notes in Computer Science, 7377:178–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31488-9_15/.

Ahmed:2012:SSH

[560] Eya Ben Ahmed, Ahlem Nabli, and Faïez Gargouri. SHACUN: Semisupervised hierarchical active clustering based on ranking constraints.

Lecture Notes in Computer Science, 7377:194–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31488-9_16/.

Wang:2012:MST

[561] Xiaochun Wang, Xia Li Wang, and D. Mitch Wilkes. A minimum spanning tree-inspired clustering-based outlier detection technique. Lecture Notes in Computer Science, 7377:209–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31488-9_17/.

Elsayed:2012:MDN

[562] Samir A. Mohamed Elsayed, Sanguthevar Rajasekaran, and Reda A. Ammar. ML-DS: a novel deterministic sampling algorithm for association rules mining. Lecture Notes in Computer Science, 7377:224–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31488-9_18/.

Paja:2012:DRD

[563] Wiesław Paja and Mariusz Wrzesień. Decision rules development using set of generic operations approach. Lecture Notes in Computer Science, 7377:236–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31488-9_19/.

Klenk:2012:RDS

[564] Sebastian Klenk, Jürgen Dippon, Andre Burkovski, and Gunther Heidemann. Redundant dictionary spaces as a general concept for the analysis of non-vectorial data. Lecture Notes in Computer Science, 7377:243– 257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31488-9_20/.

Anonymous:2012:FMbd

[565] Anonymous. Front matter. Lecture Notes in Computer Science, 7377: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31488-9/1.

Clapes:2012:UIO

[566] Albert Clapés, Miguel Reyes, and Sergio Escalera. User identification and object recognition in clutter scenes based on RGB-depth analysis. Lecture Notes in Computer Science, 7378:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31567-1_1/.

Tejera:2012:CTV

[567] Margara Tejera and Adrian Hilton. Compression techniques for 3D video mesh sequences. Lecture Notes in Computer Science, 7378:12– 25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31567-1_2/.

Hauberg:2012:SMB

[568] Søren Hauberg and Kim Steenstrup Pedersen. Spatial measures between human poses for classification and understanding. Lecture Notes in Computer Science, 7378:26–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31567-1_3/.

Bakken:2012:RTP

[569] Rune Havnung Bakken and Adrian Hilton. Real-time pose estimation using constrained dynamics. Lecture Notes in Computer Science, 7378: 37–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31567-1_4/.

Ahn:2012:NEM

[570] Junghyun Ahn, Stéphane Gobron, and David Garcia. An NVC emotional model for conversational virtual humans in a 3D chatting environment. Lecture Notes in Computer Science, 7378:47–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31567-1_5/.

Gobron:2012:EBA

[571] Stéphane Gobron, Junghyun Ahn, and David Garcia. An event-based architecture to manage virtual human non-verbal communication in 3D chatting environment. Lecture Notes in Computer Science, 7378:58–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31567-1_6/.

Carretero:2012:IGC

[572] María del Puy Carretero and Aitor Ardanza. Improving gestural communication in virtual characters. Lecture Notes in Computer Science, 7378:

69–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31567-1_7/.

Navarro:2012:MVB

[573] Sergio Navarro and Adolfo López-Méndez. Multi-view body tracking with a detector-driven hierarchical particle filter. Lecture Notes in Computer Science, 7378:82–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31567-1_8/.

Kwolek:2012:RTM

[574] Bogdan Kwolek and Tomasz Krzeszowski. Real-time multi-view human motion tracking using particle swarm optimization with resampling. Lecture Notes in Computer Science, 7378:92–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31567-1_9/.

Neophytou:2012:CSS

[575] Alexandros Neophytou and Adrian Hilton. A comparative study of surface representations used in statistical human models. Lecture Notes in Computer Science, 7378:102–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31567-1_10/.

Xu:2012:CSP

[576] Ran Xu, Priyanshu Agarwal, and Suren Kumar. Combining skeletal pose with local motion for human activity recognition. Lecture Notes in Computer Science, 7378:114–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31567-1_11/.

Loconsole:2012:NML

[577] Claudio Loconsole and Nuno Barbosa. A new marker-less 3D kinectbased system for facial anthropometric measurements. Lecture Notes in Computer Science, 7378:124–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31567-1_12/.

Kitamura:2012:ADU

[578] Kensuke Kitamura, Noriko Takemura, and Yoshio Iwai. Automatic detection of unconscious reactions to illuminance changes in illumination. Lecture Notes in Computer Science, 7378:134–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31567-1_13/.

Bennett:2012:DMH

[579] Stuart Bennett and Joan Lasenby. Deformation measurement of a human chest experiencing global motion. Lecture Notes in Computer Science, 7378:144–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31567-1_14/.

Jensen:2012:AEM

[580] Thomas Jensen and Henrik Rasmussen. Automatic estimation of movement statistics of people. Lecture Notes in Computer Science, 7378:153– 162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31567-1_15/.

Holladay:2012:GMD

[581] Seth Holladay and Parris Egbert. Granular material deposition for simulation and texturing. Lecture Notes in Computer Science, 7378:163– 172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31567-1_16/.

Mahmoud:2012:FSM

[582] Nader Mahmoud and Stephane A. Nicolau. Fast 3D structure from motion with missing points from registration of partial reconstructions. Lecture Notes in Computer Science, 7378:173–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31567-1_17/.

Ju:2012:ERN

[583] Xiangyang Ju, Balvinder Khambay, and Emer O’Leary. Evaluation of the reproducibility of non-verbal facial animations. Lecture Notes in Computer Science, 7378:184–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31567-1_18/.

Ramos:2012:CLD

[584] Francisco Ramos, Oscar Ripolles, and Miguel Chover. Continuous level of detail for large scale rendering of 3D animated polygonal models. Lecture Notes in Computer Science, 7378:194–203, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31567-1_19/.

Gong:2012:NID

[585] Wenjuan Gong and Jordi Gonzàlez. A new image dataset on human interactions. Lecture Notes in Computer Science, 7378:204–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31567-1_20/.

Anonymous:2012:FMbe

[586] Anonymous. Front matter. Lecture Notes in Computer Science, 7378: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31567-1/1.

Berkovsky:2012:PNU

[587] Shlomo Berkovsky, Jill Freyne, and Gregory Smith. Personalized network updates: Increasing social interactions and contributions in social networks. Lecture Notes in Computer Science, 7379:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31454-4_1/.

Bohnert:2012:RSM

[588] Fabian Bohnert, Ingrid Zukerman, and David W. Albrecht. Realistic simulation of museum visitors’ movements as a tool for assessing sensor-based user models. Lecture Notes in Computer Science, 7379:14– 25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31454-4_2/.

Bohnert:2012:GPT

[589] Fabian Bohnert, Ingrid Zukerman, and Junaidy Laures. GECKOmmender: Personalised theme and tour recommendations for museums. Lecture Notes in Computer Science, 7379:26–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31454-4_3/.

Cena:2012:PBI

[590] Federica Cena, Silvia Likavec, and Francesco Osborne. Property-based interest propagation in ontology-based user model. Lecture Notes in Computer Science, 7379:38–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31454-4_4/.

Galan:2012:EEE

[591] Federico Cirett Galán and Carole R. Beal. EEG estimates of engagement and cognitive workload predict math problem solving outcomes. Lecture Notes in Computer Science, 7379:51–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31454-4_5/.

Desarkar:2012:PRB

[592] Maunendra Sankar Desarkar, Roopam Saxena, and Sudeshna Sarkar. Preference relation based matrix factorization for recommender systems. Lecture Notes in Computer Science, 7379:63–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31454-4_6/.

Finnson:2012:FMT

[593] John Finnson, Jie Zhang, Thomas Tran, Umar Farooq Minhas, and Robin Cohen. A framework for modeling trustworthiness of users in mobile vehicular ad-hoc networks and its validation through simulated traffic flow. Lecture Notes in Computer Science, 7379:76–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31454-4_7/.

Gao:2012:CSU

[594] Qi Gao, Fabian Abel, Geert-Jan Houben, and Yong Yu. A comparative study of users’ microblogging behavior on Sina Weibo and Twitter. Lecture Notes in Computer Science, 7379:88–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31454-4_8/.

Gong:2012:MMD

[595] Yue Gong, Joseph E. Beck, and Carolina Ruiz. Modeling multiple distributions of student performances to improve predictive accuracy. Lecture Notes in Computer Science, 7379:102–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31454-4_9/.

Guo:2012:SEM

[596] Guibing Guo, Jie Zhang, and Daniel Thalmann. A simple but effective method to incorporate trusted neighbors in recommender systems. Lecture Notes in Computer Science, 7379:114–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31454-4_10/.

Kardan:2012:EGD

[597] Samad Kardan and Cristina Conati. Exploring gaze data for determining user learning with an interactive simulation. Lecture Notes in Computer Science, 7379:126–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31454-4_11/.

Larizza:2012:SDU

[598] Melanie Larizza, Ingrid Zukerman, Fabian Bohnert, R. Andrew Russell, and Lucy Busija. Studies to determine user requirements regarding inhome monitoring systems. Lecture Notes in Computer Science, 7379:139– 150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31454-4_12/.

Leginus:2012:ITB

[599] Martin Leginus, Peter Dolog, and Valdas Žemaitis. Improving tensor based recommenders with clustering. Lecture Notes in Computer Science, 7379:151–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31454-4_13/.

Lehmann:2012:MUE

[600] Janette Lehmann, Mounia Lalmas, Elad Yom-Tov, and Georges Dupret. Models of user engagement. Lecture Notes in Computer Science, 7379:164–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31454-4_14/.

Mandl:2012:IPU

[601] Monika Mandl and Alexander Felfernig. Improving the performance of unit critiquing. Lecture Notes in Computer Science, 7379:176–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31454-4_15/.

Musto:2012:EST

[602] Cataldo Musto, Fedelucio Narducci, Pasquale Lops, Giovanni Semeraro, and Marco de Gemmis. Enhanced semantic TV-show representation for personalized electronic program guides. Lecture Notes in Computer Science, 7379:188–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31454-4_16/.

Navalpakkam:2012:ASO

[603] Vidhya Navalpakkam, Ravi Kumar, Lihong Li, and D. Sivakumar. Attention and selection in online choice tasks. Lecture Notes in Computer Science, 7379:200–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31454-4_17/.

Nunes:2012:IEJ

[604] Ingrid Nunes, Simon Miles, Michael Luck, and Carlos J. P. de Lucena. Investigating explanations to justify choice. Lecture Notes in Computer Science, 7379:212–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31454-4_18/.

Pizzato:2012:ESP

[605] Luiz Augusto Pizzato, Joshua Akehurst, Cameron Silvestrini, and Kalina Yacef. The effect of suspicious profiles on people recommenders. Lecture Notes in Computer Science, 7379:225–236, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31454-4_19/.

Said:2012:UNM

[606] Alan Said, Brijnesh J. Jain, Sascha Narr, and Till Plumbaum. Users and noise: The magic barrier of recommender systems. Lecture Notes in Computer Science, 7379:237–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31454-4_20/.

Anonymous:2012:FMbf

[607] Anonymous. Front matter. Lecture Notes in Computer Science, 7379: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31454-4/1.

Wei:2012:FBS

[608] Lei Wei, Alexei Sourin, and Zoran Najdovski. Function-based single and dual point haptic interaction in cyberworlds. Lecture Notes in Computer Science, 7380:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32663-9_1/.

Aleshin:2012:VPM

[609] Vladimir Aleshin and Valery Afanasiev. Visual 3D perception of motion environment and visibility factors in virtual space. Lecture Notes in Computer Science, 7380:17–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32663-9_2/.

Tilmanne:2012:CCS

[610] Joëlle Tilmanne and Thierry Dutoit. Continuous control of style and style transitions through linear interpolation in hidden Markov model based walk synthesis. Lecture Notes in Computer Science, 7380:34– 54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32663-9_3/.

Viswanathan:2012:MAH

[611] Vaisagh Viswanathan and Michael Lees. Modeling and analyzing the human cognitive limits for perception in crowd simulation. Lecture Notes in Computer Science, 7380:55–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32663-9_4/.

Yampolskiy:2012:EAA

[612] Roman V. Yampolskiy and Gyuchoon Cho. Experiments in artimetrics: Avatar face recognition. Lecture Notes in Computer Science, 7380:77– 94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32663-9_5/.

Ohmori:2012:VJU

[613] Kenji Ohmori and Tosiyasu L. Kunii. Visualization of joinery using homotopy theory and attaching maps. Lecture Notes in Computer Science, 7380:95–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32663-9_6/.

Kodama:2012:EBA

[614] Toshio Kodama, Yoichi Seki, and Tosiyasu L. Kunii. Example of business applications using the numerical value and exponential calculation of the cellular data system. Lecture Notes in Computer Science, 7380:115– 133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32663-9_7/.

Harrison:2012:AIS

[615] Erika Harrison and Ali Mahdavi-Amiri. Analysis of inverse Snyder optimizations. Lecture Notes in Computer Science, 7380:134–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32663-9_8/.

Li:2012:ECS

[616] Yichen Li, Mingqiang Wei, and Jianhuang Wu. Error-controllable simplification of point cloud. Lecture Notes in Computer Science, 7380:149– 162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32663-9_9/.

Wenzhen:2012:PBH

[617] Yang Wenzhen, Pan Zhigen, and Chen Wenhua. Physically-based haptic rendering for virtual hand interaction. Lecture Notes in Computer Science, 7380:163–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32663-9_10/.

Mura:2012:AOM

[618] Gianluca Mura. The advanced open metaplastic platform for cyber art. Lecture Notes in Computer Science, 7380:179–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32663-9_11/.

Anonymous:2012:BMh

[619] Anonymous. Back matter. Lecture Notes in Computer Science, 7380: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32663-9/1.

Anonymous:2012:FMbg

[620] Anonymous. Front matter. Lecture Notes in Computer Science, 7380: I–XII, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32663-9/1.

Gao:2012:MSY

[621] Yuan Gao and Kai Salomaa. In memoriam sheng Yu. Lecture Notes in Computer Science, 7381:1–4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31606-7_1/.

Brzozowski:2012:SMC

[622] Janusz Brzozowski. In search of most complex regular languages. Lecture Notes in Computer Science, 7381:5–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31606-7_2/.

Ehrenfeucht:2012:FFP

[623] Andrzej Ehrenfeucht and Grzegorz Rozenberg. A formal framework for processes inspired by the functioning of living cells. Lecture Notes in Computer Science, 7381:25–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31606-7_3/.

Gastin:2012:APW

[624] Paul Gastin and Benjamin Monmege. Adding pebbles to weighted automata. Lecture Notes in Computer Science, 7381:28–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31606-7_4/.

Oliveira:2012:TLA

[625] José N. Oliveira. Typed linear algebra for weigthed (probabilistic) automata. Lecture Notes in Computer Science, 7381:52–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31606-7_5/.

Allauzen:2012:PTE

[626] Cyril Allauzen and Michael Riley. A pushdown transducer extension for the OpenFst library. Lecture Notes in Computer Science, 7381:66– 77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31606-7_6/.

Amavi:2012:WIR

[627] Joshua Amavi, Jacques Chabin, and Pierre Réty. Weak inclusion for recursive XML types. Lecture Notes in Computer Science, 7381:78–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31606-7_7/.

Berlinkov:2012:SAQ

[628] Mikhail V. Berlinkov. Synchronizing automata on quasi-Eulerian digraph. Lecture Notes in Computer Science, 7381:90–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31606-7_8/.

Ceccherini-Silberstein:2012:CAR

[629] Tullio Ceccherini-Silberstein, Michel Coornaert, Francesca Fiorenzi, and Zoran Šunić. Cellular automata on regular rooted trees. Lecture Notes in Computer Science, 7381:101–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31606-7_9/.

Reghizzi:2012:SLT

[630] Stefano Crespi Reghizzi and Pierluigi L. San Pietro. Strict local testability with consensus equals regularity. Lecture Notes in Computer Science, 7381:113–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31606-7_10/.

Degano:2012:NAR

[631] Pierpaolo Degano, Gian-Luigi Ferrari, and Gianluca Mezzetti. Nominal automata for resource usage control. Lecture Notes in Computer Science, 7381:125–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31606-7_11/.

Droste:2012:WNW

[632] Manfred Droste and Bundit Pibaljommee. Weighted nested word automata and logics over strong bimonoids. Lecture Notes in Computer Science, 7381:138–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31606-7_12/.

Faro:2012:FSA

[633] Simone Faro and Thierry Lecroq. A fast suffix automata based algorithm for exact online string matching. Lecture Notes in Computer Science, 7381:149–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31606-7_13/.

Fominykh:2012:PAS

[634] Fedor Fominykh and Mikhail Volkov. P(l)aying for synchronization. Lecture Notes in Computer Science, 7381:159–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31606-7_14/.

Gusev:2012:SAB

[635] Vladimir V. Gusev. Synchronizing automata of bounded rank. Lecture Notes in Computer Science, 7381:171–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31606-7_15/.

Goc:2012:ATP

[636] Daniel Goč, Dane Henshall, and Jeffrey Shallit. Automatic theoremproving in combinatorics on words. Lecture Notes in Computer Science, 7381:180–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31606-7_16/.

Ibarra:2012:HSH

[637] Oscar H. Ibarra and Nicholas Q. Tran. How to synchronize the heads of a multitape automaton. Lecture Notes in Computer Science, 7381:192– 204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31606-7_17/.

Jahn:2012:RIL

[638] Franz Jahn, Manfred Kufleitner, and Alexander Lauser. Regular ideal languages and their Boolean combinations. Lecture Notes in Computer Science, 7381:205–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31606-7_18/.

Jez:2012:HMD

[639] Artur Jeż and Andreas Maletti. Hyper-minimization for deterministic tree automata. Lecture Notes in Computer Science, 7381:217–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31606-7_19/.

Jiraskova:2012:SCC

[640] Galina Jirásková and Tomáš Masopust. On the state and computational complexity of the reverse of acyclic minimal DFAs. Lecture Notes in Computer Science, 7381:229–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31606-7_20/.

Anonymous:2012:FMbh

[641] Anonymous. Front matter. Lecture Notes in Computer Science, 7381: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31606-7/1.

Chen:2012:TVS

[642] Shizhi Chen, D. Michael Quintian, and YingLi Tian. Towards a visual speech learning system for the deaf by matching dynamic lip shapes. Lecture Notes in Computer Science, 7382:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31522-0_1/.

Kobayashi:2012:TSS

[643] Makoto Kobayashi, Takuya Suzuki, and Daisuke Wakatsuki. Teaching support software for hearing impaired students who study computer operation. Lecture Notes in Computer Science, 7382:10–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31522-0_2/.

Hladik:2012:HBO

[644] Petr Hladík and Tomáš Gůra. The hybrid book — one document for all in the latest development. Lecture Notes in Computer Science, 7382:18– 24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31522-0_3/.

Petz:2012:DCS

[645] Andrea Petz and Klaus Miesenberger. Dealing with changes in supporting students with disabilities in higher education. Lecture Notes in Computer Science, 7382:25–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31522-0_4/.

Featherstone:2012:PDS

[646] Lisa Featherstone and Simon Ball. Putting the disabled student in charge: Introduction to the special thematic session. Lecture Notes in Computer Science, 7382:33–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31522-0_5/.

Christensen:2012:BDL

[647] Lars Ballieu Christensen and Tanja Stevns. Biblus — a digital library to support integration of visually impaired in mainstream education. Lecture Notes in Computer Science, 7382:36–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31522-0_6/.

McNaught:2012:AAA

[648] Alistair McNaught and Lisa Featherstone. Alternative approaches to alternative formats — changing expectations by challenging myths. Lecture Notes in Computer Science, 7382:43–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31522-0_7/.

Wald:2012:ATE

[649] Mike Wald, E. A. Draffan, Russell Newman, Sebastian Skuse, and Chris Phethean. Access toolkit for education. Lecture Notes in Computer Science, 7382:51–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31522-0_8/.

Kushalnagar:2012:CBP

[650] Poorna Kushalnagar, Benjamin Williams, and Raja S. Kushalnagar. Community-based participatory approach: Students as partners in educational accessible technology research. Lecture Notes in Computer Science, 7382:59–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31522-0_9/.

Canton:2012:ANI

[651] Paloma Cantón, Ángel L. González, Gonzalo Mariscal, and Carlos Ruiz. Applying new interaction paradigms to the education of children with special educational needs. Lecture Notes in Computer Science, 7382:65– 72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31522-0_10/.

Faux:2012:IVD

[652] Fern Faux, David Finch, and Lisa Featherstone. InStep: a video database assessment tool. Lecture Notes in Computer Science, 7382:73– 76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31522-0_11/.

Christensen:2012:SMI

[653] Lars Ballieu Christensen, Sean J. Keegan, and Tanja Stevns. SCRIBE: a model for implementing Robobraille in a higher education institution. Lecture Notes in Computer Science, 7382:77–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31522-0_12/.

Kohlmann:2012:IBC

[654] Wiebke Köhlmann. Identifying barriers to collaborative learning for the blind. Lecture Notes in Computer Science, 7382:84–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31522-0_13/.

Kushalnagar:2012:DHS

[655] Raja S. Kushalnagar, Poorna Kushalnagar, and Jeffrey B. Pelz. Deaf and hearing students’ eye gaze collaboration. Lecture Notes in Computer Science, 7382:92–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31522-0_14/.

Borges:2012:MPE

[656] José Antonio Borges and Dolores Tomé. The Musibraille Project — enabling the inclusion of blind students in music courses. Lecture Notes in Computer Science, 7382:100–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31522-0_15/.

Wald:2012:INE

[657] Mike Wald. Important new enhancements to inclusive learning using recorded lectures. Lecture Notes in Computer Science, 7382:108–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31522-0_16/.

Fujiyoshi:2012:DNA

[658] Mamoru Fujiyoshi, Akio Fujiyoshi, Akiko Ohsawa, Toshiaki Aomatsu, and Haruhiko Sawazaki. Development of new auditory testing media with invisible 2-dimensional codes for test-takers with print disabilities. Lecture Notes in Computer Science, 7382:116–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31522-0_17/.

Gardner:2012:MAM

[659] John Gardner and Courtney Christensen. More accessible math. Lecture Notes in Computer Science, 7382:124–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31522-0_18/.

Yamaguchi:2012:AAT

[660] Katsuhito Yamaguchi and Masakazu Suzuki. Accessible authoring tool for DAISY ranging from mathematics to others. Lecture Notes in Computer Science, 7382:130–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31522-0_19/.

Gonzurova:2012:BFL

[661] Wanda Gonzúrová and Pavel Hrabák. Blind friendly LaTeX. Lecture Notes in Computer Science, 7382:138–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31522-0_20/.

Anonymous:2012:FMbi

[662] Anonymous. Front matter. Lecture Notes in Computer Science, 7382: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31522-0/1.

Knudsen:2012:MAA

[663] Lars Emil Knudsen and Harald Holone. A multimodal approach to accessible Web content on Smartphones. Lecture Notes in Computer Science, 7383:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31534-3_1/.

Manduchi:2012:MVA

[664] Roberto Manduchi. Mobile vision as assistive technology for the blind: An experimental study. Lecture Notes in Computer Science, 7383:9– 16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31534-3_2/.

Wang:2012:CBS

[665] Shuihua Wang and Yingli Tian. Camera-based signage detection and recognition for blind persons. Lecture Notes in Computer Science, 7383: 17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31534-3_3/.

Coughlan:2012:CTI

[666] James M. Coughlan and Huiying Shen. The crosswatch traffic intersection analyzer: a roadmap for the future. Lecture Notes in Computer Science, 7383:25–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31534-3_4/.

Zegarra:2012:GIM

[667] Jesus Zegarra and René Farcy. GPS and inertial measurement unit (IMU) as a navigation system for the visually impaired. Lecture Notes in Computer Science, 7383:29–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31534-3_5/.

Molina:2012:VNI

[668] Edgardo Molina, Zhigang Zhu, and Yingli Tian. Visual nouns for indoor/outdoor navigation. Lecture Notes in Computer Science, 7383:33– 40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31534-3_6/.

Shen:2012:TRT

[669] Huiying Shen and James M. Coughlan. Towards a real-time system for finding and reading signs for visually impaired users. Lecture Notes in Computer Science, 7383:41–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31534-3_7/.

Choi:2012:URC

[670] Yoonjung Choi and Ki-Hyung Hong. User requirements for camera-based mobile applications on touch screen devices for blind people. Lecture Notes in Computer Science, 7383:48–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31534-3_8/.

Boulares:2012:RPI

[671] Mehrez Boulares and Mohamed Jemni. A route planner interpretation service for hard of hearing people. Lecture Notes in Computer Science, 7383:52–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31534-3_9/.

Spindler:2012:TFP

[672] Martin Spindler, Michael Weber, Denise Prescher, Mei Miao, and Gerhard Weber. Translating floor plans into directions. Lecture Notes in Computer Science, 7383:59–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31534-3_10/.

Mehigan:2012:HWT

[673] Tracey J. Mehigan and Ian Pitt. Harnessing wireless technologies for campus navigation by blind students and visitors. Lecture Notes in Computer Science, 7383:67–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31534-3_11/.

Kulyukin:2012:ESB

[674] Vladimir Kulyukin, Tanwir Zaman, Abhishek Andhavarapu, and Aliasgar Kutiyanawala. Eyesight sharing in blind grocery shopping: Remote P2P caregiving through cloud computing. Lecture Notes in Computer Science, 7383:75–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31534-3_12/.

Almer:2012:ATF

[675] Stefan Almer, Josef Kolbitsch, Johannes Oberzaucher, and Martin Ebner. Assessment test framework for collecting and evaluating fallrelated data using mobile devices. Lecture Notes in Computer Science, 7383:83–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31534-3_13/.

Bischof:2012:NWC

[676] Werner Bischof, Elmar Krajnc, Markus Dornhofer, and Michael Ulm. NAVCOM–WLAN communication between public transport vehicles and smart phones to support visually impaired and blind people. Lecture Notes in Computer Science, 7383:91–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31534-3_14/.

Miyoshi:2012:MTR

[677] Shigeki Miyoshi, Sumihiro Kawano, Mayumi Shirasawa, Kyoko Isoda, and Michiko Hasuike. Mobile-type remote captioning system for deaf or hard-of-hearing people and the experience of remote supports after the great east Japan earthquake. Lecture Notes in Computer Science, 7383:99–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31534-3_15/.

Zamfir:2012:HAO

[678] Bogdan Zamfir, Robert Tedesco, and Brian Reichow. Handheld “app” offering visual support to students with autism spectrum disorders (ASDs). Lecture Notes in Computer Science, 7383:105–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31534-3_16/.

Bumbalek:2012:CBA

[679] Zdenek Bumbalek, Jan Zelenka, and Lukas Kencl. Cloud-based assistive speech-transcription services. Lecture Notes in Computer Science, 7383:113–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31534-3_17/.

Hwang:2012:DVU

[680] Yumi Hwang, Daejin Shin, Chang-Yeal Yang, Seung-Yeun Lee, Jin Kim, and Byunggoo Kong. Developing a voice user interface with improved usability for people with dysarthria. Lecture Notes in Computer Science, 7383:117–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31534-3_18/.

Palmer:2012:WRV

[681] Frank G. Palmer, Zhigang Zhu, and Tony Ro. Wearable rangevibrotactile field: Design and evaluation. Lecture Notes in Computer Science, 7383:125–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31534-3_19/.

Kupryjanow:2012:SSS

[682] Adam Kupryjanow, Piotr Suchomski, Piotr Odya, and Andrzej Czyzewski. System supporting speech perception in special educational needs schoolchildren. Lecture Notes in Computer Science, 7383:133– 136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31534-3_20/.

Anonymous:2012:FMbj

[683] Anonymous. Front matter. Lecture Notes in Computer Science, 7383: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31534-3/1.

Castelluccia:2012:BYA

[684] Claude Castelluccia, Mohamed-Ali Kaafar, and Minh-Dung Tran. Betrayed by your ads! Lecture Notes in Computer Science, 7384:1–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31680-7_1/.

Danezis:2012:PCS

[685] George Danezis, Markulf Kohlweiss, Benjamin Livshits, and Alfredo Rial. Private client-side profiling with random forests and hidden Markov models. Lecture Notes in Computer Science, 7384:18–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31680-7_2/.

Perez-Gonzalez:2012:USD

[686] Fernando Pérez-González and Carmela Troncoso. Understanding statistical disclosure: a least squares approach. Lecture Notes in Computer Science, 7384:38–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31680-7_3/.

Gong:2012:WDU

[687] Xun Gong, Nikita Borisov, Negar Kiyavash, and Nabil Schear. Website detection using remote traffic analysis. Lecture Notes in Computer Science, 7384:58–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31680-7_4/.

Liu:2012:ITP

[688] Wen Ming Liu, Lingyu Wang, Kui Ren, Pengsu Cheng, and Mourad Debbabi. k-indistinguishable traffic padding in Web applications. Lecture Notes in Computer Science, 7384:79–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31680-7_5/.

Gilad:2012:SDT

[689] Yossi Gilad and Amir Herzberg. Spying in the dark: TCP and Tor traffic analysis. Lecture Notes in Computer Science, 7384:100–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31680-7_6/.

Alhadidi:2012:SDF

[690] Dima Alhadidi, Noman Mohammed, Benjamin C. M. Fung, and Mourad Debbabi. Secure distributed framework for achieving ϵ-differential privacy. Lecture Notes in Computer Science, 7384:120–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31680-7_7/.

Chan:2012:DPC

[691] T.-H. Hubert Chan, Mingfei Li, Elaine Shi, and Wenchang Xu. Differentially private continual monitoring of heavy hitters from distributed streams. Lecture Notes in Computer Science, 7384:140–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31680-7_8/.

Fang:2012:ADP

[692] Chengfang Fang and Ee-Chien Chang. Adaptive differentially private histogram of low-dimensional data. Lecture Notes in Computer Science, 7384:160–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31680-7_9/.

Blass:2012:PPP

[693] Erik-Oliver Blass, Roberto Di Pietro, Refik Molva, and Melek Önen. PRISM — privacy-preserving search in MapReduce. Lecture Notes in Computer Science, 7384:180–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31680-7_10/.

Pirker:2012:PPP

[694] Martin Pirker, Daniel Slamanig, and Johannes Winter. Practical privacy preserving cloud resource-payment for constrained clients. Lecture Notes in Computer Science, 7384:201–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31680-7_11/.

Jawurek:2012:FTP

[695] Marek Jawurek and Florian Kerschbaum. Fault-tolerant privacypreserving statistics. Lecture Notes in Computer Science, 7384:221–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31680-7_12/.

Fifield:2012:ECB

[696] David Fifield, Nate Hardison, Jonathan Ellithorpe, Emily Stark, and Dan Boneh. Evading censorship with browser-based proxies. Lecture Notes in Computer Science, 7384:239–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31680-7_13/.

Nikiforakis:2012:EER

[697] Nick Nikiforakis, Steven Van Acker, Frank Piessens, and Wouter Joosen. Exploring the ecosystem of referrer-anonymizing services. Lecture Notes in Computer Science, 7384:259–278, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31680-7_14/.

Garg:2012:RCD

[698] Vaibhav Garg, L. Jean Camp, Katherine Connelly, and Lesa Lorenzen-Huber. Risk communication design: Video vs. text. Lecture Notes in Computer Science, 7384:279–298, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31680-7_15/.

McDonald:2012:UFI

[699] Andrew W. E. McDonald, Sadia Afroz, Aylin Caliskan, Ariel Stolerman, and Rachel Greenstadt. Use fewer instances of the letter “i”: Toward writing style anonymization. Lecture Notes in Computer Science, 7384:299–318, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31680-7_16/.

Anonymous:2012:BMi

[700] Anonymous. Back matter. Lecture Notes in Computer Science, 7384: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31680-7/1.

Anonymous:2012:FMbk

[701] Anonymous. Front matter. Lecture Notes in Computer Science, 7384: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31680-7/1.

Ball:2012:BFO

[702] Thomas Ball, Nikolaj Bjørner, Leonardo de Moura, Kenneth L. McMillan, and Margus Veanes. Beyond first-order satisfaction: Fixed points, interpolants, automata and polynomials. Lecture Notes in Computer Science, 7385:1–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31759-0_1/.

Marinescu:2012:HCS

[703] Paul Dan Marinescu and Cristian Cadar. High-coverage symbolic patch testing. Lecture Notes in Computer Science, 7385:7–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31759-0_2/.

Rybalchenko:2012:TAS

[704] Andrey Rybalchenko. Towards automatic synthesis of software verification tools. Lecture Notes in Computer Science, 7385:22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-31759-0_3.

Zeller:2012:MM

[705] Andreas Zeller. Mining models. Lecture Notes in Computer Science, 7385:23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-31759-0_4.

Leue:2012:CEA

[706] Stefan Leue and Mitra Tabaei Befrouei. Counterexample explanation by anomaly detection. Lecture Notes in Computer Science, 7385:24– 42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31759-0_5/.

Evangelista:2012:CSL

[707] Sami Evangelista and Lars Michael Kristensen. Combining the sweepline method with the use of an external-memory priority queue. Lecture Notes in Computer Science, 7385:43–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31759-0_6/.

Zheng:2012:CMA

[708] Hao Zheng, Emmanuel Rodriguez, Yingying Zhang, and Chris Myers. A compositional minimization approach for large asynchronous design verification. Lecture Notes in Computer Science, 7385:62–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31759-0_7/.

Ditter:2012:PSV

[709] Alexander Ditter, Milan Češka, and Gerald Lüttgen. On parallel software verification using Boolean equation systems. Lecture Notes in Computer Science, 7385:80–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31759-0_8/.

Wijs:2012:IGS

[710] Anton J. Wijs and Dragan Bošnački. Improving GPU sparse matrixvector multiplication for probabilistic model checking. Lecture Notes in Computer Science, 7385:98–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31759-0_9/.

Ruys:2012:GGA

[711] Theo C. Ruys and Pim Kars. Gossiping girls are all alike. Lecture Notes in Computer Science, 7385:117–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31759-0_10/.

Garlapati:2012:FVH

[712] Shravan Garlapati and Sandeep K. Shukla. Formal verification of hierarchically distributed agent based protection scheme in smart grid. Lecture Notes in Computer Science, 7385:137–154, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31759-0_11/.

Holzmann:2012:PSM

[713] Gerard J. Holzmann. Parallelizing the spin model checker. Lecture Notes in Computer Science, 7385:155–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31759-0_12/.

Burns:2012:PMC

[714] Ethan Burns and Rong Zhou. Parallel model checking using abstraction. Lecture Notes in Computer Science, 7385:172–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31759-0_13/.

Zeng:2012:MPA

[715] Reng Zeng, Zhuo Sun, Su Liu, and Xudong He. McPatom: a predictive analysis tool for atomicity violation using model checking. Lecture Notes in Computer Science, 7385:191–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31759-0_14/.

Sethi:2012:PMC

[716] Divjyot Sethi, Muralidhar Talupur, Daniel Schwartz-Narbonne, and Sharad Malik. Parameterized model checking of fine grained concurrency. Lecture Notes in Computer Science, 7385:208–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31759-0_15/.

Bogdoll:2012:MBG

[717] Jonathan Bogdoll, Alexandre David, Arnd Hartmanns, and Holger Hermanns. mctau: Bridging the gap between Modest and UPPAAL. Lecture Notes in Computer Science, 7385:227–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31759-0_16/.

Riener:2012:FFF

[718] Heinz Riener and Görschwin Fey. FAuST: a framework for formal verification, automated debugging, and software test generation. Lecture Notes in Computer Science, 7385:234–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31759-0_17/.

Sulzmann:2012:MCD

[719] Martin Sulzmann and Axel Zechner. Model checking DSL-generated C source code. Lecture Notes in Computer Science, 7385:241–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31759-0_18/.

Christ:2012:SIS

[720] Jürgen Christ, Jochen Hoenicke, and Alexander Nutz. SMTInterpol: An interpolating SMT solver. Lecture Notes in Computer Science, 7385:248– 254, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31759-0_19/.

Jiang:2012:SMT

[721] Yong Jiang and Zongyan Qiu. S2N: Model transformation from SPIN to NuSMV. Lecture Notes in Computer Science, 7385:255–260, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31759-0_20/.

Anonymous:2012:BMj

[722] Anonymous. Back matter. Lecture Notes in Computer Science, 7385: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31759-0/1.

Anonymous:2012:FMbl

[723] Anonymous. Front matter. Lecture Notes in Computer Science, 7385: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31759-0/1.

Jeannin:2012:CC

[724] Jean-Baptiste Jeannin and Dexter Kozen. Computing with capsules. Lecture Notes in Computer Science, 7386:1–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31623-4_1/.

Kapoutsis:2012:M

[725] Christos A. Kapoutsis. Minicomplexity. Lecture Notes in Computer Science, 7386:20–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31623-4_2/.

Platzer:2012:LAH

[726] André Platzer. Logical analysis of hybrid systems. Lecture Notes in Computer Science, 7386:43–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31623-4_3/.

Silva:2012:GAP

[727] Pedro V. Silva. Groups and automata: a perfect match. Lecture Notes in Computer Science, 7386:50–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31623-4_4/.

Arroyo:2012:UDP

[728] Fernando Arroyo, Juan Castellanos, and Victor Mitrana. Uniform distributed pushdown automata systems. Lecture Notes in Computer Science, 7386:64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31623-4_5/.

Bednarova:2012:RNC

[729] Zuzana Bednárová, Viliam Geffert, Carlo Mereghetti, and Beatrice Palano. Removing nondeterminism in constant height pushdown automata. Lecture Notes in Computer Science, 7386:76–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31623-4_6/.

Bianchi:2012:IOT

[730] Maria Paola Bianchi, Markus Holzer, Sebastian Jakobi, and Giovanni Pighizzini. On inverse operations and their descriptional complexity. Lecture Notes in Computer Science, 7386:89–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31623-4_7/.

Blanchet-Sadri:2012:DRS

[731] Francine Blanchet-Sadri and Sean Simmons. Deciding representability of sets of words of equal length. Lecture Notes in Computer Science, 7386:103–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31623-4_8/.

Brzozowski:2012:SCS

[732] Janusz Brzozowski and Baiyu Li. Syntactic complexities of some classes of star-free languages. Lecture Notes in Computer Science, 7386:117– 129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31623-4_9/.

Campeanu:2012:RBB

[733] Cezar Câmpeanu. Randomness behaviour in Blum universal static complexity spaces. Lecture Notes in Computer Science, 7386:130–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31623-4_10/.

Dassow:2012:PCS

[734] Jürgen Dassow and Ronny Harbich. Production complexity of some operations on context-free languages. Lecture Notes in Computer Science, 7386:141–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31623-4_11/.

Gao:2012:SCS

[735] Yuan Gao and Lila Kari. State complexity of star and square of union of k regular languages. Lecture Notes in Computer Science, 7386:155– 168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31623-4_12/.

Holzer:2012:SCC

[736] Markus Holzer and Sebastian Jakobi. State complexity of chop operations on unary and finite languages. Lecture Notes in Computer Science, 7386:169–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31623-4_13/.

Jez:2012:NNS

[737] Artur Jeż and Alexander Okhotin. On the number of nonterminal symbols in unambiguous conjunctive grammars. Lecture Notes in Computer Science, 7386:183–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31623-4_14/.

Jiraskova:2012:DCB

[738] Galina Jirásková and Ondřej Klíma. Descriptional complexity of biautomata. Lecture Notes in Computer Science, 7386:196–208, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-31623-4_
15/.

Malcher:2012:DCP

[739] Andreas Malcher, Katja Meckel, Carlo Mereghetti, and Beatrice Palano. Descriptional complexity of pushdown store languages. Lecture Notes in Computer Science, 7386:209–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31623-4_16/.

Manea:2012:ICG

[740] Florin Manea and Bianca Truthe. On internal contextual grammars with subregular selection languages. Lecture Notes in Computer Science, 7386:222–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31623-4_17/.

Meduna:2012:IHL

[741] Alexander Meduna, Lukáš Vrábel, and Petr Zemek. An infinite hierarchy of language families resulting from stateless pushdown automata with limited pushdown alphabets. Lecture Notes in Computer Science, 7386:236–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31623-4_18/.

Otto:2012:CPS

[742] Friedrich Otto. Centralized PC systems of pushdown automata versus multi-head pushdown automata. Lecture Notes in Computer Science, 7386:244–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31623-4_19/.

Palioudakis:2012:SCL

[743] Alexandros Palioudakis, Kai Salomaa, and Selim G. Akl. State complexity and limited nondeterminism. Lecture Notes in Computer Science, 7386:252–265, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31623-4_20/.

Anonymous:2012:FMbm

[744] Anonymous. Front matter. Lecture Notes in Computer Science, 7386: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31623-4/1.

Heinrich:2012:RAW

[745] Matthias Heinrich, Franz Josef Grüneberger, Thomas Springer, and Martin Gaedke. Reusable awareness widgets for collaborative Web applications — a non-invasive approach. Lecture Notes in Computer Science, 7387:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31753-8_1/.

Xiao:2012:NTO

[746] Feng Xiao, Tomoya Noro, and Takehiro Tokuda. News-topic oriented hashtag recommendation in Twitter based on characteristic cooccurrence word detection. Lecture Notes in Computer Science, 7387: 16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31753-8_2/.

Nebeling:2012:CWE

[747] Michael Nebeling, Stefania Leone, and Moira C. Norrie. Crowdsourced Web engineering and design. Lecture Notes in Computer Science, 7387: 31–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31753-8_3/.

Vandic:2012:SPW

[748] Damir Vandic, Flavius Frasincar, and Frederik Hogenboom. Scaling pair-wise similarity-based algorithms in tagging spaces. Lecture Notes in Computer Science, 7387:46–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31753-8_4/.

Leginus:2012:MIT

[749] Martin Leginus, Peter Dolog, Ricardo Lage, and Frederico Durao. Methodologies for improved tag cloud generation with clustering. Lecture Notes in Computer Science, 7387:61–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31753-8_5/.

Bianchini:2012:SCT

[750] Devis Bianchini, Valeria De Antonellis, and Michele Melchiori. Semantic collaborative tagging for Web APIs sharing and reuse. Lecture Notes in Computer Science, 7387:76–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31753-8_6/.

Arellano:2012:OPP

[751] Cristóbal Arellano, Oscar Díaz, and Jon Iturrioz. Opening personalization to partners: An architecture of participation for websites. Lecture Notes in Computer Science, 7387:91–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31753-8_7/.

Belchior:2012:RBA

[752] Mairon Belchior, Daniel Schwabe, and Fernando Silva Parreiras. Rolebased access control for model-driven Web applications. Lecture Notes in Computer Science, 7387:106–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31753-8_8/.

Alalfi:2012:RRB

[753] Manar H. Alalfi, James R. Cordy, and Thomas R. Dean. Recovering role-based access control security models from dynamic Web applications. Lecture Notes in Computer Science, 7387:121–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31753-8_9/.

Bozzon:2012:DMD

[754] Alessandro Bozzon, Marco Brambilla, Piero Fraternali, and Marco Tagliasacchi. Diversification for multi-domain result sets. Lecture Notes in Computer Science, 7387:137–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31753-8_10/.

Tao:2012:TSE

[755] Ke Tao, Fabian Abel, Claudia Hauff, and Geert-Jan Houben. Twinder: a search engine for Twitter streams. Lecture Notes in Computer Science, 7387:153–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31753-8_11/.

Ilina:2012:SED

[756] Elena Ilina, Claudia Hauff, Ilknur Celik, Fabian Abel, and Geert-Jan Houben. Social event detection on Twitter. Lecture Notes in Computer Science, 7387:169–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31753-8_12/.

Leprovost:2012:TSC

[757] Damien Leprovost, Lylia Abrouk, Nadine Cullot, and David Gross-Amblard. Temporal semantic centrality for the analysis of communication networks. Lecture Notes in Computer Science, 7387:177–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31753-8_13/.

Wimmer:2012:SEW

[758] Manuel Wimmer, Nathalie Moreno, and Antonio Vallecillo. Systematic evolution of WebML models by coupled transformations. Lecture Notes in Computer Science, 7387:185–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31753-8_14/.

Grigera:2012:RWA

[759] Julián Grigera, José Matías Rivero, Esteban Robles Luna, Franco Giacosa, and Gustavo Rossi. From requirements to Web applications in an agile model-driven approach. Lecture Notes in Computer Science, 7387:200–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31753-8_15/.

Koch:2012:AER

[760] Nora Koch, Alexander Knapp, and Sergej Kozuruba. Assessment of effort reduction due to model-to-model transformations in the Web domain. Lecture Notes in Computer Science, 7387:215–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31753-8_16/.

Martinez:2012:EIM

[761] Yulkeidi Martínez, Cristina Cachero, and Santiago Meliá. Evaluating the impact of a model-driven Web engineering approach on the productivity and the satisfaction of software development teams. Lecture Notes in Computer Science, 7387:223–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31753-8_17/.

Mirshokraie:2012:JJA

[762] Shabnam Mirshokraie and Ali Mesbah. JSART: JavaScript assertionbased regression testing. Lecture Notes in Computer Science, 7387:238– 252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31753-8_18/.

Comai:2012:FDH

[763] Sara Comai, Davide Mazza, and Andrea Guarinoni. A framework for the development of haptic-enhanced Web applications. Lecture Notes in Computer Science, 7387:253–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31753-8_19/.

Firmenich:2012:SUT

[764] Sergio Firmenich, Vincent Gaits, Silvia Gordillo, Gustavo Rossi, and Marco Winckler. Supporting users tasks with personal information management and Web forms augmentation. Lecture Notes in Computer Science, 7387:268–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31753-8_20/.

Anonymous:2012:FMbn

[765] Anonymous. Front matter. Lecture Notes in Computer Science, 7387: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31753-8/1.

Beyersdorff:2012:PCN

[766] Olaf Beyersdorff and Oliver Kutz. Proof complexity of non-classical logics. Lecture Notes in Computer Science, 7388:1–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31485-8_1/.

Broersen:2012:TPD

[767] Jan Broersen and Leendert van der Torre. Ten problems of Deontic logic and normative reasoning in computer science. Lecture Notes in Computer Science, 7388:55–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31485-8_2/.

DalLago:2012:SII

[768] Ugo Dal Lago. A short introduction to implicit computational complexity. Lecture Notes in Computer Science, 7388:89–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31485-8_3/.

Ghosh:2012:SGL

[769] Sujata Ghosh and R. Ramanujam. Strategies in games: a logic-automata study. Lecture Notes in Computer Science, 7388:110–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31485-8_4/.

Grossi:2012:IJA

[770] Davide Grossi and Gabriella Pigozzi. Introduction to judgment aggregation. Lecture Notes in Computer Science, 7388:160–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31485-8_5/.

Jamroga:2012:SVM

[771] Wojciech Jamroga and Wojciech Penczek. Specification and verification of multi-agent systems. Lecture Notes in Computer Science, 7388:210– 263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31485-8_6/.

Anonymous:2012:BMk

[772] Anonymous. Back matter. Lecture Notes in Computer Science, 7388: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31485-8/1.

Anonymous:2012:FMbo

[773] Anonymous. Front matter. Lecture Notes in Computer Science, 7388: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31485-8/1.

Hong:2012:PAN

[774] Xia Hong and Sheng Chen. PSO assisted NURB neural network identification. Lecture Notes in Computer Science, 7389:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31588-6_1/.

Cielecki:2012:DFA

[775] Lukasz Cielecki and Olgierd Unold. 2D discontinuous function approximation with real-valued grammar-based classifier system. Lecture Notes in Computer Science, 7389:10–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31588-6_2/.

Li:2012:SSM

[776] Zhi-Zeng Li, Xue-Yan Song, Ji-Zhou Sun, and Zhao-Tong Huang. A scatter search methodology for the aircraft conflict resolution problem. Lecture Notes in Computer Science, 7389:18–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31588-6_3/.

Xie:2012:SAD

[777] Datong Xie, Lixin Ding, Shenwen Wang, Zhaolu Guo, Yurong Hu, and Chengwang Xie. Self-adaptive differential evolution based multiobjective optimization incorporating local search and indicator-based selection. Lecture Notes in Computer Science, 7389:25–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31588-6_4/.

Zhou:2012:ADM

[778] Tianpei Zhou. Application of data mining in coal mine safety decision system based on rough set. Lecture Notes in Computer Science, 7389:34– 41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31588-6_5/.

Lu:2012:SWT

[779] Xianghua Lu and Weijing Wang. Study on Web text feature selection based on rough set. Lecture Notes in Computer Science, 7389:42–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31588-6_6/.

Bai:2012:GAD

[780] Luyi Bai and Zongmin Ma. A generalized approach for determining fuzzy temporal relations. Lecture Notes in Computer Science, 7389:49– 56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31588-6_7/.

Xing:2012:AIF

[781] Hongyan Xing and Daowen Qiu. Applications on information flow and biomedical treatment of FDES based on fuzzy sequential machines theory. Lecture Notes in Computer Science, 7389:57–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31588-6_8/.

Shao:2012:DFS

[782] Yabin Shao and Zengtai Gong. Discontinuous fuzzy systems and Henstock integrals of fuzzy number valued functions. Lecture Notes in Computer Science, 7389:65–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31588-6_9/.

Yu:2012:PAP

[783] Haiping Yu and Fengying Yang. A phased adaptive PSO algorithm for multimodal function optimization. Lecture Notes in Computer Science, 7389:73–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31588-6_10/.

Huang:2012:CSD

[784] Guoliang Huang, Xinling Shi, Zhenzhou An, and He Sun. The comparative study of different number of particles in clustering based on three-layer particle swarm optimization. Lecture Notes in Computer Science, 7389:80–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31588-6_11/.

Lee:2012:IML

[785] Yang Weon Lee. Implementation of mutual localization of multi-robot using particle filter. Lecture Notes in Computer Science, 7389:87–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31588-6_12/.

Malekzadeh:2012:OOP

[786] Milad Malekzadeh, Alireza Khosravi, Saeed Alighale, and Hamed Azami. Optimization of orthogonal poly phase coding waveform based on bees algorithm and artificial bee colony for MIMO radar. Lecture Notes in Computer Science, 7389:95–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31588-6_13/.

Dong:2012:SRM

[787] Yinli Dong and Shuisheng Zhou. SVM regularizer models on RKHS vs. on R^{m}. Lecture Notes in Computer Science, 7389:103–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31588-6_14/.

Zhang:2012:RPC

[788] Xing Zhang. Research on performance comprehensive evaluation of thermal power plant under low-carbon economy. Lecture Notes in Computer Science, 7389:112–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31588-6_15/.

Zhang:2012:CMC

[789] Hui Zhang, Qing Guo, and Costas S. Iliopoulos. Computing the minimum λ-cover in weighted sequences. Lecture Notes in Computer Science, 7389:120–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31588-6_16/.

Geng:2012:NHE

[790] Huantong Geng, Haifeng Zhu, Rui Xing, and Tingting Wu. A novel hybrid evolutionary algorithm for solving multi-objective optimization problems. Lecture Notes in Computer Science, 7389:128–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31588-6_17/.

Binh:2012:HAS

[791] Huynh Thi Thanh Binh, Pham Vu Long, Nguyen Ngoc Dat, and Nguyen Sy Thai Ha. Heuristic algorithms for solving survivable network design problem with simultaneous unicast and anycast flows. Lecture Notes in Computer Science, 7389:137–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31588-6_18/.

Li:2012:PPB

[792] Xueling Li, Min Zhu, Xiaolai Li, Hong-Qiang Wang, and Shulin Wang. Protein-protein binding affinity prediction based on an SVR ensemble. Lecture Notes in Computer Science, 7389:145–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31588-6_19/.

Wei:2012:NTS

[793] Xiaoli Wei, Xue Shi, Seongho Kim, Craig McClain, and Xiang Zhang. A novel two-stage alignment method for liquid chromatography mass spectrometry-based metabolomics. Lecture Notes in Computer Science, 7389:152–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31588-6_20/.

Anonymous:2012:FMbp

[794] Anonymous. Front matter. Lecture Notes in Computer Science, 7389: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31588-6/1.

Liu:2012:MDM

[795] Xiaoming Liu, Bo Li, Jun Liu, Xin Xu, and Zhilin Feng. Mass diagnosis in mammography with mutual information based feature selection and support vector machine. Lecture Notes in Computer Science, 7390: 1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31576-3_1/.

Lou:2012:MSE

[796] Xuyang Lou, Qian Ye, Ke Lou, and Baotong Cui. Mean square exponential stability of hybrid neural networks with uncertain switching probabilities. Lecture Notes in Computer Science, 7390:9–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31576-3_2/.

Sun:2012:TVN

[797] Ming xuan Sun. Time-varying neurocomputing: An iterative learning perspective. Lecture Notes in Computer Science, 7390:18–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31576-3_3/.

Wang:2012:KRB

[798] Hong Wang and Huijuan Shi. Knowledge reduction based on evidence reasoning theory in interval ordered information systems. Lecture Notes in Computer Science, 7390:27–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31576-3_4/.

Zhang:2012:RNA

[799] Yong-Heng Zhang and Feng Zhang. Research on new algorithm of topicoriented crawler and duplicated Web pages detection. Lecture Notes in Computer Science, 7390:35–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31576-3_5/.

Lv:2012:DSS

[800] Xiaochun Lv and Hanming Gu. Direct search simulated annealing for nonlinear global optimization of Rayleigh waves. Lecture Notes in Computer Science, 7390:43–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31576-3_6/.

Yu:2012:FQD

[801] Ke-Fan Yu and Ke-Han Zeng. Four-quadrant division with HNN for Euclidean TSP. Lecture Notes in Computer Science, 7390:51–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31576-3_7/.

Xu:2012:IGA

[802] Hai-Qin Xu, Yong-Sheng Ding, Hao Liu, and Xiao-Li Li. An immune genetic algorithm with orthogonal initialization for analog circuit design. Lecture Notes in Computer Science, 7390:59–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31576-3_8/.

Zhai:2012:NAF

[803] Yi-Kui Zhai, Ying Xu, and Jun-Ying Gan. A novel artificial fish swarm algorithm based on multi-objective optimization. Lecture Notes in Computer Science, 7390:67–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31576-3_9/.

Skrobanek:2012:CDE

[804] Paweł Skrobanek, Henryk Maciejewski, Maciej Dobrowolski, Olgierd Unold, and Ewa Walkowicz. Comparison of data exploration methods and the BLUP method in application to the animals breeding. Lecture Notes in Computer Science, 7390:74–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31576-3_10/.

Zhang:2012:TDL

[805] Shan-Wen Zhang and Chuan-Lei Zhang. Two-dimensional locality discriminant projection for plant leaf classification. Lecture Notes in Computer Science, 7390:82–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31576-3_11/.

Sun:2012:EBC

[806] Zhan-Li Sun, Ting-Ting Sun, and Yang Liu. An eigengene-based classifier committee learning algorithm for tumor classification. Lecture Notes in Computer Science, 7390:89–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31576-3_12/.

Bevilacqua:2012:ESI

[807] Vitoantonio Bevilacqua, Maurizio Triggiani, Vito Gallo, and Isabella Cafagna. An expert system for an innovative discrimination tool of commercial table grapes. Lecture Notes in Computer Science, 7390:95– 102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31576-3_13/.

Li:2012:SRS

[808] Zhaowen Li and Ninghua Gao. Soft rough sets and topologies. Lecture Notes in Computer Science, 7390:103–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31576-3_14/.

Lu:2012:STS

[809] Peng Lu, Chi Liu, Daoren Yuan, and Yusong Li. Sudden-target search algorithm of monitor area based on visual information processing model. Lecture Notes in Computer Science, 7390:111–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31576-3_15/.

Yu:2012:URT

[810] Jie Yu, Haihong Zhao, and Fangfang Liu. User real-time interest prediction based on collaborative filtering and interactive computing for academic recommendation. Lecture Notes in Computer Science, 7390:119– 127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31576-3_16/.

Cao:2012:SBF

[811] Lu Cao, Yoshinori Kobayashi, and Yoshinori Kuno. Spatial-based feature for locating objects. Lecture Notes in Computer Science, 7390:128– 137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31576-3_17/.

Luo:2012:NPH

[812] Jian-Fei Luo, Bao-Yuan Wu, Qiu-Shi Lin, Fei Shen, and Zhong-Cheng Wu. A novel F -pad for handwriting force information acquisition. Lecture Notes in Computer Science, 7390:138–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31576-3_18/.

Liu:2012:SPP

[813] Jia Liu, Hua Zhou, and Kai yu Tong. Stability of a predator-prey model with modified Holling-Type II functional response. Lecture Notes in Computer Science, 7390:145–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31576-3_19/.

Zhou:2012:UMO

[814] Qing-Hua Zhou, Jing Cui, and Juan Xie. Using multiple objective functions in the dynamic model of metabolic networks of Escherichia coli. Lecture Notes in Computer Science, 7390:151–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31576-3_20/.

Anonymous:2012:FMbq

[815] Anonymous. Front matter. Lecture Notes in Computer Science, 7390: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31576-3/1.

Achlioptas:2012:UBR

[816] Dimitris Achlioptas and Ricardo Menchaca-Mendez. Unsatisfiability bounds for random CSPs from an energetic interpolation method. Lecture Notes in Computer Science, 7391:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31594-7_1/.

Ada:2012:NMC

[817] Anil Ada, Arkadev Chattopadhyay, Omar Fawzi, and Phuong Nguyen. The NOF multiparty communication complexity of composed functions. Lecture Notes in Computer Science, 7391:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31594-7_2/.

Ambainis:2012:QSB

[818] Andris Ambainis, Artūrs Bačkurs, Kaspars Balodis, Dmitrijs Kravčenko, and Raitis Ozols. Quantum strategies are better than classical in almost any XOR game. Lecture Notes in Computer Science, 7391:25– 37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31594-7_3/.

Azar:2012:ESF

[819] Yossi Azar and Iftah Gamzu. Efficient submodular function maximization under linear packing constraints. Lecture Notes in Computer Science, 7391:38–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31594-7_4/.

Babai:2012:PTI

[820] László Babai, Paolo Codenotti, and Youming Qiao. Polynomial-time isomorphism test for groups with no Abelian normal subgroups. Lecture Notes in Computer Science, 7391:51–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31594-7_5/.

Balcan:2012:CUP

[821] Maria Florina Balcan and Yingyu Liang. Clustering under perturbation resilience. Lecture Notes in Computer Science, 7391:63–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31594-7_6/.

Barman:2012:SPC

[822] Siddharth Barman, Seeun Umboh, Shuchi Chawla, and David Malec. Secretary problems with convex costs. Lecture Notes in Computer Science, 7391:75–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31594-7_7/.

Baron:2012:NSR

[823] Joshua Baron, Rafail Ostrovsky, and Ivan Visconti. Nearly simultaneously resettable black-box zero knowledge. Lecture Notes in Computer Science, 7391:88–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31594-7_8/.

Bauwens:2012:CCM

[824] Bruno Bauwens. Complexity of complexity and maximal plain versus prefix-free Kolmogorov complexity. Lecture Notes in Computer Science, 7391:100–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31594-7_9/.

Bhaskara:2012:QPR

[825] Aditya Bhaskara, Moses Charikar, Rajsekar Manokaran, and Aravindan Vijayaraghavan. On quadratic programming with a ratio objective. Lecture Notes in Computer Science, 7391:109–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31594-7_10/.

Bose:2012:ABS

[826] Prosenjit Bose, Sébastien Collette, Rolf Fagerberg, and Stefan Langerman. De-amortizing binary search trees. Lecture Notes in Computer Science, 7391:121–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31594-7_11/.

Bringmann:2012:ESM

[827] Karl Bringmann and Konstantinos Panagiotou. Efficient sampling methods for discrete distributions. Lecture Notes in Computer Science, 7391:133–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31594-7_12/.

Buchbinder:2012:AAO

[828] Niv Buchbinder, Joseph (Seffi) Naor, R. Ravi, and Mohit Singh. Approximation algorithms for online weighted rank function maximization under matroid constraints. Lecture Notes in Computer Science, 7391:145– 156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31594-7_13/.

Byrka:2012:ILR

[829] Jaroslaw Byrka and Bartosz Rybicki. Improved LP-rounding approximation algorithm for k-level uncapacitated facility location. Lecture Notes in Computer Science, 7391:157–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31594-7_14/.

Chakrabarty:2012:TCF

[830] Deeparnab Chakrabarty and Zhiyi Huang. Testing coverage functions. Lecture Notes in Computer Science, 7391:170–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31594-7_15/.

Chan:2012:SFT

[831] T.-H. Hubert Chan, Mingfei Li, and Li Ning. Sparse fault-tolerant spanners for doubling metrics with bounded hop-diameter or degree. Lecture Notes in Computer Science, 7391:182–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31594-7_16/.

Charikar:2012:DLR

[832] Moses Charikar and Shi Li. A dependent LP-rounding approach for the k-median problem. Lecture Notes in Computer Science, 7391:194– 205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31594-7_17/.

Chekuri:2012:NWN

[833] Chandra Chekuri, Alina Ene, and Ali Vakilian. Node-weighted network design in planar and minor-closed families of graphs. Lecture Notes in Computer Science, 7391:206–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31594-7_18/.

Chen:2012:CVP

[834] Danny Z. Chen and Haitao Wang. Computing the visibility polygon of an Island in a polygonal domain. Lecture Notes in Computer Science, 7391:218–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31594-7_19/.

Chitnis:2012:DSF

[835] Rajesh Chitnis, Marek Cygan, Mohammadtaghi Hajiaghayi, and Dániel Marx. Directed subset feedback vertex set is fixed-parameter tractable. Lecture Notes in Computer Science, 7391:230–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31594-7_20/.

Anonymous:2012:FMbr

[836] Anonymous. Front matter. Lecture Notes in Computer Science, 7391: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31594-7/1.

Colini-Baldeschi:2012:MKS

[837] Riccardo Colini-Baldeschi, Monika Henzinger, Stefano Leonardi, and Martin Starnberger. On multiple keyword sponsored search auctions with budgets. Lecture Notes in Computer Science, 7392:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31585-5_1/.

Dowek:2012:TIC

[838] Gilles Dowek. A theory independent Curry–De Bruijn–Howard correspondence. Lecture Notes in Computer Science, 7392:13–15, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-31585-5_
2/.

Harel:2012:SSG

[839] David Harel. Standing on the shoulders of a giant. Lecture Notes in Computer Science, 7392:16–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31585-5_3/.

Honda:2012:STD

[840] Kohei Honda. Session types and distributed computing. Lecture Notes in Computer Science, 7392:23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-31585-5_4.

Spielman:2012:AGT

[841] Daniel A. Spielman. Algorithms, graph theory, and the solution of Laplacian linear equations. Lecture Notes in Computer Science, 7392:24–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31585-5_5/.

Vocking:2012:RMM

[842] Berthold Vöcking. Randomized mechanisms for multi-unit auctions. Lecture Notes in Computer Science, 7392:27–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31585-5_6/.

Aceto:2012:AST

[843] Luca Aceto, Arnaud Carayol, Zoltán Ésik, and Anna Ingólfsdóttir. Algebraic synchronization trees and processes. Lecture Notes in Computer Science, 7392:30–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31585-5_7/.

Alur:2012:STT

[844] Rajeev Alur and Loris D’Antoni. Streaming tree transducers. Lecture Notes in Computer Science, 7392:42–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31585-5_8/.

Arrighi:2012:CGD

[845] Pablo Arrighi and Gilles Dowek. Causal graph dynamics. Lecture Notes in Computer Science, 7392:54–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31585-5_9/.

Atserias:2012:DLB

[846] Albert Atserias and Anuj Dawar. Degree lower bounds of tower-type for approximating formulas with parity quantifiers. Lecture Notes in Computer Science, 7392:67–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31585-5_10/.

Benedikt:2012:MDC

[847] Michael Benedikt, Pierre Bourhis, and Pierre Senellart. Monadic datalog containment. Lecture Notes in Computer Science, 7392:79–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31585-5_11/.

Bojanczyk:2012:MIC

[848] Mikołaj Bojańczyk and Sławomir Lasota. A machine-independent characterization of timed languages. Lecture Notes in Computer Science, 7392:92–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31585-5_12/.

Bojanczyk:2012:RLI

[849] Mikołaj Bojańczyk and Thomas Place. Regular languages of infinite trees that are Boolean combinations of open sets. Lecture Notes in Computer Science, 7392:104–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31585-5_13/.

Bojanczyk:2012:TMT

[850] Mikołaj Bojańczyk and Thomas Place. Toward model theory with data values. Lecture Notes in Computer Science, 7392:116–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31585-5_14/.

Bouyer:2012:RRT

[851] Patricia Bouyer, Nicolas Markey, and Ocan Sankur. Robust reachability in timed automata: A Game-Based Approach. Lecture Notes in Computer Science, 7392:128–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31585-5_15/.

Brazdil:2012:MET

[852] Tomáš Brázdil, Antonín Kučera, Petr Novotný, and Dominik Wojtczak. Minimizing expected termination time in one-counter Markov decision processes. Lecture Notes in Computer Science, 7392:141–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31585-5_16/.

Broadbent:2012:PRN

[853] Christopher Broadbent. Prefix rewriting for nested-words and collapsible pushdown automata. Lecture Notes in Computer Science, 7392:153– 164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31585-5_17/.

Broadbent:2012:SMC

[854] Chris Broadbent, Arnaud Carayol, Matthew Hague, and Olivier Serre. A saturation method for collapsible pushdown systems. Lecture Notes in Computer Science, 7392:165–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31585-5_18/.

Diekert:2012:RLC

[855] Volker Diekert, Manfred Kufleitner, Klaus Reinhardt, and Tobias Walter. Regular languages are Church–Rosser congruential. Lecture Notes in Computer Science, 7392:177–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31585-5_19/.

Fearnley:2012:TPR

[856] John Fearnley and Sven Schewe. Time and parallelizability results for parity games with bounded treewidth. Lecture Notes in Computer Science, 7392:189–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31585-5_20/.

Anonymous:2012:FMbs

[857] Anonymous. Front matter. Lecture Notes in Computer Science, 7392: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31585-5/1.

Demolombe:2012:CCM

[858] Robert Demolombe. Causality in the context of multiple agents. Lecture Notes in Computer Science, 7393:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31570-1_1/.

Tosatto:2012:VNS

[859] Silvano Colombo Tosatto and Guido Boella. Visualizing normative systems: An abstract approach. Lecture Notes in Computer Science, 7393: 16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31570-1_2/.

Turrini:2012:AN

[860] Paolo Turrini. Agreements as norms. Lecture Notes in Computer Science, 7393:31–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31570-1_3/.

Governatori:2012:PWS

[861] Guido Governatori, Antonino Rotolo, and Erica Calardo. Possible world semantics for defeasible Deontic logic. Lecture Notes in Computer Science, 7393:46–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31570-1_4/.

Goncalves:2012:EIO

[862] Ricardo Gonçalves and José Júlio Alferes. An embedding of inputoutput logic in Deontic logic programs. Lecture Notes in Computer Science, 7393:61–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31570-1_5/.

Kuijer:2012:SSC

[863] Louwe B. Kuijer. Sanction semantics and contrary-to-duty obligations. Lecture Notes in Computer Science, 7393:76–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31570-1_6/.

Gabbay:2012:TDL

[864] Dov Gabbay. Temporal Deontic logic for the generalised Chisholm set of contrary to duty obligations. Lecture Notes in Computer Science, 7393:91–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31570-1_7/.

Lorini:2012:LFM

[865] Emiliano Lorini. On the logical foundations of moral agency. Lecture Notes in Computer Science, 7393:108–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31570-1_8/.

Riveret:2012:NLP

[866] Régis Riveret, Antonino Rotolo, and Giovanni Sartor. Norms and learning in probabilistic logic-based agents. Lecture Notes in Computer Science, 7393:123–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31570-1_9/.

Roy:2012:LOW

[867] Olivier Roy, Albert J. J. Anglberger, and Norbert Gratzl. The logic of obligation as weakest permission. Lecture Notes in Computer Science, 7393:139–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31570-1_10/.

Strasser:2012:ADL

[868] Christian Straßer and Mathieu Beirlaen. An Andersonian Deontic logic with contextualized sanctions. Lecture Notes in Computer Science, 7393:151–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31570-1_11/.

Herzig:2012:CGD

[869] Andreas Herzig, Tiago de Lima, and Emiliano Lorini. A computationally grounded dynamic logic of agency, with an application to legal actions. Lecture Notes in Computer Science, 7393:170–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31570-1_12/.

Kulicki:2012:DAL

[870] Piotr Kulicki and Robert Trypuz. A Deontic action logic with sequential composition of actions. Lecture Notes in Computer Science, 7393:184– 198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31570-1_13/.

Broersen:2012:DPC

[871] Jan Broersen, Dov Gabbay, and Leendert van der Torre. Discussion paper: Changing norms is changing obligation change. Lecture Notes in Computer Science, 7393:199–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31570-1_14/.

Anonymous:2012:BMl

[872] Anonymous. Back matter. Lecture Notes in Computer Science, 7393: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31570-1/1.

Anonymous:2012:FMbt

[873] Anonymous. Front matter. Lecture Notes in Computer Science, 7393: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31570-1/1.

Anonymous:2012:FMbu

[874] Anonymous. Front matter. Lecture Notes in Computer Science, 7394: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32805-3/1/1.

Nguyen:2012:RRD

[875] Hieu Cuong Nguyen and Stefan Katzenbeisser. Robust resampling detection in digital images. Lecture Notes in Computer Science, 7394:3– 15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32805-3_1/.

Kummel:2012:FSH

[876] Karl Kümmel, Tobias Scheidat, Claus Vielhauer, and Jana Dittmann. Feature selection on handwriting biometrics: Security aspects of artificial forgeries. Lecture Notes in Computer Science, 7394:16–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32805-3_2/.

Shariati:2012:SAI

[877] Saloomeh Shariati, François Koeune, and François-Xavier Standaert. Security analysis of image-based PUFs for anti-counterfeiting. Lecture Notes in Computer Science, 7394:26–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32805-3_3/.

Diong:2012:DAU

[878] Mouhamadou L. Diong, Patrick Bas, Chloé Pelle, and Wadih Sawaya. Document authentication using 2D codes: Maximizing the decoding performance using statistical inference. Lecture Notes in Computer Science, 7394:39–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32805-3_4/.

Bichsel:2012:DMA

[879] Patrik Bichsel, Jan Camenisch, Bart De Decker, Jorn Lapon, and Vincent Naessens. Data-minimizing authentication goes mobile. Lecture Notes in Computer Science, 7394:55–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32805-3_5/.

Wei:2012:NTB

[880] Zhuo Wei, Xuhua Ding, Robert Huijie Deng, and Yongdong Wu. No tradeoff between confidentiality and performance: An analysis on H.264/ SVC partial encryption. Lecture Notes in Computer Science, 7394:72– 86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32805-3_6/.

Anonymous:2012:FMbv

[881] Anonymous. Front matter. Lecture Notes in Computer Science, 7394: 87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32805-3/2/1.

Makrushin:2012:CAC

[882] Andrey Makrushin, Tobias Kiertscher, Robert Fischer, Stefan Gruhn, and Claus Vielhauer. Computer-aided contact-less localization of latent fingerprints in low-resolution CWL scans. Lecture Notes in Computer Science, 7394:89–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32805-3_7/.

Horsman:2012:MRR

[883] Graeme Horsman, Christopher Laing, and Paul Vickers. A method for reducing the risk of errors in digital forensic investigations. Lecture Notes in Computer Science, 7394:99–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32805-3_8/.

Scheidat:2012:STT

[884] Tobias Scheidat, Karl Kümmel, and Claus Vielhauer. Short term template aging effects on biometric dynamic handwriting authentication performance. Lecture Notes in Computer Science, 7394:107–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32805-3_9/.

Schmitz:2012:NAC

[885] Roland Schmitz, Shujun Li, Christos Grecos, and Xinpeng Zhang. A new approach to commutative watermarking-encryption. Lecture Notes in Computer Science, 7394:117–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32805-3_10/.

Wendzel:2012:SEC

[886] Steffen Wendzel and Jörg Keller. Systematic engineering of control protocols for covert channels. Lecture Notes in Computer Science, 7394:131– 144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32805-3_11/.

Franz:2012:ESN

[887] Elke Franz, Stefan Pfennig, and André Fischer. Efficiency of secure network coding schemes. Lecture Notes in Computer Science, 7394:145– 159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32805-3_12/.

Siad:2012:NAP

[888] Amar Siad. A new approach for private searches on public-key encrypted data. Lecture Notes in Computer Science, 7394:160–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32805-3_13/.

Maachaoui:2012:MLA

[889] M. Maachaoui, A. Abou El Kalam, C. Fraboul, and A. Ait Ouahman. Multi-level authentication based single sign-on for IMS services. Lecture Notes in Computer Science, 7394:174–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32805-3_14/.

Anonymous:2012:FMbw

[890] Anonymous. Front matter. Lecture Notes in Computer Science, 7394: 189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32805-3/3/1.

Bas:2012:BLK

[891] Patrick Bas and Teddy Furon. Are 128 bits Long keys possible in watermarking? Lecture Notes in Computer Science, 7394:191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-32805-3_15.

Perrizo:2012:PTB

[892] William Perrizo and Arjun G. Roy. Predicate-tree based pretty good privacy of data. Lecture Notes in Computer Science, 7394:192–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32805-3_16/.

Langweg:2012:OFA

[893] Hanno Langweg. OOXML file analysis of the July 22nd Terrorist Manual. Lecture Notes in Computer Science, 7394:195–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32805-3_17/.

Milutinovic:2012:PPS

[894] Milica Milutinovic, Vincent Naessens, and Bart De Decker. Privacypreserving scheduling mechanism for eHealth systems. Lecture Notes in Computer Science, 7394:198–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32805-3_18/.

Fuss:2012:CAI

[895] Jürgen Fuß, Wolfgang Kastl, Robert Kolmhofer, Georg Schönberger, and Florian Wex. Cuteforce analyzer: Implementing a heterogeneous bruteforce cluster with specialized coprocessors. Lecture Notes in Computer Science, 7394:201–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32805-3_19/.

Zuquete:2012:FEU

[896] André Zúquete, Pedro Correia, and Miguel Rocha. A framework for enforcing user-based authorization policies on packet filter firewalls. Lecture Notes in Computer Science, 7394:204–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32805-3_20/.

Anonymous:2012:FMbx

[897] Anonymous. Front matter. Lecture Notes in Computer Science, 7394: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32805-3/1.

Bastenhof:2012:PMS

[898] Arno Bastenhof. Polarized Montagovian semantics for the Lambek– Grishin calculus. Lecture Notes in Computer Science, 7395:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32024-8_1/.

Bechet:2012:TML

[899] Denis Béchet, Alexander Dikovsky, and Annie Foret. Two models of learning iterated dependencies. Lecture Notes in Computer Science, 7395: 17–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32024-8_2/.

Bransen:2012:LGC

[900] Jeroen Bransen. The Lambek–Grishin calculus is NP-Complete. Lecture Notes in Computer Science, 7395:33–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32024-8_3/.

Crysmann:2012:RIH

[901] Berthold Crysmann. Resumption and island-hood in Hausa. Lecture Notes in Computer Science, 7395:50–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32024-8_4/.

Dekhtyar:2012:IDK

[902] Michael Dekhtyar, Alexander Dikovsky, and Boris Karlov. Iterated dependencies and Kleene iteration. Lecture Notes in Computer Science, 7395:66–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32024-8_5/.

Duchier:2012:PGP

[903] Denys Duchier, Thi-Bich-Hanh Dao, Yannick Parmentier, and Willy Lesaint. Property grammar parsing seen as a constraint optimization problem. Lecture Notes in Computer Science, 7395:82–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32024-8_6/.

Graf:2012:RSC

[904] Thomas Graf. Reference-set constraints as linear tree transductions via controlled optimality systems. Lecture Notes in Computer Science, 7395:97–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32024-8_7/.

Martin:2012:HDS

[905] Scott Martin and Carl Pollard. Hyperintensional dynamic semantics. Lecture Notes in Computer Science, 7395:114–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32024-8_8/.

Mihalicek:2012:DPT

[906] Vedrana Mihaliček and Carl Pollard. Distinguishing phenogrammar from tectogrammar simplifies the analysis of interrogatives. Lecture Notes in Computer Science, 7395:130–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32024-8_9/.

Morrill:2012:GD

[907] Glyn Morrill and Oriol Valentín. Generalized discontinuity. Lecture Notes in Computer Science, 7395:146–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32024-8_10/.

Pogodalla:2012:CEA

[908] Sylvain Pogodalla and Florent Pompigne. Controlling extraction in abstract categorial grammars. Lecture Notes in Computer Science, 7395:162–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32024-8_11/.

Yi:2012:PQG

[909] Byeong-Uk Yi. Plural quantifications and generalized quantifiers. Lecture Notes in Computer Science, 7395:178–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32024-8_12/.

Yoshinaka:2012:PTL

[910] Ryo Yoshinaka and Alexander Clark. Polynomial time learning of some multiple context-free languages with a minimally adequate teacher. Lecture Notes in Computer Science, 7395:192–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32024-8_13/.

Graf:2012:LCM

[911] Thomas Graf. Locality and the complexity of minimalist derivation tree languages. Lecture Notes in Computer Science, 7395:208–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32024-8_14/.

Homola:2012:BFG

[912] Petr Homola. Building a formal grammar for a polysynthetic language. Lecture Notes in Computer Science, 7395:228–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32024-8_15/.

Kornai:2012:ED

[913] András Kornai. Eliminating ditransitives. Lecture Notes in Computer Science, 7395:243–261, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32024-8_16/.

Kuznetsov:2012:LGU

[914] Stepan Kuznetsov. Lambek grammars with the unit. Lecture Notes in Computer Science, 7395:262–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32024-8_17/.

Leiss:2012:RPA

[915] Hans Leiß. Resolving plural ambiguities by type reconstruction. Lecture Notes in Computer Science, 7395:267–286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32024-8_18/.

Martin:2012:WFA

[916] Scott Martin. Weak familiarity and anaphoric accessibility in dynamic semantics. Lecture Notes in Computer Science, 7395:287–306, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32024-8_19/.

Anonymous:2012:BMm

[917] Anonymous. Back matter. Lecture Notes in Computer Science, 7395: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32024-8/1.

Anonymous:2012:FMby

[918] Anonymous. Front matter. Lecture Notes in Computer Science, 7395: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32024-8/1.

Pernici:2012:SEE

[919] Barbara Pernici, Cinzia Cappiello, Maria Grazia Fugini, and Pierluigi Plebani. Setting energy efficiency goals in data centers: The GAMES approach. Lecture Notes in Computer Science, 7396:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33645-4_1/.

Klingert:2012:SEM

[920] Sonja Klingert, Andreas Berl, Michael Beck, Radu Serban, and Marco di Girolamo. Sustainable energy management in data centers through collaboration. Lecture Notes in Computer Science, 7396:13– 24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33645-4_2/.

vordemBerge:2012:MSD

[921] Micha vor dem Berge, Georges Da Costa, Andreas Kopecki, and Ariel Oleksiak. Modeling and simulation of data center energyefficiency in CoolEmAll. Lecture Notes in Computer Science, 7396:25– 36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33645-4_3/.

Ziegler:2012:SEE

[922] Wolfgang Ziegler. SLAs for energy-efficient data centres: The standardsbased approach of the OPTIMIS project. Lecture Notes in Computer Science, 7396:37–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33645-4_4/.

Israel:2012:GDB

[923] Ywes Israel and Thomas Leitert. The GreenIT DC-benchmarking tool: From scientific theory to real life. Lecture Notes in Computer Science, 7396:47–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33645-4_5/.

Bunse:2012:GSE

[924] Christian Bunse, Sonja Klingert, and Thomas Schulze. GreenSLAs: Supporting energy-efficiency through contracts. Lecture Notes in Computer Science, 7396:54–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33645-4_6/.

Anagnostopoulou:2012:PAR

[925] Vlasia Anagnostopoulou, Susmit Biswas, Heba Saadeldeen, Ricardo Bianchini, and Tao Yang. Power-aware resource allocation for CPUand memory-intense Internet services. Lecture Notes in Computer Science, 7396:69–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33645-4_7/.

Guazzone:2012:EVM

[926] Marco Guazzone, Cosimo Anglano, and Massimo Canonico. Exploiting VM migration for the automated power and performance management of green cloud computing systems. Lecture Notes in Computer Science, 7396:81–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33645-4_8/.

Lucanin:2012:EES

[927] Drazen Lucanin, Michael Maurer, Toni Mastelic, and Ivona Brandic. Energy efficient service delivery in clouds in compliance with the Kyoto protocol. Lecture Notes in Computer Science, 7396:93–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33645-4_9/.

Pervila:2012:IEW

[928] Mikko Pervilä, Mikko Rantanen, and Jussi Kangasharju. Implementation and evaluation of a wired data center sensor network. Lecture Notes in Computer Science, 7396:105–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33645-4_10/.

Curry:2012:ECD

[929] Edward Curry, Souleiman Hasan, Mark White, and Hugh Melvin. An environmental chargeback for data center and cloud computing consumers. Lecture Notes in Computer Science, 7396:117–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33645-4_11/.

Quan:2012:EUC

[930] Dang Minh Quan, Andrey Somov, and Corentin Dupont. Energy usage and carbon emission optimization mechanism for federated data centers. Lecture Notes in Computer Science, 7396:129–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-33645-4_12/.

Chetsa:2012:DIS

[931] Ghislain Landry Tsafack Chetsa, Laurent Lefevre, Jean-Marc Pierson, and Patricia Stolf. DNA-inspired scheme for building the energy profile of HPC systems. Lecture Notes in Computer Science, 7396:141– 152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33645-4_13/.

Anonymous:2012:BMn

[932] Anonymous. Back matter. Lecture Notes in Computer Science, 7396: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33645-4/1.

Anonymous:2012:FMbz

[933] Anonymous. Front matter. Lecture Notes in Computer Science, 7396: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33645-4/1.

Kim:2012:SAH

[934] Hyoungshick Kim, John Tang, and Ross Anderson. Social authentication: Harder than it looks. Lecture Notes in Computer Science, 7397: 1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32946-3_1/.

Chiasson:2012:MWB

[935] Sonia Chiasson and Chris Deschamps. The MVP Web-based authentication framework. Lecture Notes in Computer Science, 7397:16–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32946-3_2/.

Bonneau:2012:BPE

[936] Joseph Bonneau, Sören Preibusch, and Ross Anderson. A birthday present every eleven wallets? The security of customer-chosen banking PINs. Lecture Notes in Computer Science, 7397:25–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32946-3_3/.

Moore:2012:PPS

[937] Tyler Moore, Jie Han, and Richard Clayton. The postmodern Ponzi scheme: Empirical analysis of high-yield investment programs. Lecture Notes in Computer Science, 7397:41–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32946-3_4/.

Bogdanov:2012:DSM

[938] Dan Bogdanov, Riivo Talviste, and Jan Willemson. Deploying secure multi-party computation for financial data analysis. Lecture Notes in Computer Science, 7397:57–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32946-3_5/.

Thorpe:2012:CRB

[939] Christopher Thorpe and Steven R. Willis. Cryptographic rule-based trading. Lecture Notes in Computer Science, 7397:65–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32946-3_6/.

Lin:2012:EPP

[940] Zi Lin, Denis Foo Kune, and Nicholas Hopper. Efficient private proximity testing with GSM location sketches. Lecture Notes in Computer Science, 7397:73–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32946-3_7/.

Johnson:2012:MMI

[941] Benjamin Johnson, John Chuang, and Jens Grossklags. Metrics for measuring ISP badness: The case of spam. Lecture Notes in Computer Science, 7397:89–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32946-3_8/.

Wang:2012:CAP

[942] Tao Wang, Kevin Bauer, and Clara Forero. Congestion-aware path selection for Tor. Lecture Notes in Computer Science, 7397:98–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32946-3_9/.

Wolchok:2012:AWD

[943] Scott Wolchok, Eric Wustrow, and Dawn Isabel. Attacking the Washington, D.C. Internet Voting System. Lecture Notes in Computer Science, 7397:114–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32946-3_10/.

Bohme:2012:SAR

[944] Rainer Böhme. Security audits revisited. Lecture Notes in Computer Science, 7397:129–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32946-3_11/.

Yavuz:2012:ECR

[945] Attila A. Yavuz, Peng Ning, and Michael K. Reiter. Efficient, compromise resilient and append-only cryptographic schemes for secure audit logging. Lecture Notes in Computer Science, 7397:148–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32946-3_12/.

Dahl:2012:STP

[946] Morten Dahl, Chao Ning, and Tomas Toft. On secure two-party integer division. Lecture Notes in Computer Science, 7397:164–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32946-3_13/.

Chaabouni:2012:NIR

[947] Rafik Chaabouni, Helger Lipmaa, and Bingsheng Zhang. A noninteractive range proof with constant communication. Lecture Notes in Computer Science, 7397:179–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32946-3_14/.

Chan:2012:PPS

[948] T.-H. Hubert Chan, Elaine Shi, and Dawn Song. Privacy-preserving stream aggregation with fault tolerance. Lecture Notes in Computer Science, 7397:200–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32946-3_15/.

Slamanig:2012:DAB

[949] Daniel Slamanig. Dynamic accumulator based discretionary access control for outsourced storage with unlinkable access. Lecture Notes in Computer Science, 7397:215–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32946-3_16/.

Raykova:2012:PEA

[950] Mariana Raykova, Hang Zhao, and Steven M. Bellovin. Privacy enhanced access control for outsourced data sharing. Lecture Notes in Computer Science, 7397:223–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32946-3_17/.

Molina-Markham:2012:DPP

[951] Andres Molina-Markham, George Danezis, and Kevin Fu. Designing privacy-preserving smart meters with low-cost microcontrollers. Lecture Notes in Computer Science, 7397:239–253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32946-3_18/.

Mood:2012:MEG

[952] Benjamin Mood, Lara Letaw, and Kevin Butler. Memory-efficient garbled circuit generation for mobile devices. Lecture Notes in Computer Science, 7397:254–268, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32946-3_19/.

Mohassel:2012:ODP

[953] Payman Mohassel and Salman Niksefat. Oblivious decision programs from oblivious transfer: Efficient reductions. Lecture Notes in Computer Science, 7397:269–284, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32946-3_20/.

Anonymous:2012:FMca

[954] Anonymous. Front matter. Lecture Notes in Computer Science, 7397: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32946-3/1.

Bonneau:2012:LPM

[955] Joseph Bonneau and Ekaterina Shutova. Linguistic properties of multiword passphrases. Lecture Notes in Computer Science, 7398:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34638-5_1/.

Carlos:2012:UWH

[956] Marcelo Carlos and Geraint Price. Understanding the weaknesses of human-protocol interaction. Lecture Notes in Computer Science, 7398: 13–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34638-5_2/.

Czeskis:2012:HSD

[957] Alexei Czeskis and Jacob Appelbaum. High stakes: Designing a privacy preserving registry. Lecture Notes in Computer Science, 7398:27– 43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34638-5_3/.

Czeskis:2012:PL

[958] Alexei Czeskis and Dirk Balfanz. Protected login. Lecture Notes in Computer Science, 7398:44–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34638-5_4/.

Dong:2012:EUS

[959] Zheng Dong. Enabling users to self-manage networks: Collaborative anomaly detection in wireless personal area networks. Lecture Notes in Computer Science, 7398:53–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34638-5_5/.

Kelley:2012:CPI

[960] Patrick Gage Kelley, Sunny Consolvo, Lorrie Faith Cranor, Jaeyeon Jung, and Norman Sadeh. A conundrum of permissions: Installing applications on an Android Smartphone. Lecture Notes in Computer Science, 7398:68–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34638-5_6/.

Levesque:2012:MFS

[961] Fanny Lalonde Lévesque, Carlton R. Davis, José M. Fernandez, and Sonia Chiasson. Methodology for a field study of anti-malware software. Lecture Notes in Computer Science, 7398:80–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-34638-5_7/.

Patil:2012:MPP

[962] Sameer Patil, Yann Le Gall, Adam J. Lee, and Apu Kapadia. My privacy policy: Exploring end-user specification of free-form location access rules. Lecture Notes in Computer Science, 7398:86–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-34638-5_8/.

West:2012:SSA

[963] Andrew G. West, Pedram Hayati, Vidyasagar Potdar, and Insup Lee. Spamming for science: Active measurement in Web 2.0 abuse research. Lecture Notes in Computer Science, 7398:98–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-34638-5_9/.

Bailey:2012:REI

[964] Michael Bailey, Erin Kenneally, and David Dittrich. A refined ethical impact assessment tool and a case study of its application. Lecture Notes in Computer Science, 7398:112–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34638-5_10/.

Egelman:2012:SIY

[965] Serge Egelman, Joseph Bonneau, Sonia Chiasson, David Dittrich, and Stuart Schechter. It’s not stealing if you need it: a panel on the ethics of performing research using public data of illicit origin. Lecture Notes in Computer Science, 7398:124–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34638-5_11/.

Anderson:2012:ECI

[966] Ross Anderson. Ethics committees and IRBs: Boon, or bane, or more research needed? Lecture Notes in Computer Science, 7398:133–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34638-5_12/.

Fernandez:2012:ESD

[967] José M. Fernandez, Andrew S. Patrick, and Lenore D. Zuck. Ethical and secure data sharing across borders. Lecture Notes in Computer Science, 7398:136–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34638-5_13/.

Anonymous:2012:BMo

[968] Anonymous. Back matter. Lecture Notes in Computer Science, 7398: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34638-5/1.

Anonymous:2012:FMcb

[969] Anonymous. Front matter. Lecture Notes in Computer Science, 7398: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34638-5/1.

Abramsky:2012:LBK

[970] Samson Abramsky and Jonathan Zvesper. From Lawvere to Brandenburger-Keisler: Interactive forms of diagonalization and self-reference. Lecture Notes in Computer Science, 7399:1–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32784-1_1/.

Bonsangue:2012:DCF

[971] Marcello M. Bonsangue, Jan Rutten, and Joost Winter. Defining context-free power series coalgebraically. Lecture Notes in Computer Science, 7399:20–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32784-1_2/.

Sobocinski:2012:RPL

[972] Paweł Sobociński. Relational presheaves as labelled transition systems. Lecture Notes in Computer Science, 7399:40–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32784-1_3/.

Adamek:2012:FFT

[973] Jiří Adámek, Stefan Milius, and Lawrence S. Moss. On finitary functors and their presentations. Lecture Notes in Computer Science, 7399:51– 70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32784-1_4/.

Bacci:2012:SOS

[974] Giorgio Bacci and Marino Miculan. Structural operational semantics for continuous state probabilistic processes. Lecture Notes in Computer Science, 7399:71–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32784-1_5/.

Ciancia:2012:SAC

[975] Vincenzo Ciancia and Yde Venema. Stream automata are coalgebras. Lecture Notes in Computer Science, 7399:90–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32784-1_6/.

Jacobs:2012:TSD

[976] Bart Jacobs, Alexandra Silva, and Ana Sokolova. Trace semantics via determinization. Lecture Notes in Computer Science, 7399:109–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32784-1_7/.

Kurz:2012:ACP

[977] Alexander Kurz, Daniela Petrişan, and Paula Severi. An alphacorecursion principle for the infinitary lambda calculus. Lecture Notes in Computer Science, 7399:130–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32784-1_8/.

Marti:2012:LEC

[978] Johannes Marti and Yde Venema. Lax extensions of coalgebra functors. Lecture Notes in Computer Science, 7399:150–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32784-1_9/.

Matsikoudis:2012:TE

[979] Eleftherios Matsikoudis and Edward A. Lee. From transitions to executions. Lecture Notes in Computer Science, 7399:170–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32784-1_10/.

Pavlovic:2012:TMM

[980] Dusko Pavlovic. Tracing the man in the middle in monoidal categories. Lecture Notes in Computer Science, 7399:191–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32784-1_11/.

Picard:2012:PCG

[981] Celia Picard and Ralph Matthes. Permutations in coinductive graph representation. Lecture Notes in Computer Science, 7399:218–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32784-1_12/.

Wilkinson:2012:IMC

[982] Toby Wilkinson. Internal models for coalgebraic modal logics. Lecture Notes in Computer Science, 7399:238–258, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32784-1_13/.

Anonymous:2012:BMp

[983] Anonymous. Back matter. Lecture Notes in Computer Science, 7399: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32784-1/1.

Anonymous:2012:FMcc

[984] Anonymous. Front matter. Lecture Notes in Computer Science, 7399: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32784-1/1.

Ardagna:2012:BSM

[985] Claudio A. Ardagna, Ernesto Damiani, Leszek A. Maciaszek, Michele Missikoff, and Michael Parkin, editors. Business System Management and Engineering: From Open Issues to Applications, volume 7350 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32438-X (print), 3-642-32439-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-32439-0.

Bechet:2012:LAC

[986] Denis Béchet and Alexander Dikovsky, editors. Logical Aspects of Computational Linguistics: 7th International Conference, LACL 2012, Nantes, France, July 2–4, 2012. Proceedings, volume 7351 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31261-6 (print), 3-642-31262-4 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-31262-5.

Cox:2012:DRI

[987] Philip Cox, Beryl Plimmer, and Peter Rodgers, editors. Diagrammatic Representation and Inference: 7th International Conference, Diagrams 2012, Canterbury, UK, July 2–6, 2012. Proceedings, volume 7352 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31222-5 (print), 3-642-31223-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-31223-6.

Hirsch:2012:CST

[988] Edward A. Hirsch, Juhani Karhumäki, Arto Lepistö, and Michail Prilutskii, editors. Computer Science — Theory and Applications: 7th International Computer Science Symposium in Russia, CSR 2012, Nizhny Novgorod, Russia, July 3–7, 2012. Proceedings, volume 7353 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30641-1 (print), 3-642-30642-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-30642-6.

Karkkainen:2012:CPM

[989] Juha Kärkkäinen and Jens Stoye, editors. Combinatorial Pattern Matching: 23rd Annual Symposium, CPM 2012, Helsinki, Finland, July 3–5, 2012. Proceedings, volume 7354 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31264-0 (print), 3-642-31265-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-31265-6.

Even:2012:SIC

[990] Guy Even and Magnús M. Halldórsson, editors. Structural Information and Communication Complexity: 19th International Colloquium, SIROCCO 2012, Reykjavik, Iceland, June 30–July 2, 2012, Revised Selected Papers, volume 7355 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31103-2 (print), 3-642-31104-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-31104-8.

Salampasis:2012:MIR

[991] Michail Salampasis and Birger Larsen, editors. Multidisciplinary Information Retrieval: 5th International Retrieval Facility Conference, IRFC 2012, Vienna, Austria, July 2–3, 2012 Proceedings, volume 7356 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31273-X (print), 3-642-31274-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-31274-8.

Fomin:2012:ATS

[992] Fedor V. Fomin and Petteri Kaski, editors. Algorithm Theory — SWAT 2012: 13th Scandinavian Symposium and Workshops, Helsinki, Finland, July 4–6, 2012. Proceedings, volume 7357 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31154-7 (print), 3-642-31155-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-31155-0.

Madhusudan:2012:CAV

[993] P. Madhusudan and Sanjit A. Seshia, editors. Computer Aided Verification: 24th International Conference, CAV 2012, Berkeley, CA, USA, July 7–13, 2012 Proceedings, volume 7358 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31423-6 (print), 3-642-31424-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-31424-7.

Dawant:2012:BIR

[994] Benoît M. Dawant, Gary E. Christensen, J. Michael Fitzpatrick, and Daniel Rueckert, editors. Biomedical Image Registration: 5th International Workshop, WBIR 2012, Nashville, TN, USA, July 7–8, 2012. Proceedings, volume 7359 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31339-6 (print), 3-642-31340-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31340-0.

Artikis:2012:LPN

[995] Alexander Artikis, Robert Craven, Nihan Kesim Çiçekli, Babak Sadighi, and Kostas Stathis, editors. Logic Programs, Norms and Action: Essays in Honor of Marek J. Sergot on the Occasion of His 60th Birthday, volume 7360 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29413-8 (print), 3-642-29414-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-29414-3.

Maidment:2012:BII

[996] Andrew D. A. Maidment, Predrag R. Bakic, and Sara Gavenonis, editors. Breast Imaging: 11th International Workshop, IWDM 2012, Philadelphia, PA, USA, July 8–11, 2012. Proceedings, volume 7361 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31270-5 (print), 3-642-31271-3 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-31271-7.

Jeuring:2012:ICM

[997] Johan Jeuring, John A. Campbell, Jacques Carette, Gabriel Dos Reis, Petr Sojka, Makarius Wenzel, and Volker Sorge, editors. Intelligent Computer Mathematics: 11th International Conference, AISC 2012, 19th Symposium, Calculemus 2012, 5th International Workshop, DML 2012, 11th International Conference, MKM 2012, Systems and Projects, Held as Part of CICM 2012, Bremen, Germany, July 8–13, 2012. Proceedings, volume 7362 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31373-6 (print), 3-642-31374-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31374-5.

Li:2012:AHM

[998] Xiang-Yang Li, Symeon Papavassiliou, and Stefan Ruehrup, editors. Adhoc, Mobile, and Wireless Networks: 11th International Conference, ADHOC-NOW 2012, Belgrade, Serbia, July 9–11, 2012. Proceedings, volume 7363 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31637-9 (print), 3-642-31638-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31638-8.

Gramlich:2012:ARI

[999] Bernhard Gramlich, Dale Miller, and Uli Sattler, editors. Automated Reasoning: 6th International Joint Conference, IJCAR 2012, Manchester, UK, June 26–29, 2012. Proceedings, volume 7364 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31364-7 (print), 3-642-31365-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31365-3.

Heisel:2012:SSA

[1000] Maritta Heisel, editor. Software Service and Application Engineering: Essays Dedicated to Bernd Krämer on the Occasion of His 65th Birthday, volume 7365 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30834-1 (print), 3-642-30835-X (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-30835-2.

Zhang:2012:ABI

[1001] Huaguang Zhang, Amir Hussain, Derong Liu, and Zhanshan Wang, editors. Advances in Brain Inspired Cognitive Systems: 5th International Conference, BICS 2012, Shenyang, China, July 11–14, 2012. Proceedings, volume 7366 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31560-7 (print), 3-642-31561-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31561-9.

Wang:2012:ANNa

[1002] Jun Wang, Gary G. Yen, and Marios M. Polycarpou, editors. Advances in Neural Networks — ISNN 2012: 9th International Symposium on Neural Networks, Shenyang, China, July 11–14, 2012. Proceedings, Part I, volume 7367 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31345-0 (print), 3-642-31346-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31346-2.

Wang:2012:ANNb

[1003] Jun Wang, Gary G. Yen, and Marios M. Polycarpou, editors. Advances in Neural Networks — ISNN 2012: 9th International Symposium on Neural Networks, Shenyang, China, July 11–14, 2012. Proceedings, Part II, volume 7368 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31361-2 (print), 3-642-31362-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31362-2.

Ozbudak:2012:AFF

[1004] Ferruh Özbudak and Francisco Rodríguez-Henríquez, editors. Arithmetic of Finite Fields: 4th International Workshop, WAIFI 2012, Bochum, Germany, July 16–19, 2012. Proceedings, volume 7369 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31661-1 (print), 3-642-31662-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31662-3.

Bodlaender:2012:MAR

[1005] Hans L. Bodlaender, Rod Downey, Fedor V. Fomin, and Dániel Marx, editors. The Multivariate Algorithmic Revolution and Beyond: Essays Dedicated to Michael R. Fellows on the Occasion of His 60th Birthday, volume 7370 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30890-2 (print), 3-642-30891-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-30891-8.

Cuppens-Boulahia:2012:DAS

[1006] Nora Cuppens-Boulahia, Frédéric Cuppens, and Joaquin Garcia-Alfaro, editors. Data and Applications Security and Privacy XXVI: 26th Annual IFIP WG 11.3 Conference, DBSec 2012, Paris, France, July 11–13,2012. Proceedings, volume 7371 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31539-9 (print), 3-642-31540-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-31540-4.

Susilo:2012:ISP

[1007] Willy Susilo, Yi Mu, and Jennifer Seberry, editors. Information Security and Privacy: 17th Australasian Conference, ACISP 2012, Wollongong, NSW, Australia, July 9–11, 2012. Proceedings, volume 7372 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31447-3 (print), 3-642-31448-1 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-31448-3.

Rahaman:2012:PVD

[1008] Hafizur Rahaman, Sanatan Chattopadhyay, and Santanu Chattopadhyay, editors. Progress in VLSI Design and Test: 16th International Symposium, VDAT 2012, Shibpur, India, July 1–4, 2012. Proceedings, volume 7373 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31493-7 (print), 3-642-31494-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31494-0.

Mitrokotsa:2012:PCA

[1009] Aikaterini Mitrokotsa and Serge Vaudenay, editors. Progress in Cryptology — AFRICACRYPT 2012: 5th International Conference on Cryptology in Africa, Ifrance, Morocco, July 10–12, 2012. Proceedings, volume 7374 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31409-0 (print), 3-642-31410-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31410-0.

Prescott:2012:BBS

[1010] Tony J. Prescott, Nathan F. Lepora, Anna Mura, and Paul F. M. J. Verschure, editors. Biomimetic and Biohybrid Systems: First International Conference, Living Machines 2012, Barcelona, Spain, July 9–12, 2012. Proceedings, volume 7375 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31524-0 (print), 3-642-31525-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-31525-1.

Perner:2012:MLD

[1011] Petra Perner, editor. Machine Learning and Data Mining in Pattern Recognition: 8th International Conference, MLDM 2012, Berlin, Germany, July 13–20, 2012. Proceedings, volume 7376 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-31536-4 (print), 3-642-31537-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31537-4.

Perner:2012:ADM

[1012] Petra Perner, editor. Advances in Data Mining. Applications and Theoretical Aspects: 12th Industrial Conference, ICDM 2012, Berlin, Germany, July 13–20, 2012. Proceedings, volume 7377 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-31487-2 (print), 3-642-31488-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31488-9.

Perales:2012:AMD

[1013] Francisco J. Perales, Robert B. Fisher, and Thomas B. Moeslund, editors. Articulated Motion and Deformable Objects: 7th International Conference, AMDO 2012, Port d’Andratx, Mallorca, Spain, July 11– 13, 2012. Proceedings, volume 7378 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31566-6 (print), 3-642-31567-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-31567-1.

Masthoff:2012:UMA

[1014] Judith Masthoff, Bamshad Mobasher, Michel C. Desmarais, and Roger Nkambou, editors. User Modeling, Adaptation, and Personalization: 20th International Conference, UMAP 2012, Montreal, Canada, July 16–20, 2012. Proceedings, volume 7379 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31453-8 (print), 3-642-31454-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-31454-4.

Gavrilova:2012:TCS

[1015] Marina L. Gavrilova and C. J. Kenneth Tan, editors. Transactions on Computational Science XVI, volume 7380 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32662-5 (print), 3-642-32663-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-32663-9.

Moreira:2012:IAA

[1016] Nelma Moreira and Rogério Reis, editors. Implementation and Application of Automata: 17th International Conference, CIAA 2012, Porto, Portugal, July 17–20, 2012. Proceedings, volume 7381 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012.

CODEN LNCSD9. ISBN 3-642-31605-0 (print), 3-642-31606-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31606-7.

Miesenberger:2012:CHPa

[1017] Klaus Miesenberger, Arthur Karshmer, Petr Penaz, and Wolfgang Zagler, editors. Computers Helping People with Special Needs: 13th International Conference, ICCHP 2012, Linz, Austria, July 11–13, 2012, Proceedings, Part I, volume 7382 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31521-6 (print), 3-642-31522-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31522-0.

Miesenberger:2012:CHPb

[1018] Klaus Miesenberger, Arthur Karshmer, Petr Penaz, and Wolfgang Zagler, editors. Computers Helping People with Special Needs: 13th International Conference, ICCHP 2012, Linz, Austria, July 11–13, 2012, Proceedings, Part II, volume 7383 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31533-X (print), 3-642-31534-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31534-3.

Fischer-Hubner:2012:PET

[1019] Simone Fischer-Hübner and Matthew Wright, editors. Privacy Enhancing Technologies: 12th International Symposium, PETS 2012, Vigo, Spain, July 11–13, 2012. Proceedings, volume 7384 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31679-4 (print), 3-642-31680-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31680-7.

Donaldson:2012:MCS

[1020] Alastair Donaldson and David Parker, editors. Model Checking Software: 19th International Workshop, SPIN 2012, Oxford, UK, July 23–24, 2012. Proceedings, volume 7385 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31758-8 (print), 3-642-31759-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-31759-0.

Kutrib:2012:DCF

[1021] Martin Kutrib, Nelma Moreira, and Rogério Reis, editors. Descriptional Complexity of Formal Systems: 14th International Workshop, DCFS 2012, Braga, Portugal, July 23–25, 2012. Proceedings, volume 7386 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31622-0 (print), 3-642-31623-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-31623-4.

Brambilla:2012:WEI

[1022] Marco Brambilla, Takehiro Tokuda, and Robert Tolksdorf, editors. Web Engineering: 12th International Conference, ICWE 2012, Berlin, Germany, July 23–27, 2012. Proceedings, volume 7387 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-31752-9 (print), 3-642-31753-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-31753-8.

Bezhanishvili:2012:LLC

[1023] Nick Bezhanishvili and Valentin Goranko, editors. Lectures on Logic and Computation: ESSLLI 2010 Copenhagen, Denmark, August 2010, ESS-LLI 2011, Ljubljana, Slovenia, August 2011, Selected Lecture Notes, volume 7388 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31484-8 (print), 3-642-31485-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-31485-8.

Huang:2012:ICTa

[1024] De-Shuang Huang, Changjun Jiang, Vitoantonio Bevilacqua, and Juan Carlos Figueroa, editors. Intelligent Computing Technology: 8th International Conference, ICIC 2012, Huangshan, China, July 25–29, 2012. Proceedings, volume 7389 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31587-9 (print), 3-642-31588-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-31588-6.

Huang:2012:ICTb

[1025] De-Shuang Huang, Jianhua Ma, Kang-Hyun Jo, and M. Michael Gromiha, editors. Intelligent Computing Theories and Applications: 8th International Conference, ICIC 2012, Huangshan, China, July 25– 29, 2012. Proceedings, volume 7390 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31575-5 (print), 3-642-31576-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-31576-3.

Czumaj:2012:ALPa

[1026] Artur Czumaj, Kurt Mehlhorn, Andrew Pitts, and Roger Wattenhofer, editors. Automata, Languages, and Programming: 39th International Colloquium, ICALP 2012, Warwick, UK, July 9–13, 2012, Proceedings, Part I, volume 7391 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31593-3 (print), 3-642-31594-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31594-7.

Czumaj:2012:ALPb

[1027] Artur Czumaj, Kurt Mehlhorn, Andrew Pitts, and Roger Wattenhofer, editors. Automata, Languages, and Programming: 39th International Colloquium, ICALP 2012, Warwick, UK, July 9–13, 2012, Proceedings, Part II, volume 7392 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31584-4 (print), 3-642-31585-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-31585-5.

Aagotnes:2012:DLC

[1028] Thomas Ågotnes, Jan Broersen, and Dag Elgesem, editors. Deontic Logic in Computer Science: 11th International Conference, DEON 2012, Bergen, Norway, July 16–18, 2012. Proceedings, volume 7393 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31569-0 (print), 3-642-31570-4 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-31570-1.

DeDecker:2012:CMS

[1029] Bart De Decker and David W. Chadwick, editors. Communications and Multimedia Security: 13th IFIP TC 6/TC 11 International Conference, CMS 2012, Canterbury, UK, September 3–5, 2012. Proceedings, volume 7394 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32804-0 (print), 3-642-32805-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-32805-3.

deGroote:2012:FGI

[1030] Philippe de Groote and Mark-Jan Nederhof, editors. Formal Grammar: 15th and 16th International Conferences, FG 2010, Copenhagen, Denmark, August 2010, FG 2011, Ljubljana, Slovenia, August 2011, Revised Selected Papers, volume 7395 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32023-6 (print), 3-642-32024-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-32024-8.

Huusko:2012:EED

[1031] Jyrki Huusko, Hermann de Meer, Sonja Klingert, and Andrey Somov, editors. Energy Efficient Data Centers: First International Workshop, E 2 DC 2012, Madrid, Spain, Mai 8, 2012, Revised Selected Papers, volume 7396 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33644-2 (print), 3-642-33645-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33645-4.

Keromytis:2012:FCD

[1032] Angelos D. Keromytis, editor. Financial Cryptography and Data Security: 16th International Conference, FC 2012, Kralendijk, Bonaire, Februray 27–March 2, 2012, Revised Selected Papers, volume 7397 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32945-4 (print), 3-642-32946-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-32946-3.

Blyth:2012:FCD

[1033] Jim Blyth, Sven Dietrich, and L. Jean Camp, editors. Financial Cryptography and Data Security: FC 2012 Workshops, USEC and WECSR 2012, Kralendijk, Bonaire, March 2, 2012, Revised Selected Papers, volume 7398 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34637-5 (print), 3-642-34638-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-34638-5.

Pattinson:2012:CMC

[1034] Dirk Pattinson and Lutz Schröder, editors. Coalgebraic Methods in Computer Science: 11th International Workshop, CMCS 2012, Colocated with ETAPS 2012, Tallinn, Estonia, March 31 — April 1, 2012, Revised Selected Papers, volume 7399 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32783-4 (print), 3-642-32784-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-32784-1.